### v1

# Automatic Speech Recognition using Advanced Deep Learning Approaches: A survey

Abstract
Recent advancements in deep learning (DL) have posed a significant challenge for automatic speech recognition (ASR). ASR relies on extensive training datasets, including confidential ones, and demands substantial computational and storage resources. Enabling adaptive systems improves ASR performance in dynamic environments. DL techniques assume training and testing data originate from the same domain, which is not always true. Advanced DL techniques like deep transfer learning (DTL), federated learning (FL), and reinforcement learning (RL) address these issues. DTL allows high-performance models using small yet related datasets, FL enables training on confidential data without dataset possession, and RL optimizes decision-making in dynamic environments, reducing computation costs.
This survey offers a comprehensive review of DTL, FL, and RL-based ASR frameworks, aiming to provide insights into the latest developments and aid researchers and professionals in understanding the current challenges. Additionally, transformers, which are advanced DL techniques heavily used in proposed ASR frameworks, are considered in this survey for their ability to capture extensive dependencies in the input ASR sequence. The paper starts by presenting the background of DTL, FL, RL, and Transformers and then adopts a well-designed taxonomy to outline the state-of-the-art (SOTA) approaches. Subsequently, a critical analysis is conducted to identify the strengths and weaknesses of each framework. Additionally, a comparative study is presented to highlight the existing challenges, paving the way for future research opportunities.
keywords:
[…]
1 Introduction
1.1 Preliminary
Advancements in artificial intelligence (AI) have significantly improved human-machine interaction (HMI), especially with technologies that convert speech into executable actions. automatic speech recognition (ASR) emerges as a leading communication technology in HMI, extensively utilized by corporations and service providers for facilitating interactions through AI platforms like chatbots and digital assistants. Spoken language forms the core of these interactions, emphasizing the necessity for sophisticated speech processing in AI systems tailored for ASR. ASR technology encompasses the analysis of (i) acoustic, lexical, and syntactic aspects; and (ii) semantic understanding. The acoustic model (AM) processing includes speech coding [1], enhancement [2], and source separation [3], alongside securing speech via steganography [4, 5, 6] and watermarking [7, 8, 9]. These components are integral to audio analysis. On the other hand, the semantic model (SM), often identified as language model (LM) processing in literature, involves all natural language processing (NLP) techniques. This AI branch aims at teaching computers to understand and interpret human language, serving as the basis for applications like music information retrieval [10], sound file organization [11], audio tagging, and event detection (ED) [12], as well as converting speech to text and vice versa [13], detecting hate speech [14], and cyberbullying [15]. Employing NLP across various domains enables AI models to effectively comprehend and respond to human inputs, unveiling extensive research prospects in diverse sectors.
Abbreviations
AI
artificial intelligence
AM
acoustic model
APT
audio pyramid transformer
ASR
automatic speech recognition
AST
audio spectrogram transformer
AT
audio tagging
CAFT
client adaptive federated training
CER
character error rate
CNN
convolutional neural network
CS
code-switching
CTC
connectionist temporal classification
CV
computer vision
DA
domain adaptation
DL
deep learning
DRL
deep reinforcement learning
DTL
deep transfer learning
ED
event detection
FCF
feature correlation-based fusion
FedNST
federated noisy student training
FL
federated learning
FR
form recognition
HMI
human-machine interaction
HMM
hidden Markov models
LM
language model
LPC
linear predictive coding
LSTM
long short term memory
mAP
mean average precision
MFCC
Mel-frequency cepstral coefficient
MHSA
multi-head self-attention
ML
machine learning
NLP
natural language processing
NT
negative transfer
PESQ
perceptual evaluation of speech quality
RER
relative error rate
RL
reinforcement learning
RNN
recurrent neural network
RTF
real-time factor
S2S
sequence-to-sequence
S2S
sequence-to-sequence
SD
source domain
SER
speech emotion recognition
SM
semantic model
SNR
signal-to-noise ratio
SOTA
state-of-the-art
SS
speech security
SSAST
self-supervised audio spectrogram transformer
SWBD
switchboard
TD
target domain
TNR
true negative rate
TPR
true positive rate
TRUNet
transformer-recurrent-U network
WER
word error rate
WSJ
wall street journal
GAN
generative adversarial network
MTL
multitask learning
FMTL
federated multi-task learning
DSLM
domain-specific language modeling
LLM
large language model
DDQN
double deep Q-network
AC
actor-critic
SARSA
State–action–reward–state–action
DDPG
deep deterministic policy gradien
Recent advancements in ASR have been significantly propelled by the evolution of deep learning (DL) methodologies. An extensive range of DL models has been developed, demonstrating remarkable improvements and surpassing former SOTA achievements [16, 17]. Transformers, a notable innovation within these DL approaches, have become a cornerstone in advancing various NLP tasks, including ASR. Initially conceptualized for sequence-to-sequence (S2S) applications in NLP, their success is largely attributed to their adeptness at discerning long-range dependencies and complex patterns within sequential data. A hallmark of transformer models is their utilization of an attention mechanism, which precisely focuses on specific portions of the input sequence during prediction tasks. This mechanism is particularly effective in ASR, facilitating the detailed modeling of contextual nuances and the interconnections among acoustic signals, essential for accurate transcription. Models such as the Transformer Transducer, Conformer, and ESPnet, leveraging self-attention and parallel processing, have achieved leading performance in ASR tasks. Their robustness across diverse languages further underscores their capability to adapt to a wide range of linguistic features and acoustic variations, making transformers an exceptionally promising option for enhancing ASR systems, surpassing the constraints of conventional models.
The integration of DL with its variants in ASR introduces substantial challenges, especially concerning its application in natural HMI. Despite DL’s numerous advantages, it encounters various obstacles. The inherent complexity of DL models, which stems from their need for extensive training data to attain high performance, demands significant computational and storage resources [18]. Moreover, the issue of data scarcity in ASR reflects the inadequate quantities of training data available for exploiting complex DL algorithms effectively [19]. The paucity of annotated data further complicates the development of supervised DL-based ASR models. Additionally, the presumption that training and testing datasets originate from the same domain, possessing identical feature spaces and distribution characteristics, is often misguided. This mismatch challenges the practical deployment of DL models in real-world settings [20]. Thus, the performance of DL models may be compromised when faced with limited training datasets or discrepancies in data distribution between training and testing environments [21]. These challenges highlight the critical need for adaptive methodologies and improved data management approaches to fully harness the capabilities of DL in ASR systems.
In an effort to address existing challenges and increase the robustness and flexibility of ASR systems, novel DL methodologies have been introduced. These include deep transfer learning (DTL), deep reinforcement learning (DRL), and federated learning (FL), which collectively aim at overcoming difficulties related to the transfer of knowledge, enhancing the generalization capabilities of models, and optimizing training processes. These innovative approaches significantly broaden the operational scope of conventional DL frameworks within the ASR field.
Figure 1 highlights critical areas in speech processing where DTL, DRL, and FL can be applied. Consequently, domains such as ASR, speech enhancement (SE), hate speech detection (HSD), and speech security (SS) are closely interconnected. ASR provides acoustic parameters to NLP for HSD task, which in turn provides semantic details to ASR. Additionally, ASR can be employed in the SS domain as a steganalytic process to verify the integrity of speech [4, 22]. Furthermore, ASR and SE can mutually offer performance feedback.
Figure 1: Summary of critical areas in speech processing where DTL, DRL, FL, and transformers can be applied.
1.2 Contribution of the paper
This article offers an extensive examination of contemporary frameworks within advanced deep learning approaches, spanning the period from 2016 to 2023. These approaches include DTL, DRL, FL, and Transformers, all within the context of ASR. To the best of the authors’ knowledge, there has been no prior research paper that has intricately explored and critically evaluated contributions in the aforementioned advanced DL-based ASR until now.
In recent years, numerous survey papers have been published to assess various aspects of ASR models. Some of these surveys concentrate on specific languages, such as Portuguese [23], Indian [24], Turkish [25], Arabic [26] and tonal languages (including Asian, Indo-European and African) [27]. Additionally, Abushariah et al.’s review emphasizes bilingual ASR [28]. On the non-specific language review front, specific areas within ASR have been targeted, including ASR using limited vocabulary [29], ASR for children [30], error detection and correction [31], and unsupervised ASR [32]. Systematic reviews with a focus on neural networks [33] and deep neural networks [34] have also been proposed. In another comprehensive review, Malik et al. [35] discussed diverse feature extraction methods, SOTA classification models, and some deep learning approaches. Recently, the authors presented an ASR review focused on DTL for ASR [36]. Table 1 presents a summary of the main contributions of the proposed ASR review compared to other existing ASR reviews/surveys.
This survey article offers several significant enhancements and additions compared to previous ASR surveys. Firstly, it consolidates works that utilize both ASR and advanced DL approaches, providing a comprehensive overview of their intersection. Secondly, it provides performance evaluation results of all considered approaches. Thirdly, it includes metrics and dataset reviews used in ASR models. Furthermore, it tackles ongoing challenges and consequently proposes future directions. The main contributions of this article can be summarized as follows:
  • Presenting the background of advanced DL techniques including DTL, DRL, FL and transformers. Describing the evaluation metrics and datasets employed for validating ASR approaches.
  • Introducing a well-defined taxonomy categorizing ASR methodologies based on the domains of AM and LM.
  • Identifying challenges and gaps in advanced DL-based ASR.
  • Proposing future directions to enhance the performance of advanced DL-based ASR solutions and predicting the potential advancements in the field.
Table 1: Contribution comparison of the proposed contribution against other hand ASR review. The tick mark (✓) indicates that the specific field has been addressed, whereas the cross mark () means addressing the specific fields has been missed.
Refs Year
Description of the survey/review
Advanced DL methods
Performances
Metrics
Dataset
Current
Future
DRL
FL
TL
Transf.
evaluation
review
challe/Gaps
directions
[31] 2018
ASR review for error errors detection and correction
[34] 2019
Systematic review on DL-based speech recognition
[24] 2020
ASR survey for Indian languages
[23] 2020
ASR survey for Portuguese language
[25] 2020
ASR survey for Turkich language
[35] 2021
ASR survey
[27] 2021
ASR survey for tonal languages
[32] 2022
Unsupervised ASR review
[30] 2022
ASR Systematic review for children
[29] 2022
ASR survey for limited vocabulary
[26] 2022
ASR Systematic review for Arabic language
[28] 2022
Bilingual ASR review
[33] 2023
ASR survey on neural network techniques
[36] 2023
ASR based on DTL review
Our 2024
ASR review on advanced DL techniques
1.3 Review methodology
The methodology for the review is delineated in this segment, encompassing the search strategy and study selection. Inclusion criteria, comprising keyword alignment, creativity and impact, and uniqueness, are explicated, collectively influencing the formulation of the paper’s quality assessment protocol. To locate and compile extant advanced DL-based ASR studies, a thorough search was executed on renowned publication databases recognized for hosting top-tier scientific research articles. The exploration encompassed Scopus and Web of Science. Keywords were extracted and organized from the initial set of references through manual analysis. Employing "theme clustering," these publications were sorted based on keywords found in the "Abstract," "Title," and "Authors keywords." The outcome of this process yielded the formulation of the following query:
References=FROM "Abstract" || "Title"|| "Authors keywords" SELECT( Papers WHERE keywords= (ASR || NLP) & ( DTL || DRL || FL || Transformers)).
The symbols || and & denote OR and AND logical operations, respectively. The evaluation of publications considered the innovation level in ASR, the study’s quality, and the contributions and findings presented. This review exclusively encompassed research contributions that were published within the timeframe of 2016 to 2023.
1.4 Structure of the paper
This paper is structured into six sections. The current section provides an introduction to the paper. Section 1.3 providing background on AM and LM, and reviewing evaluation metrics and datasets utilized in ASR. Moving forward, Section 4 delves into a comprehensive review of recent advancements in ASR utilizing advanced DL approaches, including Transformers, DTL, FL and DRL. Sections 5 and 6 respectively address the existing challenges and future directions concerning advanced DL-based ASR. Finally, Section 7 presents concluding remarks summarizing the key findings of the paper.
Figure 2: Survey roadmap: A guide for navigating paper sections and subsections.
table[t!] Literature acquisition databases. to be modified Database Research Articles Conference Papers Book Chapter Total ACM 5 1 6 Elsevier 26 26 Springer 23 10 33 IEEE 39 28 67 Others 44 29 70
2 Overview of DTL techniques for speech recognition
2.1 Taxonomy of existing DTL techniques
To date, there is no standardized and comprehensive technique for classifying DTL into categories. However, DTL algorithms could be classified into several types depending on what, when, and how knowledge is transferred:
(a) What knowledge is transferred? Enquires about which characteristics of knowledge are transferable across domains or tasks. Some information is particular to certain domains or tasks, while other knowledge is shared across domains and can aid increase performance in the target task or domain. Based on this definition, DTL could be feature-based, instance-based, relation-based, or model-based.
(b) How is knowledge transferred? Enquires about which learning algorithms must be implemented to transfer the knowledge.
(c) When is knowledge transferred? Inquires as to when and under what circumstances knowledge should or should not be transferred.
Furthermore, researchers have proposed taxonomies to categorize DTL-based ASR techniques. For instance, Niu et al. [21] present a taxonomy with two levels. The first level consists of four sub-groups based on the availability of labeled data and the data modality in the source and target domains. These sub-groups include inductive DTL, transductive DTL, cross-modality DTL, and unsupervised DTL [80]. Table 2 provides a summary of these possibilities. Moreover, each sub-group at the first level can be further divided into four distinct learning types: learning on instances, learning on features, learning on parameters, and learning on relations.
[…]
Domains Tasks
Math. propriety
Sub-categories / Usage
Traditional ML/DL $\mathbb{D}_{S}=\mathbb{D}_{T}$ $\mathbb{T}_{S}=\mathbb{T}_{T}$
$X_{S}\neq X_{T}$,
$Y_{S}=Y_{T}$
ASR model trained with $X_{S}$ database and used to recognise $X_{T}$ database.
Inductive DTL $\mathbb{D}_{S}\cong\mathbb{D}_{T}$ $\mathbb{T}_{S}\neq\mathbb{T}_{T}$
$X_{S}\neq X_{T}$,
$Y_{S}\exists,Y_{T}\exists$
If $Y_{S}\exists$, DTL is multitask learning. If $Y_{S}\nexists$, DTL is self-taught learning, thus $\chi_{S}\cong\chi_{T}$.
Transductive DTL $\mathbb{D}_{S}\neq\mathbb{D}_{T}$ $\mathbb{T}_{S}=\mathbb{T}_{T}$
$P(X_{S})\neq P(X_{T})$,
$Y_{S}\exists,Y_{T}\nexists$ ,
$\chi_{S}=\chi_{T}$
When $\chi_{S}=\chi_{T}$, DTL is is related to DA. If $\mathbb{D}_{T}\exists!$ and $\mathbb{T}_{T}\exists!$, DTL is used for sample selection bias or covariate shift.
Cross-modality DTL $\mathbb{D}_{S}\neq\mathbb{D}_{T}$ $\mathbb{T}_{S}\neq\mathbb{T}_{T}$
$P(Y_{S}/X_{S})\neq P(Y_{T}/X_{T})$,
$Y_{S}\neq Y_{T}$, $\chi_{S}\neq\chi_{T}$
i,e. the dataset $X_{S}$ of $\mathbb{D}_{S}$ is speech data, and the dataset $X_{T}$ of $\mathbb{D}_{T}$ is text data.
Unsupervised DTL $\mathbb{D}_{S}\varsubsetneq\mathbb{D}_{T}$ $\mathbb{T}_{S}\varsubsetneq\mathbb{T}_{T}$
$Y_{S}\nexists,Y_{T}\nexists$
DTL used for clustering, dimensionality reduction, and density estimation, etc.
Inductive DTL In comparison to classical ML, which may be used as a reference for DTL comparison, and given that the target tasks $\mathbb{T}_{T}$ are distinct from the source tasks $\mathbb{T}_{S}$, the goal of inductive DTL is to enhance the target prediction function $\mathbb{F}_{T}$ in the TD, mentioned above in subsection LABEL:sub12. However, the SD $\mathbb{D}_{S}$ and TD $\mathbb{D}_{T}$ may not always be the same (Table 2). The inductive DTL can be stated similarly to the following two cases, depending on whether labeled or unlabeled data is available:
(a) Multi-task DTL: The SD has a huge labeled database ($X_{S}$ labeled with $Y_{S}$), which is a distinctive form of multi-task learning. However, with multi-task approaches, many tasks $(T_{1},T_{2},\dots,T_{n})$ are learned at the same time (in parallel), including both source and target activities (tasks).
(b) Sequential DTL: ( Commonly known as self-taught learning) Dataset is not labeled in the SD ($X_{S}$ is not labeled with $Y_{S}$) but the labels are available in the destination domain ($X_{T}$ is labeled with $Y_{T}$). Sequential learning is a DL system that can be realized in two steps for classification purposes. The first step is the feature representation transfer, which is learned from a large collection of the unlabeled datasets, and the second stage is when this learned representation is applied to labeled data to accomplish classification tasks. Hence, sequential DTL is a method of sequentially learning a number of activities (Tasks). The spaces between the source and destination domains may differ. For example, let suppose we have a pre-trained model $M$ and consider applying DTL to a number of tasks $(T_{1},T_{2},\dots,T_{n})$. We learn a specific task $\mathbb{T}_{T}$ at each time step $t$, which is slower than multi-task learning. However, when not all the tasks are present during training time, it might be beneficial. Sequential DTL is additionally classified into several types [alyafeai2020survey]:
  • 1-
    Fine-tuning: The principle is to learn a new function $\mathbb{F}_{T}$ that translates the parameters $\mathbb{F}_{T}(W_{S})=W_{T}$ by using $M$, given a pre-trained model $M_{S}$ having $W_{S}$ as weights and target task $\mathbb{T}_{T}$ having $W_{T}$ as weights. The settings can be adjusted across all layers or just some of (Fig. 3 (a)). The learning rate for each layer could be distinct (discriminative fine tuning). A new set of parameters $K$ could be added to most of the tasks so that $\mathbb{F}_{T}(W_{T},K)=W_{S}\circ K$.
  • 2-
    Adapter modules: Given an $M_{S}$ model that has been pre-trained and output $W_{S}$, for a target task $\mathbb{T}_{T}$. The adapter module aims to lunch a different set of parameters $K$ that is too much less than $W_{S}$, i.e, $K\ll W_{S}$. $K$ and $W_{S}$ must have the ability to be decomposed into more compact modules such that, $W_{S}=\{w\}_{n}$ and $K=\{k\}_{n}$. The adapter module permit learning the following new function $\mathbb{F}_{T}$:
    $\mathbb{F}_{T}(K,W_{S})=k_{1}^{\prime}\circ w_{1}\circ\dots k_{n}^{\prime}\circ w_{n}.$ (1)
    According to equation 1, during the adaptation procedure, the set of original weights $W_{S}=\{w\}_{n}$ is left unaltered, but the set of weights K is changed to $K^{\prime}=\{k^{\prime}\}_{n}$. The principle of the adaptation domain is illustrated in Fig. 3 (b).
  • 3-
    Feature based: Interested only in learning concepts and representations, at various levels, such as, word, character, phrase, or paragraph embedding $E$. The collection of $E$ based on a model $M$ remains unaltered, i.e., $\mathbb{F}_{T}(W_{S},E)=E\circ W^{\prime}$, in the way that $W^{\prime}$ is fine-tuned. For example, researchers have applied the generative adversarial network (GAN) principle to DTL where, the generators send features from the SD and the TD to a discriminator, which determines the source of the features and feeds the result back to the generators until they can no longer be distinguished. In this procedure, GAN obtains the common properties of two domains, as shown in Fig. 3 (c).
  • 4-
    Zero-shot: Is the easiest method among all of the others. Making the assumption that the parameters $W_{S}$ can’t be modified or add $K$ as a new parameters to a pre-trained model $M_{S}$ using $W_{S}$. To put this into context, in zero-shot there is no training technique to optimize or learn new parameters.
Figure 3: Structures of: (a) Fine-tuning, (b) DA, and (c) DTL-based GAN.
2.1.1 Transductive DTL
Compared to the traditional ML, which can be considered as a reference for DTL comparison, and given that the TDs $\mathbb{D}_{T}$ are distinct from the SDs $\mathbb{D}_{S}$. The SD has a labeled dataset ($X_{S}$ labeled with $Y_{S}$), whereas the TD has no labeled dataset, the source and target tasks are equal ( Table 2). The goal of transductive DTL is to build the target prediction function $\mathbb{F}_{T}$ in the $\mathbb{D}_{T}$ by knowledge of the $\mathbb{D}_{S}$ and $\mathbb{T}_{T}$. Furthermore, the transductive DTL environment may be further classified into two categories depending on different conditions between the source and destination domains:
(a) Domain adaptation (DA): The feature spaces across domains, $\chi_{S}$ and $\chi_{T}$, are the identical, but the marginal probability distributions of the input dataset are not, $P(Y_{S}/X_{S})\neq P(Y_{T}/X_{T})$. For example, an assessment may be done on the topic of the resort in the $\mathbb{D}_{S}$ and it will be used to train a model for restaurants in the $\mathbb{D}_{T}$. DA is mostly effective when the $\mathbb{T}_{T}$ has a distinct distribution or there is a scarcity of labeled data.
(b) Cross-modality DTL: Also known as cross-lingual DTL in the spoken language field, most DTL methods, more or less, a connection in feature spaces or label spaces is required between $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$. In other words, DTL can only occur when the source and destination data are both in the same modality, like video, speech, or text. cross-lingual DTL, in contrast to all other DTL approaches, is one of the most complicated issues in DTL. It is assumed that the feature spaces of the source and destination domains are completely distinct ($\chi_{S}\neq\chi_{T}$), as in speech-to-image, image-to-text, and text-to-speech. Furthermore, the label spaces of source $Y_{S}$ and destination $Y_{S}$ domains might differ ($Y_{S}\neq Y_{T}$).
(c) Unsupervised DTL: Intends to enhance the learning of the target predictive function $\mathbb{F}_{T}$ in $\mathbb{D}_{T}$ using the knowledge in $\mathbb{D}_{S}$ and $\mathbb{T}_{S}$, where $\mathbb{T}_{S}$ different from $\mathbb{T}_{T}$ but related, and $Y_{S}$ and $Y_{T}$ are not visible, given a SD $\mathbb{D}_{S}$ with a learning task $\mathbb{T}_{S}$, a TD $\mathbb{D}_{T}$ and a matching learning task $\mathbb{T}_{T}$ ($\mathbb{D}_{S}$ different from $\mathbb{D}_{T}$, but related).
2.1.2 Adversarial DTL
In contrast to the methods described above for DTL, adversarial learning [wang2020transfer] aids in the learning of more transferable and discriminative representations. The work in [ganin2016domain], was the first that introduced the domain-adversarial neural network (DANN). Instead of using a predefined distance function like maximum mean discrepancy (MMD), the core idea is to use a domain-adversarial loss in the network. This has greatly aided the network’s ability to learn more discriminative data. Many studies have used domain-adversarial training as a result of DANN’s idea [bousmalis2016domain, chen2019joint, long2017deep, zhang2018collaborative]. All of the previous work ignores the different effects of marginal and conditional distributions in adversarial TL, whereas in [wang2020transfer], the proposed scheme, named dynamic distribution alignment (DDA), can dynamically evaluate the importance of each distribution.
3 Background
3.1 Acoustic and language models
The AM is in charge of capturing the sound characteristics of different phonetic units. This involves generating statistical measures for characteristic vector sequences from the audio waveform. Various techniques, such as linear predictive coding (LPC), Cepstral analysis, filter-bank analysis, Mel-frequency cepstral coefficients, wavelet analysis, and others, can be used to extract these features [37]. In the processing stage, a decoder (search algorithm) uses the acoustic lexicon and LM to create the hypothesized word or phoneme. You can see the overall process illustrated in Figure 4.
[…]
LM provide probabilities of sequences of words, crucial for ASR systems to predict the likelihood of subsequent words in a sentence [38]. A domain-specific LM is trained on text data from the target domain to capture its unique vocabulary and grammatical structures [39]. For n-gram models, this involves calculating the conditional probability of a word given the previous $n-1$ words [40]:
$P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}$ (2)
In the context of ASR, the LM complements the AM by providing linguistic context. The combined probability from the AM and the LM helps in determining the most likely transcription for a given audio input during the decoding process. The frequently utilized LM in ASR systems is the backoff n-gram model.
Figure 4: Diagram illustrating the end-to-end framework for ASR.
3.2 Evaluation criteria in ASR
To assess the effectiveness and suitability of ASR techniques, researchers have employed diverse methods. Some of these encompass well-established DL metrics, including accuracy, F1-score, recall (sensitivity or true positive rate (TPR)), precision (also known as positive predictive value), and specificity (commonly referred to as true negative rate (TNR)) [41]. These metrics serve as crucial evaluation criteria for experimental outcomes, as evidenced in studies such as [42, 43, 44]. Additionally, there are ASR-specific metrics, which are detailed in Table LABEL:tab:metrics.
Metric
Formula
Description
WER
$\mathrm{\frac{S+D+I}{N}=\frac{S+D+I}{H+S+D}}.$
The word error rate (WER) serves as a frequently utilized metric to assess the performance of Automatic Speech Recognition (ASR). It is computed by determining the ratio of incorrectly recognized words to the overall number of processed words [45, 17, 46]. In the given context, $\mathrm{I,D,S,H}$, and $\mathrm{N}$ denote the quantities of insertions, deletions, substitutions, hits, and input words, respectively. Instead of WER, the character error rate (CER) has been employed, while adhering to the same evaluation principle.
PESQ and MOS-LQO
$\mathrm{MOS-LQO}=0.999+\frac{4.999-0.999}{1+e^{-1.4945.PESQ+4.6607}}$
PESQ serves as an objective technique for evaluating the perceived quality of speech [47]. The assessment involves assigning numerical scores within the range of -0.5 to 4.5. Additionally, a correlation can be established between MOS and PESQ scores, giving rise to a novel evaluation metric termed the mean opinion score-listening quality objective (MOS-LQO), also identified as PESQ Rec.862.1. [22]
RTF
RTF=Total Processing TimeTotal $\mathrm{RTF=\frac{\text{Total Processing Time}}{\text{Total Duration}}}$
[…]
RTF serves as a standard metric to assess the processing time cost of an ASR system. It represents the average processing time required for one second of speech
RER
$\frac{(E_{\text{baseline}}-E_{\text{proposed}})}{E_{\text{baseline}}}\times 100\%$
The relative error rate (RER) expresses the percentage error rate achieved by the proposed DL model compared to the baseline. $E_{\text{baseline}}$ is the error rate of the baseline model. $E_{\text{proposed}}$ is the error rate of the proposed model or method.
D
$\frac{\sum_{i=1}^{n}\Big{(}1-\frac{\sum_{j=1}^{n}a_{ij}\cdot|i-j|}{\max(|i-1|,|i-2|,\ldots,|i-n|)}\Big{)}}{n}\hskip 28.45274pt$
Diagonal centrality of an attention matrix (D) is defined as the mean value across the centrality of all its rows. where $j$ represents the index of each column, $n$ signifies the length of the input sequence, $a_{ij}$ denotes the attention weight between the $i$-th and $j$-th elements of the input sequence, and $|i-j|$ signifies the distance between the $i$-th and $j$-th elements of the input sequence [48].
3.3 ASR datasets
Various datasets have been employed in the literature for diverse ASR tasks. Table 4 presents a selection of datasets utilized for DTL-based ASR applications, along with their respective characteristics. It is important to note that the table primarily includes publicly accessible repositories. Furthermore, it is worth mentioning that certain datasets have undergone multiple updates and improvements over time, leading to their enhanced development.
Table 4: List of publicly available datasets used for advanced DL-based ASR applications
. Dataset Used by Default ASR task Characteristics LibriSpeech [49, 50, 51] Train and assess systems for recognizing speech. The collection consists of 1000 hours of speech recorded at a 16 kHz sampling rate, sourced from audiobooks included in the LibriVox project. DCASE [52, 53] Identifying acoustic environments and detecting sound occurrences. Comprise 8 coarse-level and 23 fine-level urban sound categories, collected in New York City in 2020 using 50 acoustic sensors. WSJ [48] Acoustic scene and sound event corpus Comprises an extensive 81 hours of meticulously curated read speech training data. SWBD [48] Conversational telephone speech corpus Is a comprehensive collection, boasting a substantial 260 hours of conversational telephone speech training data. AISHELL [54, 55, 56] Chinese Mandarin speech corpus 400 participants from diverse Chinese accent regions recorded in a quiet indoor space using high-fidelity microphones, later downsampled to 16kHz. CHIME3 [57] SR for distant microphone in real-world settings. Includes around 342 hours of English speech with noisy transcripts and 50 hours of noisy environment recordings. Google-SC [57] Speech commands with a restricted range of words. The dataset contains 105,829 one-second utterances of 35 words categorized by frequency. Each utterance is stored as a one-second WAVE format file with 16-bit single-channel at 16KHz rate. It involves 2,618 speakers AURORA4 [57] Compare front-ends for large vocabulary recognition performance. Aurora-4 is a speech recognition dataset derived from the WSJ corpus, offering four conditions (Clean, channel, noisy, channel+noisy) with two microphone types and six noise types, totaling 4,620 utterances per set. Car-env [57] Vehicle environment sound Is a dataset from Korea that spans 100 hours of recordings in a vehicle. It comprises brief commands, with an average of 1.6 words per command. HKUST [50, 56] Classify Mandarin speech into standard and accented types. Comprises roughly 149 hours of telephone conversations in Mandarin. AudioSet [51] Audio event recognition Includes 1,789,621 segments of 10 seconds each (equivalent to 4,971 hours). It consists of at least 100 instances clustered into 632 audio classes, with only 485 audio event categories clearly identified. AWIC-19 [58] Arabic words recognition It comprises 770 recordings featuring isolated Arabic words. TED2 [59] English corpus for ASR The dataset was first made available in May 2012, for training, it comprising 118 hours, 4 minutes, and 48 seconds of training data from 666 speakers, containing approximately 1.7 million words.
4 Advanced ASR methods and applications
Traditional statistical LMs, such as backoff n-gram LMs, have been widely used due to their simplicity and reliability. However, bidirectional encoder representations from transformers (BERT), which utilize attention models, have shown better contextual understanding compared to single-direction LMs, as demonstrated in the work of Devlin et al. [60].
In terms of AM, deep learning-based models like the deep neural network-hidden Markov model (DNN-HMM) and the connectionist temporal classification (CTC) have made significant advancements. DNN-HMM models have been extensively studied in ASR research, while CTC is an end-to-end training method that does not require pre-alignment and only needs input and output sequences. The S2S model has also been successful in solving ASR tasks without using an LM or pronunciation dictionary, as described in Chiu et al. [61].
ASR systems often face performance degradation in certain situations due to the "one-model-fits-all" approach. Additionally, the lack of diverse and sufficient training data affects AM performance. To overcome these constraints and improve the resilience and flexibility of ASR systems, advanced DL methodologies such as DTL and it sub-field domain adaptation (DA), DRL, and FL have surfaced. These innovative methodologies collectively address issues concerning knowledge transfer, model generalization, and training effectiveness, offering remedies that expand upon the capabilities of traditional DL models within the ASR sphere. Thus, many research studies have focused on enhancing existing ASR systems by applying the aforementioned algorithms. Figure 5 provides an overview of the current SOTA advanced DL-based ASR and its most useful related schemes in both AM and LM.
Figure 5: Overview of advanced DL-driven ASR algorithms and their commonly utilized models.
4.1 Transformer-based ASR
The Transformer stands as a prominent deep learning model extensively employed across diverse domains, including NLP, computer vision (CV), and speech processing. Originally conceived for machine translation as a S2S model, it has evolved to find applications in various fields. The Transformer heavily relies on the self-attention mechanism, enabling it to capture extensive dependencies in input sequences. The standard Transformer model incorporates the query–key–value (QKV) attention mechanism. In this setup, given matrix representations of queries $\mathbf{Q}\in\mathbb{R}^{N\times D_{k}}$, keys $\mathbf{K}\in\mathbb{R}^{M\times D_{k}}$, and values $\mathbf{V}\in\mathbb{R}^{M\times D_{v}}$, the scaled dot-product attention is defined as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{k}}}\right)\mathbf{V}$
Here, $N$ and $M$ represent the lengths of queries and keys (or values), and $D_{k}$ and $D_{v}$ denote the dimensions of keys (or queries) and values. The softmax operation is applied row-wise to the matrix $\mathbf{A}$. Within the Transformer architecture, three attention mechanisms exist based on the source of queries and key–value pairs:
  • Self-attention: In the Transformer encoder, queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$ all equal the outputs of the previous layer, denoted as $\mathbf{X}$ in Eq. (2).
  • Masked Self-attention: In the Transformer decoder, self-attention is constrained, allowing queries at each position to attend only to key–value pairs up to and including that position. This is accomplished by implementing a mask function, before normalization, on the attention matrix $\hat{\mathbf{A}}=\exp\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{k}}}\right)$, where illegal positions are masked out by setting $\hat{A}_{ij}=-\infty$ if $i<j$. This type of self-attention is often referred to as autoregressive or causal attention.
  • Cross-attention: In cross-attention, queries originate from the results of the preceding (decoder) layer, while keys and values stem from the outputs of the encoder.
Numerous studies in the ASR field have introduced transformer-based approaches, encompassing both the acoustic and language domains. In the subsequent subsections, we delve into a comprehensive review and detailed analysis of several cutting-edge techniques within each of these categories.
Table 5: Summary of some proposed work in transformer-based ASR. The symbol ($\uparrow$) denotes result increase, whereas ($\downarrow$) signifies result decrease. In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
Ref.
Based on
Speech recognition task
Transformer
AM/LM
Result with metric
[55]
CNN
Solve the problem of code-switching
Multi-head attention
LM
RER= 10.2%
[56]
VGGnet
Compress ASR parameters and speeds up the inference time
Low-rank multi-head attention
AM
CER= 13.09%
[57]
DNN-HMM
Improve ASR
Attention
AM
RER= 4.7$\downarrow$
[62]
Emformer
Large scale ASR
Attention
AM
RERR= 26%
[63]
TRUNet
Sound source separation
TNet
AM
PESQ= 0.22$\uparrow$
[64]
MHSA
Improve speech/ASR
[…]
AM
PESQ= 0.96$\uparrow$
[58]
HMM
Improve ASR
Acoustic Encoder
AM
Acc= 96%
[65]
RNN-T
Acoustic re-scoring scenario
Transformer- Transducer
AM
Acc= 8$\uparrow$
[66]
CTC
ASR, ST, Acoustic ED
All-in-one
AM
WER=0.3$\uparrow$
[67]
CNN
Speech recognition with low latency, reduced frame rate, and streamability.
Transformer-Transducer
AM
WER= 3.6%
[68]
CTC alignment
Retrieve the acoustic embedding at the token level for better ASR
Attention
AM
[…]
WER= 2.3%
[69]
RNN-LSTM
Improve the efficiency of end-to-end ASR
Attention
LM
CER=1.98%
[59]
CTC Alignment
Enhance the performance of end-to-end ASR
Autoregressive Transformer
AM
RTF= 0.0134
WER= 2.7%
4.1.1 Acoustic domain
The study [57] reveals the Transformer model’s increased susceptibility to input sparsity compared to the convolutional neural network (CNN). The authors analyze the performance decline, attributing it to the Transformer’s structural characteristics. Additionally, they introduce a novel regularization method to enhance the Transformer’s resilience to input sparsity. This method directly regulates attention weights through silence label information in forced-alignment, offering the advantage of not requiring extra module training and excessive computation.
The paper [50] addresses a limitation in Transformer-based end-to-end modeling for ASR tasks, where intermediate features from multiple input streams may lack diversity. The proposed solution introduces a multi-level acoustic feature extraction framework, incorporating shallow and deep streams to capture both traditional features for classification and speaker-invariant deep features for diversity. A feature correlation-based fusion (FCF) strategy, employed to combine intermediate features across both the frequency and time domains, correlates and combines these features before feeding them into the Transformer encoder-decoder module.
The proposed masked autoencoding audio spectrogram Transformer (MAE-AST) operates solely on unmasked tokens [51], utilizing a large encoder. It concatenates mask tokens with encoder output embeddings, feeding them into a shallow decoder. Fine-tuning for downstream tasks involves using only the encoder, eliminating the decoder’s reconstruction layers. MAE-AST represents a significant improvement over the self-supervised audio spectrogram transformer (SSAST) model for speech and audio classification. Addressing the high masking ratio issue, the method achieves a 3$\times$ speedup and 2$\times$ memory usage reduction. During downstream tasks, the approach consistently outperforms SSAST. To identify varities of sounds types, Bai et al. [52] introduce SE-Trans, a cross-task model for environmental sound recognition, encompassing acoustic scene classification, urban sound tagging, and anomalous sound detection. Utilizing attention mechanisms and Transformer encoder modules, SE-Trans learns channel-wise relationships and temporal dependencies in acoustic features. The model incorporates FMix data augmentation, involving the creation of a binary mask from a randomly sampled complex matrix with a low-pass filter. SE-Trans achieves outstanding performance in ESR tasks, proven through evaluations on DCASE challenge databases, underscoring its robustness and versatility in environmental sound recognition.
Automated audio captioning (AAC) involves generating textual descriptions for audio recordings, covering sound events, acoustic scenes, and event relationships. Current AAC systems typically employ an encoder-decoder architecture, with the decoder crafting captions based on extracted audio features. Chen et al. in their paper [53] introduces a novel approach that enhances caption generation by leveraging multi-level information extracted from the audio clip. The proposed method consists of a CNN encoder with multi-level feature extraction (channel attention, spatial attention), A module specialized in predicting keywords to generate guidance information at the word level and Transformer decoder. Figure 6 depicts the overall architecture incorporating the three mentioned modules. Results demonstrate significant improvements in various metrics, achieving SOTA performance during the cross-entropy training stage.
Figure 6: An example of CNN-based transformer for automated audio captioning [53].
Adversarial audio involves manipulating sound to deceive or compromise machine learning (ML) systems, exploiting vulnerabilities in audio recognition models. Both [70, 71] work are built to combat adversarial noise using transformers. The authors in [70] employed a vision transformer customized for audio signals to identify speech regions amidst challenging acoustic conditions. To enhance adaptability, they incorporated an augmentation module as an additional head in the transformer, integrating low-pass and band-pass filters. Experimental results reveal that the augmented vision architecture achieves an F1-score of up to 85.2% when using a low-pass filter, surpassing the baseline vision transformer, which attains an F1-score of up to 81.2%, in speech detection. However, in [71] the authors present an adversarial detection framework using an attention-based transformer mechanism to identify adversarial audio. Spectrogram features are segmented and integrated with positional information before input into the transformer encoder, achieving 96.5% accuracy under diverse conditions such as noisy environments, black-box attacks, and white-box attacks.
The paper [72] introduces a parallel-path transformer model to address computation cost challenges for speech separation tasks. Using improved feed-forward networks and transformer modules, it employs a parallel processing strategy with intra-chunk and inter-chunk transformers. This enables parallel local and global modeling of speech signals, enhancing overall system performance by capturing short and long-term dependencies.
A Hybrid ASR approach outlines the conceptualization and execution of a technique that integrates neural network methodologies into advanced continuous speech recognition systems. This integration is built upon hidden Markov modelss with the aim of enhancing their overall performance. Wang et al. [73] introduce and assess transformer-based AMs for hybrid speech recognition. The approach incorporates various positional embedding methods and an iterated loss for training deep transformers. Demonstrating superior performance on the Librispeech benchmark, the suggested transformer-based AM outperforms the best hybrid result by 19% to 26% relative with a standard n-gram LM.
[…]
TRUNet proposed in [63] represents an innovative approach to end-to-end multi-channel reverberant sound source separation. The model incorporates a recurrent-U network that directly estimates multi-channel filters from input spectra, enabling the exploitation of spatial and spectro-temporal diversity in sound source separation. In Figure 7, the block diagram illustrates the architecture of TRUNet’s transformer. The transformer network (TNet) encompasses three variations: (i) TNet–Cat, which concatenates multi-channel spectra, treating them as a single input. This approach allows for the direct utilization of spatial information between channels. (ii) TNet–RealImag, utilizing two separate transformer stacks for real and imaginary parts, respectively. Queries and keys are computed from the multi-channel spectra. Despite this, the method may not fully exploit spatial information, such as phase differences, directly. (iii) TNet–MagPhase, analogous to TNet–RealImag, but employing spectral magnitude and spectral phase instead of real and imaginary parts. This variation proves superior in extracting spatial information from complex-valued spectra, resulting in maximum enhanced performance in sound source separation when employing TRUNet-MagPhase architecture.
Figure 7: An example of source separation scheme based on Transformer [63].
[…]
The study proposed by gong et al. [75] explores self-attention-based neural networks like the audio spectrogram transformer (AST) for audio tasks. It introduces a self-supervised framework, improving AST performance by 60.9% on various speech classification tasks such as ASR, speaker recognition, and more, reducing reliance on labeled data. The approach marks a pioneering effort in audio self-supervised learning. Likewise, a novel augmented memory self-attention addresses limitations of transformer-based acoustic modeling in streaming applications has been proposed in [76], outperforming existing streamable methods by over 15% in relative error reduction on benchmark datasets.
Shareef et al. in [58] propose a collaborative training method for acoustic encoders in Arabic ASR systems for speech-impaired children, achieving a 10% relative accuracy improvement on phoneme alignment in the output sequence. Pioneering in recognizing impaired children’s Arabic speech. Similarly in [77], collaboratively training acoustic encoders of various sizes for on-device ASR improves efficiency and reduces redundancy. Using co-distillation with an auxiliary task, collaborative training achieves up to 11% relative WER improvement on LibriSpeech corpus.
Transducer models, in the context of ASR, map input sequences (acoustic features) to output sequences (transcriptions). Unlike traditional S2S models, transducers can handle variable-length input and output sequences more efficiently. The study [65] explores attention masking in transformer-transducer-based ASR, comparing fixed masking with chunked masking in terms of accuracy and latency. The authors claim that variable masking is the viable choice in acoustic rescoring scenarios. Similarly, to adapt the Transformer for streaming ASR, the authors in [67] employ the Transducer framework for streamable alignments. Using a unidirectional Transformer with interleaved convolution layers for audio encoding, they model future context and gradually downsample input to reduce computation cost, while limiting history context length.
Moving on, The work [66] introduces an all-in-one AM based on the Transformer architecture, combined with the CTC to ensure a sequential arrangement and utilize timing details. It addresses ASR, audio tagging (AT), and acoustic ED simultaneously. The model demonstrates superior performance, showcasing its suitability for comprehensive acoustic scene transcription.
Winata et al. [56] propose a memory-efficient Transformer architecture for end-to-end speech recognition. It significantly reduces parameters, boosting training speed by over 50% and inference time by 1.35$\times$ compared to baseline. Experiments show better generalization, lower error rates, and outperformance existing schemes without external language or acoustic models. Growing demand for on-device ASR systems prompts interest in model compression.compression.
4.1.2 Language domain
Self-attention models, such as Transformers, excel in speech recognition and reveal an important pattern. As upper self-attention layers are replaced with feed-forward layers, resembling CLDNN architecture in [48], experiments on wall street journal (WSJ) and switchboard (SWBD) datasets show no performance drop and minor gains. The novel proposed metric of attention matrix diagonality indicates increased diagonality in lower to upper encoder self-attention layers. The authors conclude that a global view appears unnecessary for training upper encoder layers in speech recognition Transformers when lower layers capture sufficient contextual information. The study conducted by Hrinchuk et al. [49] presents a proficient postprocessing model for ASR with a Transformer-based encoder-decoder architecture, initialized with the weights of pre-trained BERT model [36]. The model effectively refines ASR output, demonstrating substantial performance gains through strategies like extensive data augmentation and pretrained weight initialization. On the LibriSpeech benchmark dataset, significant reductions in WERs are observed, particularly on noisier evaluation dataset portions, outperforming baseline models and approaching the performance of Transformer-XL neural language model re-scoring with 6-gram.
Figure 8: Some end-to-end models: Basic CTC , RNN-Transducer, and attention architectures [78].
[…]
CTC is an architecture and principle commonly used in S2S tasks (Figure 8), such as ASR. It enables alignment-free training by introducing a blank symbol and allowing variable-length alignments between input and output sequences. During training, the model learns to align the input sequence with the target sequence, and the blank symbol accounts for multiple possible alignments. CTC is particularly effective in tasks with variable-length outputs, making it well-suited for applications like speech recognition where the duration of spoken words may vary. Figure Figure 8 (a) illustrates the operational principles and components of CTC. This latter has been used in many ASR schemes, for example, Deng et al. [54] presents the innovative pretrained Transformer S2S ASR architecture, which integrates self-supervised pretraining techniques for comprehensive end-to-end ASR. Employing a hybrid CTC/attention model, it maximizes the potential of pretrained AM and LM. The inclusion of a CTC branch aids in the encoder’s convergence during training and considers all potential time boundaries in beam searching. The encoder is initiated with wav2vec2.0, and the introduction of a one-cross decoder (OCD) mitigates reliance on acoustic representations, enabling initialization with pretrained DistilGPT2 and overcoming the constraint of conditioning on acoustic features.
[…]
CS takes place when a speaker switches between words of two or more languages within a single sentence or across sentences. Zhou et al. [55] introduces a multi-encoder-decoder Transformer, for code-switching (CS) problem. It employs language-specific encoders and attention mechanisms to enhance acoustic representations, pre-trained on monolingual data to address limited CS training data. Hadwan et al. research [69] employ an attention-based encoder-decoder transformer, to enhance end-to-end ASR for the Arabic language, focusing on Qur’an recitation. The proposed model incorporates a multi-head attention mechanism and Mel filter bank for feature extraction. For constructing a LM, recurrent neural network (RNN) and long short term memory (LSTM) techniques were employed to train an n-gram word-based LM. The study introduces a new dataset, yielding SOTA results with a low character error rate.
4.2 TL-based ASR
Overall, DTL consists of training a DL model on a specific domain (or task) and then transferring the acquired knowledge to a new, similar domain (or task). In what follows, we present some of the definitions that are essential to understand the principle of DTL for ASR applications.
[…]
Figure 9: Deep transfer learning principle.
4.2.1 Acoustic domain
Schneider et al. [81] explored unsupervised pre-training for speech recognition using wav2vec model on large unlabeled audio data. The learned representations enhanced AM training with a simple CNN optimized through noise contrastive binary classification. In [82], a source filter warping data augmentation strategy is proposed to enhance the robustness of children’s speech ASR. The authors constructed an end-to-end acoustic model using the XLS-R wav2vec 2.0 model, pre-trained in a self-supervised manner on extensive cross-lingual corpora of adult speech. The work proposed in [83] introduces a multi-dialect acoustic model employing soft-parameter-sharing multi-task learning, a transductive DTL subcategory, within the Transformer architecture. Auxiliary cross-attentions aid dialect ID recognition, providing dialect information. Adaptive cross-entropy loss automatically balances multi-task learning. Experimental results demonstrate a significant reduction in error rates compared to various single- and multi-task models on multi-dialect speech recognition and dialect ID recognition tasks. Similarly, in the realm of computer vision, CNNs models like ConvNeXt have outperformed cutting-edge transformers, partly due to the integration of depthwise separable convolutions (DSC). DSC, which approximates regular convolutions, enhances the efficiency of CNNs in terms of time and memory usage without compromising accuracy—in some cases, even enhancing it. The study [84] introduces DSC into the pre-trained audio model family for audio classification on AudioSet (target task), demonstrating its advantages in balancing accuracy and model size. Xin et al. [85] introduce an audio pyramid Transformer with an attention tree structure, with four branches, to reduce computational complexity in fine-grained audio spectrogram processing. It proposes a DA transfer learning approach for weakly supervised sound event detection, a sub-field of ASR, enhancing localization performance by aligning feature distributions between frame and clip domains with a DA detection loss.
4.2.2 Language domain
The methodology is founded on the utilization of the BERT model [86], which involves pretraining language models and demonstrates improved performance across various downstream tasks. DTL approaches for language models, specifically employed in the domain of voice recognition, are referred to as LM adaptation. These approaches aim to bridge the gap between the source distribution $\mathbb{D}_{S}$ and the target distribution $\mathbb{D}_{T}$. Song et al. [87] present a novel learning-to-rescore (L2RS) approach, which relies on two main components: (i) utilizing diverse textual data from SOTA NLP models, such as BERT, and (ii) automatically determining their weights to rescore the N-best lists for ASR systems.
Recent advancements in S2S models have shown promising results for training monolingual ASR systems. The CTC and encoder-decoder models are two popular architectures for end-to-end ASR. Additionally, joint training of these architectures in a multi-task hybrid approach has been explored, demonstrating improved overall performance. For instance, the architecture illustrated in Fig. 8 (a) comprises S2S layers. The encoder network of S2S consists of a series of RNNs that generate embedding vectors, while the RNN decoder utilizes these vectors to produce final results. The RNN also benefits from prior predictions ($P_{i},i=0,\dots,n$), enhancing the accuracy of subsequent predictions. Moving on, a novel TL-based approach that enhances end-to-end speech recognition has been proposed in [88]. The novelty lies in applying DTL through multilingual training and multi-task learning at two levels. The initial stage utilizes nonnegative matrix factorization (NMF), instead of a bottleneck layer, and multilingual training for high-level feature extraction. The subsequent stage employs joint CTC-attention models on these features, where the CTC was transferred to the target attention-based model. The scheme demonstrated superior performance on TIMIT but requires testing on high-resource data. Further optimization is needed for standard end-to-end training. In addition, Integrating both AM and LM methodologies has the potential to enhance or construct an effective DTL-based ASR model, as demonstrated in [46, 89, 90, 91, 92, 93].
Table 6: In contemporary cutting-edge frameworks, diverse pre-trained models are utilized for distinct tasks within the field. These frameworks employ different DTL approaches and assess their efficacy using specific metrics. The symbol ($\uparrow$) result increase, whereas ($\downarrow$) signifies result decrease. In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
Scheme
Based on
Speech recognition task ($\mathbb{T}_{T}$)
AM/LM
Adaptation
Result with metric
[84]
ConvNeXt-Tiny
Audio classification
AM
DA
mAP= 0.471
[85]
APT
Sound event detection
AM
DA
F1= 79.6%
[49]
BERT (Jasper)
Speech-to-text
LM
TL
WER= 14%
[54]
DistilGPT2
Improve ASR
Both
Fine-tuning
CER= 4.6%
[93]
XLRS Wave2vec
Improve ASR in low resource language
Both
Fine-tuning
5.6% WER $\downarrow$
[82]
XLRS Wave2vec
Improve ASR for children’s speech
AM
TL
WER = 4.86%
[90]
S2S
Speaker adaptation
Both
Features norm.
[…]
[83]
Transformer
Multi-dialect model aids recognizing diverse speech dialects effectively
AM
Multi-task learning
Acc= 100%
[94]
S2S
Enhancing the existing multilingual S2S model.
LM
DTL
4%CER $\downarrow$ 6% WER
[95]
PaSST
Audio tagging and event detection
AM
Fine-tuning
F1= 64.85%
[81]
Wav2vec
WSJ data speech
AM
affine transform.
[…]
[96]
ARoBERT
Spoken language understanding
LM
Fine-tuning
F1-score=92.56%
Abbreviations: Transformer (T)
4.3 FL-based ASR
[…]
FL revolutionizes AI model training by enabling collaboration without the need to share sensitive training data. Traditional centralized approaches are evolving towards decentralized models, where ML algorithms are trained collaboratively on edge devices like mobile phones, laptops, or private servers [97]. The mathematical formulation of FL focuses on training a single global model across multiple devices or nodes (clients) while keeping the data localized. The objective is to minimize a global loss function that is typically the weighted sum of the local loss functions on all clients. The standard FL problem can be formulated as:
$\min_{\theta}F(\theta)=\min_{\theta}\sum_{k=1}^{K}\frac{n_{k}}{N}F_{k}(\theta)$ (3)
In this context, $\theta$ denotes the parameters of the global model to be learned, $K$ represents the total number of clients, $n_{k}$ signifies the number of data samples at client $k$, $N=\sum_{k=1}^{K}n_{k}$ stands for the total number of data samples across all clients, and $F_{k}(\theta)$ indicates the local loss function computed on the data of client $k$. In FL, the goal is to find the global model parameters $\theta$ that minimize the global loss function $F(\theta)$, which is an aggregate of the local loss functions from all participating clients. This process typically involves iterative updates to the model parameters using algorithms like federated averaging (FedAvg), where clients compute gradients or updates based on their local data and send these updates to a central server. The server then aggregates these updates to improve the global model.
  • Horizontal federated fearning (HFL): In HFL, clients train a shared global model using their respective datasets, characterized by the same feature space but different sample spaces. Each client utilizes a local AI model, and their updates are aggregated by a central server without exposing raw data. The HFL training process involves: (1) initialization, (2) local training, (3) encryption of gradients, (4) secure aggregation, and (5) global model parameter updates. The objective function minimizes a global loss across all parties’ datasets [97].
  • Vertical federated learning (VFL): VFL trains models on datasets sharing the same sample space but having different feature spaces. Through entity data alignment (EDA) and encrypted model trained (EMT), VFL allows clients to cooperatively train models without sharing raw data. The training process involves the same steps as HFL [97].
Figure 10: Federated learning principle: (a) Horizontal FL, (b) Vertical FL.
FL presents a paradigm shift in AI training, promoting collaboration while respecting data privacy. The working principle of HFL, and VFL is depicted on Figure 10, they cater to various data distribution scenarios, offering flexible solutions for decentralized and secure ML. The application of these FL frameworks extends across diverse domains, promising improved model accuracy and privacy preservation. The first work introducing FL in ASR is presented in [98]. The authors introduced a FL platform that is easily generalizable, incorporating hierarchical optimization and a gradient selection algorithm to enhance training time and SR performance. Guliani et al. [99] proposed a strategy to compensate non-independent and identically distributed (non-IID) data in federated training of ASR systems. The proposed strategy involved random client data sampling, which resulted in a cost-quality trade-off. Zhu et al. [100] addressed also FL-based ASR in non-IID scenarios with personalized FL. They introduced two approaches: adapting personalization layer-based FL for ASR, involving local layers for personalized models, and proposing decoupled federated learning (DecoupleFL). DecoupleFL reduces computation on clients by shifting the computation burden to the server. Additionally, it communicates secure high-level features instead of model parameters, reducing communication costs, particularly for large models. In [101], the authors proposed a client-adaptive federated training scheme to mitigate data heterogeneity when training ASR models. Nguyen et al. [102] used FL to train an ASR model based on a wav2vec 2.0 model pre-trained by self supervision. Yang et al. [103] proposed a decentralized feature extraction approach in federated learning. This approach is built upon a quantum CNN (QCNN) composed of a quantum circuit encoder for feature extraction, and an RNN based end-to-end AM. This framework takes advantage of the quantum learning progress to secure models and to avoid privacy leakage attacks. Gao et al. [104] tackled a challenging and realistic ASR federated experimental setup with clients having heterogeneous data distributions, featuring thousands of different speakers, acoustic environments, and noises. Their empirical study focused on attention-based S2S End-to-End (E2E) ASR models, evaluating three aggregation weighting strategies: standard FedAvg, loss-based aggregation, and a novel WER-based aggregation. Table 7 summarizes the most recent FL-based ASR techniques.
Table 7: Summary of recent proposed work in FL-based ASR. All the schemes are suggested for AM. The symbol ($\uparrow$) result increase, whereas ($\downarrow$) signifies result decrease. In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
Scheme
Based on
Speech recognition task
FL technique
Metric and result
[98]
S2S
Improve ASR
FedAvg
WER = 6%
[99]
End-to-end RNN-T
ASR on non-IDD data
FedAvg
WER=
[100]
CNN+Transformer extractor
ASR on non-IDD data
DecoupleFL
2.3- 3.4% WER $\downarrow$ compared with FedAvg
[101]
LSTM
ASR on non-IDD data
CAFT
WER = 15.13%
[102]
wav2vec 2.0
Improve ASR
FedAvg
WER= 10.92%
EER= 5-20%
[103]
QCNN and RNN
Improve privacy-preservation in ASR
VFL
Accuracy = 95.12%
[104]
S2S
ASR on heterogeneous data distributions
FedAvg
WER= 19.98-23.86%
[105]
S2S
ASR on private and unlabelled user data.
FedNST
WER= 22.5%
[106]
Wav2vec 2.0 and Whisper
ASR & KWS for child exploitation settings
FedAvg
WER = 11-25%
[107]
Kaldi and backoff n-gram
Improve privacy-preservation in ASR
Merging models
WER= 17.7%
[108]
TDNN
Improve privacy-preservation in ASR
Aggregation
EER = 1-2%.
[109]
Non-Streaming & Streaming Conformer
Reduce client ASR model size
Federated Dropout
6-22% $\downarrow$ Client size model; 34-3% WER $\downarrow$
4.4 RL-based ASR
[…]
RL is a ML paradigm where an agent learns optimal decision-making by interacting with an environment. The agent receives feedback in the form of rewards or penalties, adapting its behavior to maximize cumulative reward over time through a trial-and-error process. reinforcement learning (RL) involves several key concepts, as defined in the following terms: Environment model, serving as a representation of contextual dynamics; State (s), denoting the current situation perceived by the agent; Observation (o), a subset of the state directly perceived by the agent; Action (a), the decision made by the agent in response to the environment; Policy ($\pi$), describing how the agent converts environmental conditions into actions; Agent, the entity making decisions based on current states and past experiences; Reward, numerical values assigned by the environment to the agent based on state-action interactions. Figure 11 illustrate the principle of RL.
Figure 11: RL principle.
Markov Decision Process (MDP) is a fundamental framework for dynamic and stochastic decision-making, characterized by state space $S$, action space $\mathbb{A}$, transition probabilities $\mathbb{P}$, and a reward function $R$. The primary objective in an MDP is to identify an optimal policy $\pi^{*}$ maximizing the expected discounted total reward over time, expressed as $\pi^{*}=\max_{\pi}\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}(s_{t},\pi(s_{t}))]$, where $\gamma$ is the discount factor. MDPs find extensive applications in addressing uncertainties in intelligent systems within dynamic wireless environments, including spectrum management, cognitive radios, and wireless security.
In the field of ASR, RL has primarily been proposed to tackle discrepancies between training and testing phases. Two main discrepancies leading to potential performance deterioration have been identified: 1) The conventional use of the cross-entropy criterion maximizes log-likelihood during training, while performance is assessed by WER, not log-likelihood; 2) The teacher-forcing method, which relies on ground truth during training, implies that the model has never encountered its own predictions before testing. RL addresses these discrepancies by bridging the gap between the training and testing phases. Several RL-based approaches for ASR have been proposed [110, 111, 112, 113, 114, 115, 116, 117, 118, 119]. For example, in [110], the authors introduced a RL-based optimization method for the S2S ASR task called Self-critical sequence training (SCST). This method can be conceptualized as a sequential decision model, depicted in Figure 12. The entire encoder-decoder neural network is treated as an agent. At each time step $t$, the current state $s_{t}$ is formed by concatenating the acoustic feature $x_{t}$ and the previous prediction $Y_{t-1}$. The output token serves as the action, updating the generated hypotheses sequence. SCST associates training loss and WER using a WER-related reward function, calculating the reward $r_{t}$ at each token generation step by comparing it with the ground truth sequence $Y^{*}$. SCST uses the test-time beam search algorithm to sample hypotheses for reward normalization, assigning positive weights to high-reward hypotheses that outperform the current test-time system and negative weights to low-reward hypotheses. The framework is illustrated in Figure 12.
In ASR, RL has been primarily proposed to address mismatches between training and testing phases. Two main mismatches leading to potential performance degradation are identified: 1) The commonly used cross-entropy criterion maximizes log-likelihood during training, while performance is evaluated by WER, not log-likelihood; 2) The teacher-forcing method, relying on ground truth during training, implies that the model has never encountered its own predictions before testing. RL bridges the gap between the training and testing phases, addressing these mismatches.
Several RL-based ASR approaches have proposed [110, 111, 112, 113, 114, 115, 116, 117, 118, 119]. For instance, In [110], the authors presented a RL-based optimization method for S2S ASR task called self-critical sequence training (SCST). This can be viewed as a sequential decision model as shown in Fig. 12. The whole encoder-decoder neural network can be viewed as an agent. In each time step $t$, acoustic feature $x_{t}$ and previous $Y_{t-1}$ prediction are concatenated as current state $s_{t}$. The output token is the action at that will update the generated hypotheses sequence. After comparing it with ground truth sequence $Y^{*}$, a reward $r_{t}$ of this time step is calculated. SCST associates the training loss and WER using WER-related reward function, which considers the intermediate reward at each token generation step. Furthermore, SCST utilizes the test-time beam search algorithm to sample a set of hypotheses for reward normalization. As a result, the high-reward hypotheses that outperform the current test-time system are given positive weights, while the low-reward hypotheses are given negative weights.
The entire encoder-decoder neural network is treated as an agent. At each time step $t$, the current state $s_{t}$ is formed by concatenating the acoustic feature $x_{t}$ and the previous prediction $Y_{t-1}$. The output token serves as the action, updating the generated hypotheses sequence. SCST associates training loss and WER using a WER-related reward function, calculating the reward $r_{t}$ at each token generation step by comparing it with the ground truth sequence $Y^{*}$. SCST uses the test-time beam search algorithm to sample hypotheses for reward normalization, assigning positive weights to high-reward hypotheses that outperform the current test-time system and negative weights to low-reward hypotheses. The framework is illustrated in Figure 12.
Figure 12: SCST sequential decision model of ASR [110].
In [111], the authors developed a RL framework for speech recognition systems using the policy gradient method. They introduced a RL method within this framework, incorporating user feedback through hypothesis selection. Tjandra et al. [112, 113] also employed policy gradient RL to train a S2S ASR model. In [114], the authors constructed a generic RL-based AutoML system. This system automatically optimizes per-layer compression ratios for a SOTA attention-based end-to-end ASR model, which consists of multiple LSTM layers. The compression method employed in this work is singular value decomposition (SVD) low-rank matrix factorization. The authors improved this approach by combining iterative compression with AutoML-based rank searching, achieving over 5 x ASR compression without degrading the WER [115]. Shen et al. [116] suggested employing RL to optimize a speech enhancement model based on recognition results, aiming to directly enhance ASR performance. AutoML-based low-rank factorization (LRF) achieves up to 3.7× speedup. In the shade of this, Mehrotra et al. in their work [115] propose an iterative AutoML-based LRF that employs RL for the iterative search, surpassing 5× compression without degrading WERs, advancing ASR. Table 8 summarizes the recent RL-based ASR techniques.
Table 8: Summary of recent proposed works in RL-based ASR. The symbol ($\uparrow$) result increase, whereas ($\downarrow$) signifies result decrease. In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
Scheme
Model-based
ASR Tasks
RL technique
Metric and result
[110]
S2S Conformer
Improve ASR
Policy gradient
8.7%-7.8% WER $\downarrow$ over Baseline model
[111]
DNN-HMM
Improve ASR for AM
Policy gradient
WER=23.82-25.43%
[112]
S2S
Improve ASR FOR am
Policy gradient
CER=6.10%
[113]
S2S
Improve ASR for AM
Policy gradient
CER=6.10%
[114]
End-to-end encoder-attention- decoder
Improve ASR Model compression for AM
Policy gradient
Up to $\sim$3x compression; WER=8.06%
[115]
End-to-end encoder-attention- decoder
Improve ASR Model compression for AM
Policy gradient
Up to $\sim$5x compression; WER=8.19%
[116]
CD-DNN-HMM AM & SRI LM
Speech enhancement for ASR for AM and LM
Q-learning
12.40% and 19.23% CER $\downarrow$ at 5 and 0 dB SNR conditions.
[117]
LSTM
Improve ASR for dialogue state tracking
DQN
Acc= 3.1$\uparrow$
[118]
S2S
Improve ASR for AM
Policy gradient
CER=8.7%
[119]
Wav2vec 2.0
Improve ASR for AM
Policy gradient
4% WER $\downarrow$
DTL for acoustic models (AMs) Two types of deep learning-based AMs commonly used in ASR are the end-to-end model and the layered deep neural network-hidden Markov model (DNN-HMM) model. The DNN component is responsible for extracting high-level features, such as MFCCs and HMM lexical sequences, from acoustic signals. These features are then decoded into transcripts. The DNN takes acoustic characteristics as input and produces context-dependent lexical units (tri-phonemes) that are used as input for the downstream HMM component.
In contrast, the end-to-end model is a DNN-based technique that directly takes acoustic characteristics (features) as input and outputs the recognition rate without the need for separate components like the DNN and HMM. A typical end-to-end framework for voice recognition is illustrated in Figure 13. The neural network in this framework generates embeddings from the input features, which are then passed through a stack of recurrent layers. These recurrent layers analyze patterns based on prior and current input information to produce the final output. The network is trained using back-propagation with the CTC loss [mridha2021study]. Regarding domain transfer learning (DTL), three main strategies have been commonly employed with AMs.
Figure 13: An example of end-to-end source model for DTL-based ASR [mridha2021study].
4.4.1 Feature normalisation based-DTL
The concept behind linear transformation is to normalize speech characteristics through a linear mapping process. This involves adding a transformation network or layer to an existing network. The transformation network is an effective method for adapting neural networks. Typically, the last hidden layer is designed to act as a bottleneck, reducing the number of parameters to be adjusted by using fewer neurons. Figure 14 illustrates this architecture. The transformation network can be either a linear input network or a linear output network.
Figure 14: The basics of transformation in DTL involve estimating the weights connected to the links within the dashed rectangles, while leaving the remaining weights unchanged. In the case of mono-task DTL, feature normalization is applied as shown in Figure (a). For multi-task DTL, feature normalization is also performed, as illustrated in Figure (b).
Assuming that the last hidden network (LHN) functions as a feature extractor and the output layer serves as a discriminative source model ($M_{S}$), the weights of the linear transformation matrix $W_{L}$ in the output layer correspond to the parameters of the target model, $M_{T}$. The representation of $M_{T}$ can be expressed as follows [huang2015maximum]:
$\mathrm{M_{T}=softmax(W_{L}M_{S})}.$ (4)
The activation of the last hidden layer in $M_{S}$ can serve as the extracted feature representation for the hidden layers in $M_{T}$. Transforming the model parameters using a transformation matrix $W_{LHN}$ to generate an adapted set of model parameters is equivalent to incorporating an extended last hidden layer after the existing last hidden layer [huang2015maximum]:
$\mathrm{M_{T}=softmax(W_{LHN}W_{L}M_{S})}.$ (5)
Several ASR research studies have utilized linear transformation strategies to improve performance. In one example [elaraby2016deep], the authors aimed to enhance a computer-aided language learner (CAPL) system for teaching Arabic pronunciation for Quran recitation regulations. They implemented various improvements, including speaker adaptive training (SAT), integrating a hybrid DNN-HMM model, combining the hybrid DNN with minimum phone error (MPE), and using a grammar-based decoding graph during testing. Another study [mimura2016joint] employed a multi-target learning approach to optimize the output of a denoising auto-encoder (DAE) and the input of a DNN. The DAE was trained in the first stage to reduce input-related errors propagated to the DNN. Then, a unified network of DAE and DNN was fine-tuned for phone state ASR, with an additional target of input voice augmentation applied to the DAE. Additionally, in [ma2017approaches], an adaptation layer was introduced for fine-tuning, and non-linearity was incorporated to learn a more complex function than a linear transformation in the softmax layer. During fine-tuning, adjustments were made to the cluster softmax and NNadapt layers while keeping the other layers unchanged. Table 6 provides a summary of additional schemes that employ linear transformation techniques, along with their performance details.
4.4.2 Conservative training for DTL
The conservative training approach is commonly used for accent adaptation in ASR. It is an efficient method that requires only a limited amount of spoken data to achieve satisfactory results. However, it can lead to an excessive number of parameters, which may disrupt the model’s structure. To address this, KL-Divergence (KLD) is widely employed as a DNN-based DTL algorithm for ASR. KLD regularization provides a mathematical framework for DL training, aiming to minimize the loss and make the output distributions of the source model $M_{S}$ and target model $M_{T}$ more similar. The KL-divergence prevents overfitting and helps to keep the adapted $M_{T}$ close to the domain of $M_{S}$ [89, 90]. In terms of model-based DTL with KLD-regularization, assuming the loss functions for training DNNs in the source and target domains are $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$, respectively, the conservative approach can be summarized mathematically as follows [89, 90]:
$\mathbb{D}_{T}^{KLD}=(1-\rho)\mathbb{D}_{S}+\frac{\rho}{N}\sum P_{S}(x_{T}/x_{T})logP_{T}(y_{T}/x_{T}),$ (6)
In the context of DTL-based ASR, the conservative training approach with KLD-regularization is utilized to build a target AM. This approach involves training DNNs on speech samples collected from the target domain $\mathbb{D}_{T}$, denoted as ($x_{T},y_{T}$), where $N$ represents the number of speech samples in $\mathbb{D}_{T}$. The hyper-parameter $\rho$ controls the transfer ratio from the source domain $\mathbb{D}_{S}$. An example of applying this DTL-based ASR approach with conservative training and KLD-regularization can be found in [89] for building a target AM. Additionally, in [90], both KLD and LHN techniques were employed for speaker adaptation using a pre-trained seq2seq ASR model. Table 6 summarizes other techniques that employ conservative training along with their corresponding performance details.
4.4.3 Subspace-based DTL
Subspace-based DTL approaches in ASR utilize unsupervised techniques such as PCA, SVD, and NMF for data dimensionality reduction. These techniques aim to identify a subspace of model parameters or transformations that capture essential information. PCA maps high-dimensional data to lower-dimensional subspaces while preserving correlation and maximizing variance. SVD, similar to PCA, condenses networks by selecting a specific number of singular values. However, SVD introduces nonlinearity in the weight matrix, which can lead to information loss when constructing a linear projection layer. NMF algorithms require at least one nonnegative matrix and express the target matrix as a weighted sum of base matrix columns, making it more interpretable than SVD and PCA. In the study [88], Convex Nonnegative Matrix Factorization (CNMF), a variant of NMF, is employed to extract high-level features. Subsequently, DTL is applied to both high and low-level features through multilingual training and multi-task learning. These techniques yield significant performance improvements compared to state-of-the-art ASR studies. Other studies utilizing subspace techniques and their respective performance details can be found in Table 6.
LDA-based DTL In the case of discrete data collection, generative probabilistic models, such as the latent Dirichlet allocation (LDA), are employed. Typically, LDA is a hierarchical Bayesian model with three levels, where each item in a collection is represented as a finite mixture over a set of underlying topics. Each topic, in turn, is modeled as an infinite mixture of topic probabilities. To capture the relationship between words and create language models for specific documents, topic model-based techniques, such as LDA, have been utilized as described in [song2019topic]. In the work presented in [hentschel2018feature], LDA features are transformed using a linear layer consisting of a weight matrix and a bias vector. These transformed features are then utilized as inputs in the LHN (Local Hidden Node) during the training and evaluation of the network.
4.4.4 NNLM-based DTL
Neural network language models (NN-LMs) generally outperform count-based LM models in automatic speech recognition (ASR) across various tasks. Specifically, when applied to N-best rescoring, NN-LMs achieve WER as mentioned in [hentschel2018feature]. Adapting NN-LMs to new domains poses a research challenge, and current approaches can be categorized as either model-based or feature-based adaptations. In feature-based adaptation, auxiliary features are incorporated into the input of an NN-LM, while model-based adaptation involves fine-tuning and adapting network layers. The authors of [91] propose a recurrent neural network-based LM (RNN-LM) approach where both types of adaptation are explored. As an illustrative example, Figures 15 (a) and (b) provide detailed explanations of the adopted RNN-based DTL in the study. Furthermore, the authors in [92] suggest a DNN-based model for LM modification in ASR, employing a factorized time-delay neural network (TDNN-F). Specifically, the TDNN-F model is trained using a combination of cross-entropy and lattice-free maximum mutual information objective functions (LF-MMI). The effectiveness of TDNN-F is demonstrated in the recognition of English child speech [92].
Figure 15: DTL-based RNNLM with different adaptation techniques, where the out-of-vocabulary (OOV) node represents an input word that does not belong to the specified vocabulary but can be included in the input. Similarly, out-of-shortlist (OOS) nodes can also be included in the output: (a) RNNLM with LHN adaptation layer. (b) RNNLM with feature-based adaptation layer.
LSTM-based DTL Generally, the NNLM models used in ASR are still trained on a sentence-level corpus, despite the attempts to train them at the document level. This is due to various factors; for example, a more extended context may not be relevant for enhancing next-word prediction in conventional ASR systems. It is also challenging to gather training data representing extended session contexts in many conversational circumstances. Long-span models are becoming more common in scenarios where they are beneficial. Long-span models will likely help scenarios, such as transcriptions of conversations and meetings, human-to-human communication, and document production by voice [parthasarathy2019long]. LSTM models are widely employed, and their architectures are well-suited to variable-length sequences. Therefore, they can exploit extreme long-range dependencies without using n-gram approximation. For instance, by employing equal context, the authors in [tuske2018investigation] demonstrate that the deep 4-gram LSTM outperforms big interpolated count models by performing considerably better backing off and smoothing. In another example, the central part of a shared encoder is constructed using BLSTM [94].
Cross-domain ASR
4.4.5 Cross-language DTL
Cross-language DTL is a particular application of cross-modality DTL. It serves as a commonly employed method for constructing ASR models for low resource languages by leveraging models trained on other languages. This approach is based on the assumption that phoneme features can be shared across languages. Additionally, a generic ASR model can be adapted to a specific narrow domain using DTL techniques. To address the issue of data sparsity, various knowledge transfer methods are explored in [liu2019investigation] with the assistance of high-resource languages. These methods include DTL and fine-tuning, where a well-trained neural network initializes the parameters of the LHN. Progressive neural networks (Prognets) are also examined, as they are resistant to the forgetting effect and excel at knowledge transfer due to the presence of lateral connections in the network architecture. Furthermore, the utilization of cross-lingual DNNs is explored, where bottleneck features are extracted to enhance the effectiveness of the ASR system. Recent approaches related to ASR using cross-language DTL and their respective performances are summarized in Tables 6 and 9.
Table 9: A summary of the recent ASR-based cross-language DTL technique. Whereas the marks ($\uparrow$) and ($\downarrow$) indicate improvement and reduction, respectively. If many scenarios has been conducted in one metric, only the best result is mentioned.
Scheme
Model-based
ASR Tasks ($\mathbb{T}_{T}$)
Characteristic
Performance
[yusuf2019low]
EDML
Framework of performing the DTL that reduces the impact of the prevalence of out-of vocabulary terms.
query-by-example task
[…]
[liu2019investigation]
Prognets
Improving ASR scheme quality by overcoming the data sparsity problems by means of high-resource languages.
Fine-tuning
LHN adapt.
[…]
[feng2019low]
FNN and CNN architectures
Indo-European speech samples used to improve the identification of African languages.
Using PLP coeff.
Fine-tuning
[…]
[sahraeian2018cross]
DNN
Weighted averaging schemes are used to combine the ensemble’s constituents, with the combination weights being trained to minimize the cross-entropy objective function.
Weights interpolation
[…]
[wilkinson2020semi]
CNN-GMM-HMM
Fully-automatic segmentation, semi-supervised training of ASR systems for five-lingual code-switched speech
Semi-supervised
[…]
4.4.6 DTL-based ASR for emotion recognition:
Combining language information with acoustic data has been shown to enhance the accuracy of speech emotion recognition (SER). Therefore, integrating both systems can be advantageous to enhance the capability of ASR systems to handle emotional speech and provide linguistic input to SER systems. Figure 16 illustrates a hybrid ASR-SER system [fayek2016deep], where a spectrogram is inputted into shared convolutional layers, followed by specialized layers that have shared levels to facilitate interaction between the two systems. This integration allows for improved performance in handling emotional speech and leveraging linguistic features.
Figure 16: A linguistic-paralinguistic hybrid system [fayek2016deep]. (a) ASR-SER Hybrid system. (b) An example of DTL-based ASR-SER hybrid system.
The approach presented in [tits2018asr] utilizes the internal representation of a speech-to-text system to explore the connection between valence/arousal and different modalities through the use of DTL. A speech-to-text or ASR system learns to map audio speech signals to their corresponding transcriptions. By employing DTL, the proposed method can estimate valence and arousal using features learned from an ASR task. This approach offers the advantage of combining large datasets of speech with transcriptions with smaller datasets annotated with emotional dimensions. In a similar vein, the work described in [ananthram2020multi] fine-tunes a TDNN-based speaker recognition model for the task of emotion detection using the crema-D multi-modal emotion dataset and canonical label clustering. By adapting the model using fine-tuning, the authors aim to improve its performance on emotion detection. The study presented in [boateng2020speech] focuses on extracting features from audio segments with extreme positive and negative ratings, as well as the ending of the audio. They employ the peak-end rule and a DTL approach to extract acoustic features. The authors utilize a pre-trained CNN speech model called YAMNet and a linear SVM for binary classification of partner valence.
4.4.7 Cross-corpus SER (CC-SER)
In the context of Speech Emotion Recognition (SER), it is typically assumed that speech utterances in the training and testing domains are recorded under the same conditions. However, in real-world scenarios, speech data is often collected from diverse environments or devices, resulting in a domain discrepancy that adversely affects recognition performance. As a solution, researchers have recently investigated the problem of cross-corpus SER (CC-SER) and explored various DTL models.
For example, [song2019transfer] proposes a transfer linear subspace learning (TLSL) scheme to develop a CC-SER framework. This approach facilitates the learning of shared feature space for both the source and target domains. The similarity between different corpora is estimated using a nearest-neighbor graph algorithm. Additionally, a feature grouping method is devised to partition emotional features into highly transferable parts (HTP) and low transferable parts (LTP).
In the case of unsupervised CC-SER examined in [liu2018unsupervised], only the training data is annotated. The authors introduce a domain-adaptive subspace learning (DoSL) technique to learn a projection matrix that transforms the source and target speech data from the initial domain to the labeled domain. This allows the classifier trained on the labeled source domain data to effectively predict the emotional states of the unlabeled target domain data. Furthermore, [liu2021transfer] presents an improvement to the DoSL-based CC-SER method by introducing transfer subspace learning (TRaSL). In another study by [luo2019cross], a semi-supervised CC-SER approach is proposed using non-negative matrix factorization (NMF). This approach incorporates training corpus labels into NMF and seeks a latent low-rank feature space where the differences in conditional and marginal distributions between the two corpora can be simultaneously minimized.
Advancing further, [zhang2019transfer] proposes a transfer sparse discriminant subspace learning (TSDSL) method to discover a shared feature subspace among multiple corpora. This is achieved by incorporating the $\ell_{2,1}$-norm penalty and discriminative learning, facilitating the identification of the most discriminative characteristics across the corpora. Similarly, [luo2020nonnegative] introduces a non-negative matrix factorization-based transfer subspace learning (NMFTSL) scheme. The goal is to minimize the distances between the marginal and conditional distributions in the common subspace. The distances are estimated using the maximum mean discrepancy (MMD) criterion. In [zhang2021cross], a joint transfer subspace learning and regression (JTSLR) technique is employed. It learns a latent subspace using discriminative MMD as the discrepancy metric, followed by modeling the relationships between features and annotations using a regression function in the latent subspace. A label graph is utilized to enhance knowledge transfer from the source domain (SD) data to the target domain (TD) data. Similarly, [chen2019target] presents a target-adapted subspace learning (TaSL) approach for CC-SER. It aims to find a projection subspace that enables more accurate regression of labels from the features. This effectively bridges the gap in feature distributions between the TD and SD. The projection matrix is optimized by combining $\ell_{1}$-norm and $\ell_{2,1}$-norm penalty terms with other regularization terms. Furthermore, [zhao2021cross] employs sparse subspace transfer learning (SSTL) to develop a CC-SER technique. It learns a robust common subspace projection using discriminative subspace learning and transfers knowledge from the source corpus to the target corpus through sparse reconstruction based on $\ell_{2,1}$-norm. The target samples are suitably represented as linear combinations of the SD data. On a different note, [braunschweiler2021study] investigates the impact of cross-corpus data complementation and data augmentation on the performance of SER models. The study focuses on six emotional speech corpora, considering factors such as single and multiple speakers, as well as variations in emotion style (natural, elicited, and acted).
4.5 Adversarial TL-based ASR
In many cases, the source model is trained using multilingual training, where abundant speech data is available in multiple languages [88, 94]. Multilingual training involves shared hidden layers (SHL) and language-specific layers or classifier layers for different languages. The SHL of the source model acts as a feature converter, transforming language-specific features into a common feature space [liu2019investigation]. However, there might be language-dependent features present in the common feature space, which hinders effective cross-lingual knowledge transfer. To address this issue, language-adversarial training is employed. Adversarial training helps in creating a language-invariant feature space. Once the source model is prepared, the first $n$ SHL can be transferred to the target model of an unknown language. In [yi2018language], the authors propose language-adversarial transfer learning as a solution to mitigate the performance degradation of the target model caused by shared features that may contain unnecessary language-dependent information. Fig. 17 illustrates the architecture of the suggested language-adversarial transfer learning method [yi2018language]. The source model, also known as the adversarial SHL model, is shown on the left, while the target model is depicted on the right. The presence of an additional language discriminator in the SHL model is denoted. The fully connected layer is represented as FC. The gradient reversal layer (GRL) ensures that the feature distributions across all languages are made as similar as possible for the language discriminator. The output labels of the language discriminator indicate the languages.
Figure 17: An example of proposed model architecture language-adversarial TL for limited ASR resource [yi2018language]. Senones refers to feature cluster’s name, representing similar acoustic states/events.
To enhance the effectiveness of ASR in low-resource scenarios, a combination of semi-supervised training and language adversarial TL is explored in [kumar2021exploration]. The research presented in [yi2020adversarial] suggests utilizing adversarial transfer learning to improve punctuation prediction performance. Specifically, a pre-trained BERT model is employed to transfer bidirectional representations to punctuation prediction models. The proposed approach is applied to the ASR task as the target task. Table 10 provides a summary of the performance achieved in recent studies involving adversarial TL for ASR.
Table 10: A summary of the recent ASR-based adversarial-language TL technique. Whereas the marks ($\uparrow$) and ($\downarrow$) indicate improvement and reduction, respectively. If many scenarios has been conducted in one metric, only the best result is mentioned.
Scheme
Model-based
ASR Tasks ($\mathbb{T}_{T}$)
Characteristic
Performance
[kumar2021exploration]
Adversarial SHL-Mode
( SincNet-CNN-LiGRU )
Use three Indian languages ( Hindi, Marathi, and Bengali ) cross-lingual to improve Hindi ASR
Semi-supervised
WER=5.5%
(25.65% WER $\downarrow$)
[yi2018language]
SHL Model
Improve the performance of low-resource ASR
Cross-lingual
10.1 % WER $\downarrow$
[yi2020adversarial]
BERT
Improve the performance of punctuation predicting
Multi-task
9.4% F1-score $\uparrow$
4.6 DTL-based ASR for medical diagnosis
DTL-based ASR has made significant advancements in the field of medicine, particularly in the early detection of diseases. These advancements have been observed in various medical domains, as outlined in Table 11, which includes:
4.6.1 Heart sound classification
The components of the heart sound encompass various elements. The first (S1) and second (S2) heart sounds are considered normal, whereas the third (S3) and fourth (S4) heart sounds are often associated with murmurs, and ejection clicks typically indicate certain illnesses or abnormalities. Koike et al. [koike2020audio] introduced a novel DTL approach based on Probabilistic Audio Neural Networks (PANNs), which involves pre-training a model on a large-scale audio dataset for the purpose of classifying heart sounds. Another approach for heart sound classification was proposed by Boulares et al. [boulares2020transfer]. Their method utilizes DTL on the Pascal public dataset, without any denoising or cleaning procedures, to establish an experimental benchmark. The primary objective is to provide a foundation of experimental results that can serve as a starting point for future research on cardiovascular disease (CVD) recognition using phonocardiogram (PCG)-based cardiac cycle vibration sounds. This proposed scheme addresses the absence of a CVD recognition benchmark and the lack of objective comparability among classification results, which tend to vary significantly.
4.6.2 Parkinson disease detection
Parkinson’s disease (PD) is a progressive neurodegenerative condition with a global impact. Accurate diagnosis of PD is crucial for improving daily activities and extending patients’ lives. However, predicting symptom changes and their impact on patients’ lives is challenging due to the variability in symptoms and disease progression. Traditional PD detection methods are often manual and require specialized expertise. Language impairment and atrophy are common early symptoms in over 90% of PD patients, characterized by reduced voice volume, monotonous and rapid speech, and eventual loss of audibility. Karaman et al. [43] proposed a robust automated PD detection approach using DTL-based ASR, where pre-trained models like SqueezeNet1_1, ResNet101, and DenseNet161 were fine-tuned and retrained. This scheme showed promising results in PD detection. To address the scarcity of speech data for PD and the distribution inconsistency among subjects, Li et al. [42] introduced a two-step unsupervised DTL algorithm called two-step sparse transfer learning (TSTL). This algorithm helps extract useful information from large amounts of unlabeled speech data, align the distribution of training and test sets, and preserve the original sample structure simultaneously. Another strategy proposed by Qing et al. [45] involves enhancing ASR for Parkinson’s patients using a pre-trained long short-term memory (LSTM) neural network model. To mitigate overfitting and reduce WER, the scheme employs frequency spectrogram masking data augmentation.
4.6.3 Other medical diagnosis
In the field of speech-based depression prediction, Harati et al. [harati2021speech] proposed a DTL method that utilizes a lightweight encoder and transfers only the encoder weights. This approach simplifies the runtime model for depression prediction. For individuals with dysarthria, Xiong et al. [xiong2020source] developed an improved DTL framework for robust personalized speech recognition models. They adapted the CNN-TDNN-F ASR AM onto the target dysarthric speakers using neural network weight adaptation. In the context of dysarthria speaking identification, Takashima et al. [takashima2019knowledge] proposed a method that transfers two types of knowledge from different datasets: the language-dependent characteristics of unimpaired speech and the language-independent characteristics of dysarthric speech. They focused on Japanese people with articulation disorders. Additionally, Sertolli et al. [sertolli2021representation] presented a novel feature representation for health states identification using an end-to-end DTL-based ASR framework. They utilized ASR DNNs as feature extractors, combined multiple feature representations using compact bilinear pooling (CBP), and employed an optimized RNN classifier for inference.
Table 11: A summary of a DTL-based ASR technique in medical diagnosis, whereas the marks ($\uparrow$) and ($\downarrow$) indicate the improvement and reduction, respectively. If many scenarios have been conducted in one metric, only the best result is mentioned.
Scheme
Model-based
ASR Tasks ($\mathbb{T}_{T}$) $\uparrow$
DTL Type
Performance
[43]
DenseNet-161
PD detection
Fine-tuning
Accuracy= 91.17%
[42]
TSTL-based CSC&SF
PD speech diagnosis
Unsupervised
Accuracy= 97.50%
[45]
Proposed four layers
PD speech
Fine-tuning
[…]
[koike2020audio]
PANN CNN14
Heart sound classification
Fine-tuning
UAR= 89.7%
[boulares2020transfer]
InceptionResNet-v2
PCG-based CVD classification
Fine-tuning
Accuracy= 0.89%
[harati2021speech]
EH-AC
Depression prediction
LHN (encoder weights)
27% AUC $\uparrow$
[xiong2020source]
CNN-TDNN-F
Dysarthric speech
Neural weight adapter
[…]
[takashima2019knowledge]
LAS
Dysarthric Speech
Multilingual
[…]
[sertolli2021representation]
Wav2Letter and
DeepSpeech
Health states classification
Transductive
UAR= 73.0% (8.6% $\uparrow$)
4.7 DTL-based ASR attacks and security
Adversarial examples are created by introducing small perturbations or noise to valid audio files or speech characteristics in order to manipulate or deceive ASR systems. These perturbations, although often imperceptible to the human auditory system and may be perceived as background noise by the ASR model, can lead to correct or incorrect classification of the inputs. For instance, by introducing a slight disturbance to the speech "At the still point, there the dance is," an ASR system may produce the incorrect transcription "At the tail point, there the tense is" [hu2019adversarial]. The concept behind attacking ASR systems lies in the vulnerability of these models to adversarial examples, which has prompted speech researchers to explore the creation of such examples. By generating adversarial examples for different representations of speech in the time or frequency domain, which capture various speech features and can be used as inputs to neural networks, researchers can manipulate ASR performance, either improving it or decreasing it. DTL plays a key role in achieving transferability, whereby adversarial examples crafted to attack a source model can also be effective in attacking target models that classify the same type of data. In this context, adversarial attacks on DTL-based ASR models can be categorized into two groups, as illustrated in Fig. 18.
Figure 18: Possible adversarial attacks in DTL-based ASR schemes.
4.7.1 Positive adversarial attacks
Encompassing various techniques that aim to enhance or ensure the efficacy of current ASR systems, these methods strive to improve the robustness and performance of acoustic models. In [sun2018training], the authors introduced a methodology that combines natural data with adversarial data to train a resilient acoustic model. Specifically, they focused on utilizing MFCC features and employed a gradient-based approach to generate adversarial MFCC features, taking into account the network model and input parameters for each mini-batch. By applying the teacher/student training concept, the neural network was trained using a combination of natural data and the generated adversarial data. The effectiveness of the proposed approach was validated through experiments conducted on the CHiME-4 and Aurora-4 tasks, utilizing a customized convolutional neural network (CNN).
4.7.2 Negative adversarial attacks
Encompassing various techniques that aim to diminish or undermine the effectiveness of existing ASR systems, these methods pose potential threats to the performance and robustness of acoustic models. In [abdullah2021hear], the authors propose a framework that utilizes the Google (Phone) model as the source model and investigates the impact of adversarial attacks on the target model (DeepSpeech 1). The adversarial attacks are applied to the audio waveform, specifically after the signal decomposition and thresholding processes, and the resulting manipulated input is fed into the ASR source model. According to [hu2019adversarial], the adversarial attack models can be categorized into two types based on the adversary’s objectives, knowledge, and background, as illustrated in Figure 18:
  • Adversary knowledge: It is divided into two types: white-box attack, which assumes the adversary has complete knowledge of $M_{T}$ including its architecture, training weights, and parameters; and black-box attack, which assumes the adversary has no access to $M_{T}$ and only knows its output like a regular user.
  • Adversarial specificity: It is divided into two types: non-targeted attack, which aims to make the adversarial example’s $M_{T}$ predict any incorrect class with the sole goal of compromising the ASR algorithm, and targeted attack, which aims to deceive $M_{T}$ into assigning the adversarial example to a specific class selected by the attacker, imposing specific instructions on the ASR scheme.
Moving on, Table 12 shows a summary of DTL-based adversarial models for existing works.
Table 12: A summary of the recent DTL-based ASR for adversarial attacks. Whereas the marks ($\uparrow$)and ($\downarrow$), indicate improvement and reduction respectively.
Scheme
Model-based
ASR Tasks ($\mathbb{T}_{T}$)
Target object
Adversary knowledge
Adversarial specificity
Performance
[sun2018training]
Aurora-4
Boost
MFCC
White-box
Targeted
23% WER $\downarrow$
[schonherr2018adversarial]
DNN-HMM (Kaldi)
Boost
Waveform
White-box
Targeted
Accuracy= 98 %
[zelasko2021adversarial]
DeepSpeech
Fool
Waveform
White-box
Targeted
[…]
[subramanian2020study]
VGG13
Fool (Dense_mel)
Mel-spectrogram
White-box
Non-targeted
SNR=29.06 dB
[carlini2018audio]
DeepSpeech
Fool (Speech-to-Text)
Waveform
White-box
Targeted
Attack success rate= 100%
[abdullah2021hear]
Google
Fool
(Deep-Speech 1)
Waveform
Black-box
Non-targeted
Attack success rate=87%
[kwon2019selective]
DeepSpeech
Fool (victime DeepSpeech)
Mel-frequency cepstrum
White-box
Targeted
Attack success rate=91.67%)
4.8 DTL-based ASR for other applications
DTL-based ASR has been applied in various fields. In the work by Boes et al. [12], DTL is explored for audio tagging and sound event detection tasks, where pre-trained auditory and visual features are incorporated into a baseline system using feature fusion. Arora et al. [arora2017study] use DTL to address the lack of annotated databases for audio event detection. Wang et al. [wang2020cross] propose an environment adaptation technique using DTL for deep speech enhancement models. Chen et al. [chen2018transfer] employ DTL for wearable devices to evaluate social speech in natural daily situations. Wu et al. [wu2020self] investigate self-supervised pre-trained speech for speech translation, while Zhu et al. [zhu2021conwst] propose a self-supervised bidirectional distillation system for low-resource speech translation. DTL is used in the speaker verification field by Hong et al. [hong2017transfer] for discriminative learning, and in the detection of marine mammal sounds by Lu et al. [lu2021detection] to classify different species using DTL with the AlexNet pre-trained model. Lastly, Azizah et al. [azizah2020hierarchical] employ hierarchical DTL for multilingual text-to-speech synthesis using DNNs for low-resource languages.
5 Open Issues and of Key challenges
Integrating advanced techniques like DTL, FL, and RL into ASR systems presents exciting opportunities but comes with its set of challenges. This section delves into the distinct challenges associated with each approach, emphasizing the critical areas that demand attention and innovation.
5.1 DTL and domain adaptation
This section discusses challenges and concepts related to DTL and DA in speech recognition, including distribution shift, feature space adaptation, label distribution shift, catastrophic forgetting, domain-invariant feature learning, sample selection bias, and hyperparameter optimization.
When applying a model trained on one domain (source) to another (target), a distribution shift often occurs, referred to as domain shift. Formally, if $P_{S}(X,Y)$ and $P_{T}(X,Y)$ represent the joint distributions of features $X$ and labels $Y$ in the source and target domains, respectively, the challenge arises when $P_{S}(X,Y)\neq P_{T}(X,Y)$. To address this, techniques focus on learning a transformation of the feature space to minimize the difference between the source and target distributions. This involves finding a mapping function $f:X\rightarrow Z$, where $Z$ is a latent space in which the distributions of transformed features $f(X_{S})$ and $f(X_{T})$ are more similar, quantified using measures such as the maximum mean discrepancy (MMD). Label distribution shift occurs when the distributions of labels ($P_{S}(Y)$ vs. $P_{T}(Y)$) differ, even if the feature distributions align. This poses challenges, especially with underrepresented classes in the target domain. Addressing this mathematically involves adjusting the model or learning process, possibly by re-weighting the loss function based on class distribution estimates. Catastrophic Forgetting is a risk during fine-tuning on a new domain, where the model may lose its performance on the original task. Balancing loss functions ($L_{S}$ for the source and $L_{T}$ for the target) is crucial, often weighted by a hyperparameter $\lambda$ to control their importance. Domain-invariant feature learning aims to learn features invariant across domains while remaining predictive. It involves optimizing a feature extractor $f$ and predictor $g$ to minimize the $D_{S}$ loss $L_{S}$ and the domain discrepancy (e.g., MMD). The problem of sample selection bias occurs in selecting samples for DA, affecting the effectiveness of adaptation strategies. Mathematically, addressing this bias involves weighting or selecting samples to minimize it, often using importance sampling or re-weighting techniques. Hyperparameter optimization is critical in DA, where the choice of hyperparameters (e.g., $\lambda$) significantly impacts performance. Finding the optimal hyperparameters typically involves solving complex optimization problems using techniques like grid search, random search, or Bayesian optimization on a validation set.
MMD^2[f(X_S), f(X_T)] = ∥1nS∑_i=1^n_Sϕ(x_S_i) - 1nT∑_j=1^n_Tϕ(x_T_j)^2
where $\phi(\cdot)$ represents the mapping to a reproducing kernel Hilbert space (RKHS), and $n_{S}$, $n_{T}$ are the numbers of samples in the source and target domains, respectively.
Moreover, the unification of DTL in ASR poses a challenge due to the varied mathematical formulations used in different studies. While efforts have been made to unify definitions and formulations, further work is needed for a consistent understanding of DTL. Speech-based DTL processing faces challenges compared to image-based processing due to potential mismatches between source and target databases arising from factors like language, speakers, age groups, ethnicity, and acoustic environments. The CTC approach, while promising, is limited by the assumption of frame independence. Cross-lingual DTL challenges include incorporating linguistic characteristics from multiple sources and integrating knowledge at different hierarchical levels, considering linguistic differences. Finally, computational burden remains a significant challenge in DTL and DA processes. Knowledge transfer between domains can incur additional computational costs, especially considering the extensive computational resources required for deep architectures inherent in DTL techniques.
5.2 FL-based
FL has significant potential for ASR systems, particularly in enhancing privacy and personalization. However, deploying this technology in ASR also comes with a set of challenges. Typically, in FL, data is inherently decentralized and can vary significantly across devices. This heterogeneity in speech data—due to differences in accents, dialects, languages, and background noise—can make it challenging to train a model that performs well across all nodes. Ensuring robustness and generalization of the ASR model under these conditions is a complex task. Moving on, FL requires periodic communication between the central server and the devices to update the model. For ASR systems, where models can be quite large, this can result in substantial communication overhead. Optimizing the efficiency of these updates, in terms of both bandwidth usage and energy consumption, especially on mobile devices, is a significant challenge. Besides, although FL is designed to enhance privacy by not sharing raw data, there are still privacy challenges. For instance, it’s possible to infer sensitive information from model updates. Ensuring that these updates do not leak private information about the users’ speech data is a critical concern that requires sophisticated privacy-preserving techniques like differential privacy or secure multi-party computation.
Additionally, one of the advantages of FL is the ability to personalize models based on local data. However, balancing personalization with the need for a generally effective model—especially in a diverse ecosystem with varying speech patterns—is challenging. Achieving this balance without compromising the model’s overall performance or the personalization benefits is a key challenge. Moving forward, FL systems need to manage potentially thousands or millions of devices participating in the training process. Scalability issues, including managing updates from such a large and potentially unreliable network of devices, ensuring consistent model improvements, and handling devices joining or leaving the network, are significant technical hurdles. Lastly, in FL, the distribution of data across devices is often non-identically distributed (non-IID). This means that the speech data on one device might be very different from that on another, leading to challenges in training a model that generalizes well across all devices. Overcoming the bias introduced by non-IID data is a major challenge in FL for ASR.
5.3 RL-based
Using RL in ASR systems offers promising avenues for improvement but also presents several challenges. Specifically, one of the primary challenges in applying RL to ASR is the issue of sparse and delayed rewards. In many ASR tasks, the system only receives feedback (rewards or penalties) after processing lengthy sequences of speech, making it difficult to attribute the reward to specific actions or decisions. This delay complicates the learning process, as the model struggles to identify which actions led to successful outcomes. Moreover, balancing exploration, trying new actions to discover their effects, with exploitation, using known actions that yield the best results, is a critical challenge in RL. In the context of ASR, this means the system must balance between adhering to known speech patterns and exploring new patterns or interpretations. Overemphasis on exploration can lead to inaccurate transcriptions, while excessive exploitation may prevent the model from adapting to new speakers or accents. Additionally, RL models typically require a significant amount of interaction data to learn effectively. In ASR, obtaining large volumes of labeled speech data, especially with user feedback, can be challenging and expensive. Additionally, RL algorithms can be sample-inefficient, meaning they need a lot of data before they start performing well, which can be a bottleneck in practical applications.
Moving forward, most ASR systems are built using supervised learning techniques that rely on vast amounts of annotated data. Integrating RL into these systems poses technical challenges, as it requires a different training paradigm that focuses on learning from user interactions and feedback rather than static datasets. Besides, using RL in ASR often involves collecting and analyzing user feedback and interactions to improve the model. This raises concerns about user privacy and data security, as sensitive information might be inadvertently captured and used for training. Ensuring that data is handled securely and in compliance with privacy regulations is a significant challenge.
Designing an appropriate reward function that accurately reflects the desired outcomes in ASR is challenging. The reward function must capture the nuances of speech recognition, such as accuracy, naturalness, and user satisfaction, which can be difficult to quantify. Poorly designed reward functions can lead to suboptimal learning outcomes or unintended behaviors. Lestly, ASR systems are used in a wide range of environments, from quiet offices to noisy streets. RL models need to adapt to these varying conditions, but training them to handle such diversity can be complex. The environment’s variability requires models that can generalize well across different acoustic conditions, which remains a challenge for RL-based ASR systems.
6 Future directions
6.1 Personalized data augmentation for dysarthric and older people
While ASR technologies have advanced significantly, especially in recognizing typical speech patterns, they still struggle to accurately identify speech from individuals with dysarthria or older adults [120]. Gathering extensive datasets from these groups is challenging due to mobility limitations often associated with these populations. In this context, personalized data augmentation plays a crucial role [121, 122]. Personalized data augmentation tailores the training process to accommodate the unique speech patterns and challenges associated with these groups. Dysarthria, a motor speech disorder, and the natural aging process can lead to speech that deviates from the normative models typically used to train ASR systems, making accurate recognition difficult [123]. Personalized data augmentation introduces a wider range of speech variations into the training dataset, including those specific to dysarthric speakers or older adults. This can include variations in speech rate, pitch, articulation, and clarity. By training on this augmented dataset, the ASR system learns to recognize and accurately transcribe speech that exhibits these characteristics [124]. Moreover, this helps the ASR models generalize better to unseen examples of speech from dysarthric speakers or older adults. This enhanced generalization is crucial for real-world applications where the system encounters a wide range of speech variations. Moving forward, personalized data augmentation can employ specific techniques tailored to the needs of dysarthric speakers or older adults, such as simulating the slurring of words, varying speech tempo, or introducing background noise [125], commonly challenging for these groups [126]. Techniques like pitch perturbation, temporal stretching, and adding noise can simulate real-world conditions more accurately for these users. A typical example is presented in [127], where a unique approach utilizes speaker-dependent generative adversarial networks has been proposed.
6.2 Multitask learning for ASR
[…]
MTL enhances the performance of ASR systems by leveraging the inherent relatedness of multiple learning tasks to improve the generalization of the primary ASR task. This approach allows the ASR model to learn shared representations that capture underlying patterns across different but related tasks, leading to several key benefits [128]. Typically, MTL encourages the ASR system to learn representations that are beneficial across multiple tasks. This can lead to more robust feature extraction, as the model is not optimized solely for transcribing speech but also for other related tasks, such as speaker identification or emotion recognition. This shared learning process helps in capturing a broader range of speech characteristics, which can improve the ASR system’s ability to handle varied speech inputs. Moving one, By simultaneously learning related tasks, MTL acts as a form of regularization, reducing the risk of overfitting on the primary ASR task. This is because the model must find a solution that performs well across all tasks, preventing it from relying too heavily on noise or idiosyncrasies specific to the training data of the main task. Besides, learning auxiliary tasks alongside the main ASR task can improve the model’s generalization capabilities. For example, learning to identify the speaker or the language can provide additional contextual clues that help the ASR system better understand and transcribe ambiguous audio signals.
Additionally, MTL can make more efficient use of available data by leveraging auxiliary tasks for which more data might be available. In scenarios where annotated data for ASR is limited, incorporating additional tasks with more abundant data can help improve the learning process and performance of the ASR system. Moreover, MTL allows ASR systems to better handle acoustic variability in speech, such as accents, dialects, or noisy environments, by incorporating tasks that directly or indirectly encourage the model to learn features that are invariant to these variations. Last but not least, modern ASR systems often employ DL architectures that can benefit from end-to-end learning strategies. MTL fits naturally into this paradigm, allowing for the joint optimization of multiple objectives within a single model architecture. This can simplify the training process and reduce the need for separately trained models or handcrafted features.
6.3 Federated multi-task learning and distillation for ASR
[…]
FMTL extends the concept of FL by allowing each client to learn a personalized model that addresses its specific task, while still benefiting from collaboration with other clients.This approach recognizes the heterogeneity in clients’ data distributions and tasks. Mathematically and compared with FL (Equation 3), federated multi-task learning (FMTL) can be formulated as:
$\min_{\theta_{1},\theta_{2},...,\theta_{K}}\sum_{k=1}^{K}F_{k}(\theta_{k})+\lambda R(\theta_{1},\theta_{2},...,\theta_{K})$ (7)
Different from FL, $R(\theta_{1},\theta_{2},...,\theta_{K})$ is a regularization term that encourages some form of similarity or sharing among the model parameters of different tasks, promoting collaboration among clients. $\lambda$ is a regularization coefficient that balances the trade-off between fitting the local data well and collaborating with other clients. FMTL has task-specific model parameters $\theta_{k}$ for each client, where only a single global model parameter $\theta$ in FL.
In this regard, FMTL offers a promising approach to improving ASR systems while also enhancing privacy and security measures. This learning paradigm extends the traditional FL model by enabling the simultaneous training of multiple tasks across distributed devices or nodes, without the need to share raw data [129]. FMTL leverages data from a wide range of devices and users, each potentially offering unique speech data, accents, dialects, and noise conditions. This diversity helps in training more robust ASR models that can perform well across various speech patterns and environments [130]. By learning from many tasks simultaneously, FMTL can personalize ASR models to individual users or specific groups without compromising the model’s general performance [131]. This is particularly beneficial for users with unique speech patterns, such as those with accents or speech impairments. Moreover, FMTL encourages the development of compact models that can handle multiple tasks efficiently. For ASR systems, this means that a single model can potentially perform speech recognition, speaker identification, and even emotion detection, reducing the computational overhead on client devices [132].
On the other hand, in FMTL, raw data remains on the user’s device and does not need to be shared or transferred to a central server. This inherently reduces the risk of data breaches and unauthorized access, as sensitive speech data is not centralized [133]. Additionally, FMTL can be combined with differential privacy techniques to further anonymize the model updates sent from devices to the central server. This ensures that the shared information does not reveal sensitive details about the data or the user, enhancing privacy protection [134]. Moving on, the aggregation process in FMTL can be secured using cryptographic techniques, ensuring that the aggregated model updates cannot be traced back to individual users. This secure aggregation process protects user privacy while allowing the benefits of collective learning [135]. Lastly, by aggregating model updates from a wide range of tasks and users, FMTL can improve the system’s robustness to malicious attempts at data poisoning. The diversity of inputs helps in diluting the impact of any adversarial data introduced to compromise the model.
End-to-End ASR models Continuing to develop and refine end-to-end deep learning models that directly map speech inputs to text outputs without the need for intermediate representations (like phonetic transcriptions) [mamyrbayev2023hybrid]. Advances in models such as Transformer and Conformer architectures, which can capture long-range dependencies in speech, are promising for improving ASR accuracy and efficiency [liu2023sfa]. Specifically, end-to-end models in ASR aim to directly convert speech input into text output using a single neural network architecture. This simplifies the ASR pipeline by bypassing traditional stages such as acoustic, pronunciation, and language modeling [yang2023attention].
An end-to-end ASR model maps an input sequence of speech features $X=(x_{1},x_{2},\ldots,x_{T})$ to an output sequence of tokens $Y=(y_{1},y_{2},\ldots,y_{N})$. The model function $f$ with parameters $\theta$ aims to minimize the difference between the predicted sequence and the ground truth. Two common loss functions are used [almadhor2023e2e]:
  • CTC Loss: Connectionist Temporal Classification (CTC) introduces a ’blank’ token for aligning sequences. The CTC loss is given by:
    $\mathcal{L}_{\text{CTC}}(\theta)=-\sum_{(X,Y)\in\mathcal{D}}\log P(Y|X;\theta)$
    where $\mathcal{D}$ is the dataset, and $P(Y|X;\theta)$ is the sequence probability.
  • Attention-based Seq2Seq Loss: This uses an encoder-decoder architecture with attention to predict each token. The loss is the negative log-likelihood:
    $\mathcal{L}_{\text{Seq2Seq}}(\theta)=-\sum_{(X,Y)\in\mathcal{D}}\log P(Y|X;\theta)$
Besides, model architectures can be described as follows:
  • Recurrent Neural Networks (RNNs) including Long Short-Term Memory (LSTM) networks for temporal dependencies [vander2023using].
  • Convolutional Neural Networks (CNNs) for capturing local patterns in speech features.
  • Transformer Models utilize self-attention mechanisms for parallel processing and superior performance [wang2022optimizing].
Gradient-based methods like stochastic gradient descent (SGD) or Adam are used to minimize the loss function, optimizing the model parameters $\theta$ to improve the speech-to-text mapping. However, end-to-end models face challenges such as data scarcity for low-resource languages, computational resource demands, and integrating external language models for linguistic improvements [qu2023emphasizing].
6.4 Domain-specific language modeling
[…]
$P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}$ (8)
Additionally, data augmentation plays a crucial role in ASR-based DA. Typically, data augmentation in domain-specific ASR involves generating synthetic training examples by altering existing recordings or using text-to-speech (TTS) systems to create new audio samples from domain-specific texts. Mathematically, augmentation techniques can include time stretching, pitch shifting, adding noise, or simulating room acoustics, each represented as transformations $T$ applied to the original audio signal $x(t)$:
$x^{\prime}(t)=T(x(t);\phi)$ (9)
where $\phi$ represents the parameters of the transformation. Moving on, regularization techniques are applied to prevent overfitting to the domain-specific data, ensuring the model maintains generalization capabilities. Techniques such as dropout, L1/L2 regularization, or elastic net involve adding terms to the loss function or modifying the optimization process to penalize large weights or complex models. For instance, L2 regularization adds a penalty equal to the square of the magnitude of coefficients:
$L^{\prime}(D;\theta)=L(D;\theta)+\lambda\|\theta\|^{2}$ (10)
where $\lambda$ is a regularization parameter.
While DSLM offers substantial benefits, challenges remain, such as data scarcity in niche domains, the need for continual model updates to keep pace with evolving language use, and ensuring privacy and security in sensitive domains like healthcare and finance. Addressing these challenges through innovative approaches in model training, data augmentation, and privacy-preserving technologies will be crucial for the advancement of domain-specific ASR systems.
Privacy preservation With the advancements in ML, DL, and DTL, ASR systems have become more practical and scalable. However, these systems also introduce serious privacy concerns due to the abundance of sensitive acoustic and textual information in speech data.
While open-source and offline ASR systems can mitigate privacy risks, online DTL-based systems can amplify these threats. Additionally, the transcription performance of offline and open-source ASR systems is typically inferior to cloud-based ASR systems, especially in real-world scenarios [ahmed2020preech]. In this context, the $D_{S}$ data may contain sensitive information that needs to be protected. Thus, preserving users’ privacy during the knowledge transfer from the source domain (SD) to the target domain (TD) becomes a critical issue.
To address this challenge, future research efforts should focus on integrating effective security and privacy protection strategies. Examples include decentralized DTL approaches utilizing blockchain technology [ul2020decentralized, wang2021enabling] and federated DTL methods [zhang2021federated, maurya2021federated]. These approaches aim to ensure privacy-preserving knowledge transfer in DTL-based ASR systems.
6.5 Interpretation of DTL models
DTL-based ASR models, despite their success, are often considered as "black box" systems lacking interpretability. This lack of interpretability raises doubts about the credibility and repeatability of their decisions, making it crucial to explain the reasoning behind their predictions. The field of explainable and interpretable ML/DL has gained increasing interest in various applications, including speech processing.
Several studies have explored the interpretability of DTL models. In one study [ramakrishnan2016towards], an agent was designed to explain how it learns a new task using prior common knowledge, aiming to enhance users’ trust and acceptance of the system results. Another study [kim2019the] defined interpretable features in a DTL algorithm and examined the relationship between the SD and TD in the task, focusing on the interpretability of the pretrained DTL model.
The work by Lee et al. [lee2021interpretable] introduced a knowledge distillation approach that generated interpretable embedding procedure (IEP) knowledge based on PCA and transferred it to the student network using a message passing neural network. This approach enhanced interpretability while maintaining accuracy through multi-task learning. Additionally, Carr et al. proposed an interpretable staged TL (iSTL) scheme for accurate and explainable classification of optical coherence tomography (OCT) scans with a small sample size [carr2021interpretable]. iSTL outperformed DTL techniques on unseen data, utilizing clinical features for predictions with interpretable attention maps.
Delving deeper into techniques for FL distillation (optimizing model size) within FL frameworks is essential. This exploration involves researching methods to compress neural ASR models effectively while ensuring their performance remains intact, especially tailored for edge devices with storage and computational constraints. It is imperative to investigate the trade-offs associated with reducing model size while maintaining performance metrics like WER. Developing strategies to strike a balance between downsizing models and preserving satisfactory performance levels within federated learning environments is crucial.
6.6 Recent RL techniques for ASR
Exploring incremental DRL approaches [136, 137, 138] in DRL-based ASR systems, could be very interesting. This approach involves the model continuously learning from newly acquired data and dynamically adjusting its ASR functionalities over time. By incrementally updating its knowledge base, the model can enhance its performance without necessitating full retraining, thus enabling continual enhancement of ASR systems. This capability not only fosters greater resilience and adaptability in speech recognition capabilities but also offers potential applications in scenarios where real-time adaptation to changing conditions is crucial, such as in noisy environments or with varying speaker accents. Moreover, incremental RL can potentially lead to more efficient use of computational resources, as the model only needs to focus on learning from new data, rather than reprocessing the entire dataset. Further research in this area could unlock new possibilities for ASR systems to evolve and improve over time, ultimately enhancing their usability and effectiveness in diverse real-world settings.
Although some ASR schemes based on RL have been proposed, there remains a notable scarcity in the application of RL techniques to enhance ASR methods. While policy gradient and Q-learning are commonly employed, the realm of RL encompasses various subcategories such as double deep Q-network (DDQN), actor-critic (AC), deep deterministic policy gradien (DDPG), and more [139], which hold promise for advancing ASR with innovative approaches. Researchers are encouraged to delve into these diverse DRL-based methods to further enrich the field of ASR for both AM and LM fields.
6.7 Online DTL
Online DTL combines the principles of TL and online learning with deep neural networks, enabling models to adapt in real-time to new tasks or data distributions. This approach is beneficial in dynamic environments where data arrives sequentially. DL models, specifically Deep Neural Networks (DNNs), learn through optimizing the weights $\theta$ to minimize a loss function $L$, which measures the discrepancy between predicted outputs $\hat{y}$ and true outputs $y$: $\theta^{*}=\arg\min_{\theta}L(D;\theta)$. On the other hand, TL improves learning in a new target task through the transfer of knowledge from a related source task, adapting a pre-trained model $\theta_{S}$ on $D_{S}$ to a $D_{T}$: $\theta_{T}^{*}=\arg\min_{\theta}L(D_{T};\theta_{T})$. In this regard, online learning updates the model incrementally as new data $(x_{t},y_{t})$ arrives:
$\theta_{t+1}=\theta_{t}-\alpha_{t}\nabla_{\theta}L(y_{t},f(x_{t};\theta_{t}))$ (11)
Where $\alpha_{t}$ is the learning rate, and $\nabla_{\theta}L$ is the gradient of the loss with respect to $\theta$. Moving on, online TL integrates these concepts to continuously adapt a deep learning model to new tasks or data streams, often involving techniques such as feature extraction, fine-tuning, model adaptation, and continual learning. The adaptation process at each time step $t$ can be viewed as $\theta_{t+1}^{*}=\arg\min_{\theta}L(D_{T_{t}};\theta_{T_{t}})$. Where $D_{T_{t}}$ represents the data available at time $t$, including new target domain data.
The approach of online DTL offers a forward-looking solution to this issue. Particularly, discrepancies in class distributions and the representation of features between the $D_{S}$ and $D_{T}$ amplify the complexity of online DTL [140]. To navigate the complexities mentioned, online DTL has been dissected into two primary methodologies. The first, known as homogeneous online DTL, operates on the premise of a unified feature space across both domains. Conversely, heterogeneous online DTL acknowledges the distinct feature spaces intrinsic to each domain [141]. An exemplary solution to the challenges of heterogeneous online DTL includes leveraging unlabeled instances of co-occurrence to forge a connective bridge between the $D_{S}$ and $D_{T}$, facilitating the precursor to knowledge transfer [142]. Furthering the discourse, online DTL augmented with extreme learning machines introduces a novel framework [143]. Addressing the challenge of limited data in the $D_{T}$, the technique of transfer learning with lag (TLL), rooted in shallow neural network embeddings, has been applied. This method ensures the continuity of knowledge transfer, notwithstanding fluctuations in the feature set.
6.8 Transformers and LLMs-based ASR
[…]
Similarly, the LM aspect of LLM based on DTL can enhance ASR by leveraging TL. Pre-training LLM on vast text corpora equips it with extensive language representations, aiding in addressing diverse language challenges. Fine-tuning enables adaptation to specific language characteristics, improving transcription accuracy and contextual appropriateness. Both AM and LM fine-tuning can benefit from DA, incorporating target domain data to tailor models, reducing domain mismatch, and enhancing effectiveness and generalization. Utilizing LLM for objective ASR testing and MOS evaluation involves compiling diverse datasets, fine-tuning LLM, and integrating it into the ASR system. Evaluation metrics like CER and Pearson’s correlation gauge system performance, guiding further fine-tuning iterations for improved results. This iterative process ensures the ASR system’s continual enhancement and accurate MOS scale generation. Researchers are invited to explore these gaps and advance the integration of more advanced AI techniques such as DRL and FL into both AM and LM domains within transformer-based models. Additionally, there is a need for further investigation into Transformers and DTL-based ASR schemes specifically tailored for the LM domain. Closing these gaps will contribute to the development of more robust and effective language models across various applications.
7 Conclusion
In conclusion, recent advancements in deep learning have presented both challenges and opportunities forASR. Traditional ASR systems require extensive training datasets, often including confidential information, and consume significant computational resources. However, the demand for adaptive systems capable of performing well in dynamic environments has spurred the development of advanced deep learning techniques such as DTL, FL, and DRL, with all their variant techniques. These advanced techniques address issues related to DA, privacy preservation, and dynamic decision-making, thereby enhancing ASR performance and reducing computational costs.
This survey has provided a comprehensive review of DTL, FL, and RL-based ASR frameworks, offering insights into the latest developments and helping researchers and professionals understand current challenges. Additionally, the integration of transformers, powerful DL models, has been explored for their ability to capture complex dependencies in ASR sequences. By presenting a structured taxonomy and conducting critical analyses, this paper has shed light on the strengths and weaknesses of existing frameworks, as well as highlighted ongoing challenges. Moving forward, further research is needed to overcome these challenges and unlock the full potential of advanced DL techniques in ASR. Future work should focus on refining existing approaches, addressing privacy concerns in FL, improving RL algorithms for ASR optimization, and exploring innovative ways to leverage transformers for more efficient and accurate speech recognition. By continuing to innovate and collaborate across disciplines, we can push the boundaries of ASR technology and realize its transformative impact on various fields, including healthcare, communication, and accessibility.

### v2

@@
 DL techniques assume training and testing data originate from the same domain, which is not always true.
-Advanced DL techniques like deep transfer learning (DTL), federated learning (FL), and reinforcement learning (RL) address these issues.
-DTL allows high-performance models using small yet related datasets, FL enables training on confidential data without dataset possession, and RL optimizes decision-making in dynamic environments, reducing computation costs.
-This survey offers a comprehensive review of DTL, FL, and RL-based ASR frameworks, aiming to provide insights into the latest developments and aid researchers and professionals in understanding the current challenges.
+Advanced DL techniques like deep transfer learning (DTL), federated learning (FL), and deep reinforcement learning (DRL) address these issues.
+DTL allows high-performance models using small yet related datasets, FL enables training on confidential data without dataset possession, and DRL optimizes decision-making in dynamic environments, reducing computation costs.
+This survey offers a comprehensive review of DTL, FL, and DRL-based ASR frameworks, aiming to provide insights into the latest developments and aid researchers and professionals in understanding the current challenges.
 Additionally, Transformers, which are advanced DL techniques heavily used in proposed ASR frameworks, are considered in this survey for their ability to capture extensive dependencies in the input ASR sequence.
-The paper starts by presenting the background of DTL, FL, RL, and Transformers and then adopts a well-designed taxonomy to outline the state-of-the-art (SOTA) approaches.
+The paper starts by presenting the background of DTL, FL, DRL, and Transformers and then adopts a well-designed taxonomy to outline the state-of-the-art (SOTA) approaches.
 Subsequently, a critical analysis is conducted to identify the strengths and weaknesses of each framework.
@@
 ASR technology encompasses the analysis of (i) acoustic, lexical, and syntactic aspects; and (ii) semantic understanding.
-The acoustic model (AM) processing includes speech coding [1], enhancement [2], and source separation [3], alongside securing speech via steganography [4, 5, 6] and watermarking [7, 8, 9].
+The acoustic model (AM) processing includes speech coding [1], enhancement [2], source separation [3, 2], alongside securing speech via steganography [4, 5, 6] and watermarking [7, 8, 9].
 These components are integral to audio analysis.
 On the other hand, the semantic model (SM), often identified as language model (LM) processing in literature, involves all natural language processing (NLP) techniques.
-This AI branch aims at teaching computers to understand and interpret human language, serving as the basis for applications like music information retrieval [10], sound file organization [11], audio tagging, and event detection (ED) [12], as well as converting speech to text and vice versa [13], detecting hate speech [14], and cyberbullying [15].
+This AI branch aims at teaching computers to understand and interpret human language, serving as the basis for applications like music information retrieval [10], sound file organization [11], audio tagging (AT), and event detection (ED) [12], as well as converting speech to text and vice versa [13], detecting hate speech [14], and cyberbullying [15].
 Employing NLP across various domains enables AI models to effectively comprehend and respond to human inputs, unveiling extensive research prospects in diverse sectors.
 Abbreviations
+AAC
+automated audio captioning
+AC
+actor-critic
 AI
@@
 audio tagging
+BERT
+bidirectional encoder representations from transformers
 CAFT
@@
 character error rate
+CLDNN
+convolutional long-short term deep neural network
 CNN
@@
 domain adaptation
+DDPG
+deep deterministic policy gradien
+DDQN
+double deep Q-network
 DL
 deep learning
+DNN
+deep neural network
 DRL
 deep reinforcement learning
+DSC
+depthwise separable convolutions
+DSLM
+domain-specific language modeling
 DTL
@@
 event detection
+ESPNet
+end-to-end speech processing network
 FCF
 feature correlation-based fusion
+FedAvg
+federated averaging
 FedNST
@@
 federated learning
+FMTL
+federated multi-task learning
 FR
 form recognition
+GAN
+generative adversarial network
+GLDPT
+global-local dual-path Transformer
+HFL
+horizontal federated learning
 HMI
@@
 hidden Markov models
+HSD
+hate speech detection
+ISTFT
+inverse short-time Fourier transform
+LLM
+large language model
 LM
@@
 linear predictive coding
+LRF
+low-rank factorization
 LSTM
 long short term memory
+MAE-AST
+masked autoencoding audio spectrogram Transformer
 mAP
 mean average precision
+MDP
+Markov decision process
 MFCC
@@
 machine learning
+MMD
+maximum mean discrepancy
+MOS-LQO
+mean opinion score-listening quality objective
+MTL
+multitask learning
 NLP
 natural language processing
+non-IID
+non-identically distributed
 NT
@@
 perceptual evaluation of speech quality
+QCNN
+quantum CNN
 RER
 relative error rate
-RL
-reinforcement learning
 RNN
@@
 sequence-to-sequence
-S2S
-sequence-to-sequence
+SARSA
+state–action–reward–state–action
+SCST
+self-critical sequence training
 SD
 source domain
+SE
+speech enhancement
 SER
@@
 self-supervised audio spectrogram transformer
+STFT
+short-time Fourier tranform
+SVD
+singular value decomposition
 SWBD
 switchboard
+TBE
+two-branch encoder
 TD
 target domain
+TNet
+Transformer network
 TNR
@@
 transformer-recurrent-U network
+TSTNN
+Transformer-based neural network
+VFL
+vertical federated learning
 WER
@@
 wall street journal
-GAN
-generative adversarial network
-MTL
-multitask learning
-FMTL
-federated multi-task learning
-DSLM
-domain-specific language modeling
-LLM
-large language model
-DDQN
-double deep Q-network
-AC
-actor-critic
-SARSA
-State–action–reward–state–action
-DDPG
-deep deterministic policy gradien
 Recent advancements in ASR have been significantly propelled by the evolution of deep learning (DL) methodologies.
@@
 This mechanism is particularly effective in ASR, facilitating the detailed modeling of contextual nuances and the interconnections among acoustic signals, essential for accurate transcription.
-Models such as the Transformer Transducer, Conformer, and ESPnet, leveraging self-attention and parallel processing, have achieved leading performance in ASR tasks.
+Models such as the Transformer Transducer, Conformer, and end-to-end speech processing network (ESPNet), leveraging self-attention and parallel processing, have achieved leading performance in ASR tasks.
 Their robustness across diverse languages further underscores their capability to adapt to a wide range of linguistic features and acoustic variations, making Transformers an exceptionally promising option for enhancing ASR systems, surpassing the constraints of conventional models.
@@
 Recently, the authors presented an ASR review focused on DTL for ASR [36].
-Table 1 presents a summary of the main contributions of the proposed ASR review compared to other existing ASR reviews/surveys.
-This survey article offers several significant enhancements and additions compared to previous ASR surveys.
-Firstly, it consolidates works that utilize both ASR and advanced DL approaches, providing a comprehensive overview of their intersection.
+Table 1 presents a summary of the main contributions of the proposed DTL review compared to other existing DTL reviews/surveys.
+This survey article offers several significant enhancements and additions compared to previous DTL surveys.
+Firstly, it consolidates works that utilize both DTL and advanced DL approaches, providing a comprehensive overview of their intersection.
 Secondly, it provides performance evaluation results of all considered approaches.
-Thirdly, it includes metrics and dataset reviews used in ASR models.
+Thirdly, it includes metrics and dataset reviews used in DTL models.
 Furthermore, it tackles ongoing challenges and consequently proposes future directions.
@@
   • Proposing future directions to enhance the performance of advanced DL-based ASR solutions and predicting the potential advancements in the field.
-Table 1: Contribution comparison of the proposed contribution against other hand ASR review.
+Table 1: Contribution comparison of the proposed contribution against other hand DTL review.
 The tick mark (✓) indicates that the specific field has been addressed, whereas the cross mark () means addressing the specific fields has been missed.
-Refs Year
-Description of the survey/review
-Advanced DL methods
-Performances
-Metrics
-Dataset
-Current
-Future
-DRL
-FL
-TL
-Transf.
-evaluation
-review
-challe/Gaps
-directions
-[31] 2018
-ASR review for error errors detection and correction
-[34] 2019
-Systematic review on DL-based speech recognition
-[24] 2020
-ASR survey for Indian languages
-[23] 2020
-ASR survey for Portuguese language
-[25] 2020
-ASR survey for Turkich language
-[35] 2021
-ASR survey
-[27] 2021
-ASR survey for tonal languages
-[32] 2022
-Unsupervised ASR review
-[30] 2022
-ASR Systematic review for children
-[29] 2022
-ASR survey for limited vocabulary
-[26] 2022
-ASR Systematic review for Arabic language
-[28] 2022
-Bilingual ASR review
-[33] 2023
-ASR survey on neural network techniques
-[36] 2023
-ASR based on DTL review
-Our 2024
-ASR review on advanced DL techniques
+Refs Year Description of the survey/review Advanced DL methods Performances Metrics Dataset Current Future
+DRL FL DTL Transf. evaluation review challe/Gaps directions
+[31] 2018 ASR review for errors detection and correction
+[34] 2019 Systematic review on DL-based speech recognition
+[24] 2020 ASR survey for Indian languages
+[23] 2020 ASR survey for Portuguese language
+[25] 2020 ASR survey for Turkish language
+[35] 2021 ASR survey
+[27] 2021 ASR survey for tonal languages
+[32] 2022 Unsupervised ASR review
+[30] 2022 ASR Systematic review for children
+[29] 2022 ASR survey for limited vocabulary
+[26] 2022 ASR Systematic review for Arabic language
+[28] 2022 Bilingual ASR review
+[33] 2023 ASR survey on neural network techniques
+[36] 2023 ASR based on DTL review
+Our 2024 ASR review on advanced DL techniques
 1.3 Review methodology
@@
 The current section provides an introduction to the paper.
-Section 1.3 providing background on AM and LM, and reviewing evaluation metrics and datasets utilized in ASR.
+Section 2 providing background on AM and LM, and reviewing evaluation metrics and datasets utilized in ASR.
 Moving forward, Section 3 delves into a comprehensive review of recent advancements in ASR utilizing advanced DL approaches, including Transformers, DTL, FL and DRL.
-Sections 5 and 6 respectively address the existing challenges and future directions concerning advanced DL-based ASR.
+Sections 4 and 5 respectively address the existing challenges and future directions concerning advanced DL-based ASR.
 Finally, Section 6 presents concluding remarks summarizing the key findings of the paper.
+Figure 2 presents a structured roadmap, offering a comprehensive guide to assist readers in navigating through the various sections and subsections of the paper.
 Figure 2: Survey roadmap: A guide for navigating paper sections and subsections.
-table[t!] Literature acquisition databases. to be modified Database Research Articles Conference Papers Book Chapter Total ACM 5 1 6 Elsevier 26 26 Springer 23 10 33 IEEE 39 28 67 Others 44 29 70
-2 Overview of DTL techniques for speech recognition
-2.1 Taxonomy of existing DTL techniques
-To date, there is no standardized and comprehensive technique for classifying DTL into categories.
-However, DTL algorithms could be classified into several types depending on what, when, and how knowledge is transferred:
-(a) What knowledge is transferred?
-Enquires about which characteristics of knowledge are transferable across domains or tasks.
-Some information is particular to certain domains or tasks, while other knowledge is shared across domains and can aid increase performance in the target task or domain.
-Based on this definition, DTL could be feature-based, instance-based, relation-based, or model-based.
-(b) How is knowledge transferred? Enquires about which learning algorithms must be implemented to transfer the knowledge.
-(c) When is knowledge transferred? Inquires as to when and under what circumstances knowledge should or should not be transferred.
-Furthermore, researchers have proposed taxonomies to categorize DTL-based ASR techniques.
-For instance, Niu et al.
-[21] present a taxonomy with two levels.
-The first level consists of four sub-groups based on the availability of labeled data and the data modality in the source and target domains.
-These sub-groups include inductive DTL, transductive DTL, cross-modality DTL, and unsupervised DTL [80].
-Table 2 provides a summary of these possibilities.
-Moreover, each sub-group at the first level can be further divided into four distinct learning types: learning on instances, learning on features, learning on parameters, and learning on relations.
-[…]
-Domains Tasks
-Math. propriety
-Sub-categories / Usage
-Traditional ML/DL $\mathbb{D}_{S}=\mathbb{D}_{T}$ $\mathbb{T}_{S}=\mathbb{T}_{T}$
-$X_{S}\neq X_{T}$,
-$Y_{S}=Y_{T}$
-ASR model trained with $X_{S}$ database and used to recognise $X_{T}$ database.
-Inductive DTL $\mathbb{D}_{S}\cong\mathbb{D}_{T}$ $\mathbb{T}_{S}\neq\mathbb{T}_{T}$
-$X_{S}\neq X_{T}$,
-$Y_{S}\exists,Y_{T}\exists$
-If $Y_{S}\exists$, DTL is multitask learning. If $Y_{S}\nexists$, DTL is self-taught learning, thus $\chi_{S}\cong\chi_{T}$.
-Transductive DTL $\mathbb{D}_{S}\neq\mathbb{D}_{T}$ $\mathbb{T}_{S}=\mathbb{T}_{T}$
-$P(X_{S})\neq P(X_{T})$,
-$Y_{S}\exists,Y_{T}\nexists$ ,
-$\chi_{S}=\chi_{T}$
-When $\chi_{S}=\chi_{T}$, DTL is is related to DA. If $\mathbb{D}_{T}\exists!$ and $\mathbb{T}_{T}\exists!$, DTL is used for sample selection bias or covariate shift.
-Cross-modality DTL $\mathbb{D}_{S}\neq\mathbb{D}_{T}$ $\mathbb{T}_{S}\neq\mathbb{T}_{T}$
-$P(Y_{S}/X_{S})\neq P(Y_{T}/X_{T})$,
-$Y_{S}\neq Y_{T}$, $\chi_{S}\neq\chi_{T}$
-i,e. the dataset $X_{S}$ of $\mathbb{D}_{S}$ is speech data, and the dataset $X_{T}$ of $\mathbb{D}_{T}$ is text data.
-Unsupervised DTL $\mathbb{D}_{S}\varsubsetneq\mathbb{D}_{T}$ $\mathbb{T}_{S}\varsubsetneq\mathbb{T}_{T}$
-$Y_{S}\nexists,Y_{T}\nexists$
-DTL used for clustering, dimensionality reduction, and density estimation, etc.
-Inductive DTL In comparison to classical ML, which may be used as a reference for DTL comparison, and given that the target tasks $\mathbb{T}_{T}$ are distinct from the source tasks $\mathbb{T}_{S}$, the goal of inductive DTL is to enhance the target prediction function $\mathbb{F}_{T}$ in the TD, mentioned above in subsection LABEL:sub12.
-However, the SD $\mathbb{D}_{S}$ and TD $\mathbb{D}_{T}$ may not always be the same (Table 2).
-The inductive DTL can be stated similarly to the following two cases, depending on whether labeled or unlabeled data is available:
-(a) Multi-task DTL: The SD has a huge labeled database ($X_{S}$ labeled with $Y_{S}$), which is a distinctive form of multi-task learning.
-However, with multi-task approaches, many tasks $(T_{1},T_{2},\dots,T_{n})$ are learned at the same time (in parallel), including both source and target activities (tasks).
-(b) Sequential DTL: ( Commonly known as self-taught learning) Dataset is not labeled in the SD ($X_{S}$ is not labeled with $Y_{S}$) but the labels are available in the destination domain ($X_{T}$ is labeled with $Y_{T}$).
-Sequential learning is a DL system that can be realized in two steps for classification purposes.
-The first step is the feature representation transfer, which is learned from a large collection of the unlabeled datasets, and the second stage is when this learned representation is applied to labeled data to accomplish classification tasks.
-Hence, sequential DTL is a method of sequentially learning a number of activities (Tasks).
-The spaces between the source and destination domains may differ.
-For example, let suppose we have a pre-trained model $M$ and consider applying DTL to a number of tasks $(T_{1},T_{2},\dots,T_{n})$.
-We learn a specific task $\mathbb{T}_{T}$ at each time step $t$, which is slower than multi-task learning.
-However, when not all the tasks are present during training time, it might be beneficial.
-Sequential DTL is additionally classified into several types [alyafeai2020survey]:
-  • 1-
-    Fine-tuning: The principle is to learn a new function $\mathbb{F}_{T}$ that translates the parameters $\mathbb{F}_{T}(W_{S})=W_{T}$ by using $M$, given a pre-trained model $M_{S}$ having $W_{S}$ as weights and target task $\mathbb{T}_{T}$ having $W_{T}$ as weights.
-The settings can be adjusted across all layers or just some of (Fig.
-3 (a)).
-The learning rate for each layer could be distinct (discriminative fine tuning).
-A new set of parameters $K$ could be added to most of the tasks so that $\mathbb{F}_{T}(W_{T},K)=W_{S}\circ K$.
-  • 2-
-    Adapter modules: Given an $M_{S}$ model that has been pre-trained and output $W_{S}$, for a target task $\mathbb{T}_{T}$.
-The adapter module aims to lunch a different set of parameters $K$ that is too much less than $W_{S}$, i.e, $K\ll W_{S}$.
-$K$ and $W_{S}$ must have the ability to be decomposed into more compact modules such that, $W_{S}=\{w\}_{n}$ and $K=\{k\}_{n}$.
-The adapter module permit learning the following new function $\mathbb{F}_{T}$:
-    $\mathbb{F}_{T}(K,W_{S})=k_{1}^{\prime}\circ w_{1}\circ\dots k_{n}^{\prime}\circ w_{n}.$ (1)
-    According to equation 1, during the adaptation procedure, the set of original weights $W_{S}=\{w\}_{n}$ is left unaltered, but the set of weights K is changed to $K^{\prime}=\{k^{\prime}\}_{n}$.
-The principle of the adaptation domain is illustrated in Fig.
-3 (b).
-  • 3-
-    Feature based: Interested only in learning concepts and representations, at various levels, such as, word, character, phrase, or paragraph embedding $E$.
-The collection of $E$ based on a model $M$ remains unaltered, i.e., $\mathbb{F}_{T}(W_{S},E)=E\circ W^{\prime}$, in the way that $W^{\prime}$ is fine-tuned.
-For example, researchers have applied the generative adversarial network (GAN) principle to DTL where, the generators send features from the SD and the TD to a discriminator, which determines the source of the features and feeds the result back to the generators until they can no longer be distinguished.
-In this procedure, GAN obtains the common properties of two domains, as shown in Fig.
-3 (c).
-  • 4-
-    Zero-shot: Is the easiest method among all of the others.
-Making the assumption that the parameters $W_{S}$ can’t be modified or add $K$ as a new parameters to a pre-trained model $M_{S}$ using $W_{S}$.
-To put this into context, in zero-shot there is no training technique to optimize or learn new parameters.
-Figure 3: Structures of: (a) Fine-tuning, (b) DA, and (c) DTL-based GAN.
-2.1.1 Transductive DTL
-Compared to the traditional ML, which can be considered as a reference for DTL comparison, and given that the TDs $\mathbb{D}_{T}$ are distinct from the SDs $\mathbb{D}_{S}$.
-The SD has a labeled dataset ($X_{S}$ labeled with $Y_{S}$), whereas the TD has no labeled dataset, the source and target tasks are equal ( Table 2).
-The goal of transductive DTL is to build the target prediction function $\mathbb{F}_{T}$ in the $\mathbb{D}_{T}$ by knowledge of the $\mathbb{D}_{S}$ and $\mathbb{T}_{T}$.
-Furthermore, the transductive DTL environment may be further classified into two categories depending on different conditions between the source and destination domains:
-(a) Domain adaptation (DA): The feature spaces across domains, $\chi_{S}$ and $\chi_{T}$, are the identical, but the marginal probability distributions of the input dataset are not, $P(Y_{S}/X_{S})\neq P(Y_{T}/X_{T})$.
-For example, an assessment may be done on the topic of the resort in the $\mathbb{D}_{S}$ and it will be used to train a model for restaurants in the $\mathbb{D}_{T}$.
-DA is mostly effective when the $\mathbb{T}_{T}$ has a distinct distribution or there is a scarcity of labeled data.
-(b) Cross-modality DTL: Also known as cross-lingual DTL in the spoken language field, most DTL methods, more or less, a connection in feature spaces or label spaces is required between $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$.
-In other words, DTL can only occur when the source and destination data are both in the same modality, like video, speech, or text.
-cross-lingual DTL, in contrast to all other DTL approaches, is one of the most complicated issues in DTL.
-It is assumed that the feature spaces of the source and destination domains are completely distinct ($\chi_{S}\neq\chi_{T}$), as in speech-to-image, image-to-text, and text-to-speech.
-Furthermore, the label spaces of source $Y_{S}$ and destination $Y_{S}$ domains might differ ($Y_{S}\neq Y_{T}$).
-(c) Unsupervised DTL: Intends to enhance the learning of the target predictive function $\mathbb{F}_{T}$ in $\mathbb{D}_{T}$ using the knowledge in $\mathbb{D}_{S}$ and $\mathbb{T}_{S}$, where $\mathbb{T}_{S}$ different from $\mathbb{T}_{T}$ but related, and $Y_{S}$ and $Y_{T}$ are not visible, given a SD $\mathbb{D}_{S}$ with a learning task $\mathbb{T}_{S}$, a TD $\mathbb{D}_{T}$ and a matching learning task $\mathbb{T}_{T}$ ($\mathbb{D}_{S}$ different from $\mathbb{D}_{T}$, but related).
-2.1.2 Adversarial DTL
-In contrast to the methods described above for DTL, adversarial learning [wang2020transfer] aids in the learning of more transferable and discriminative representations.
-The work in [ganin2016domain], was the first that introduced the domain-adversarial neural network (DANN).
-Instead of using a predefined distance function like maximum mean discrepancy (MMD), the core idea is to use a domain-adversarial loss in the network.
-This has greatly aided the network’s ability to learn more discriminative data.
-Many studies have used domain-adversarial training as a result of DANN’s idea [bousmalis2016domain, chen2019joint, long2017deep, zhang2018collaborative].
-All of the previous work ignores the different effects of marginal and conditional distributions in adversarial TL, whereas in [wang2020transfer], the proposed scheme, named dynamic distribution alignment (DDA), can dynamically evaluate the importance of each distribution.
 2 Background
@@
 For n-gram models, this involves calculating the conditional probability of a word given the previous $n-1$ words [40]:
-$P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}$ (2)
+$\scriptsize{P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}}$ (1)
 In the context of ASR, the LM complements the AM by providing linguistic context.
@@
 These metrics serve as crucial evaluation criteria for experimental outcomes, as evidenced in studies such as [42, 43, 44].
-Additionally, there are ASR-specific metrics, which are detailed in Table LABEL:tab:metrics.
-Metric
-Formula
-Description
-WER
-$\mathrm{\frac{S+D+I}{N}=\frac{S+D+I}{H+S+D}}.$
-The word error rate (WER) serves as a frequently utilized metric to assess the performance of Automatic Speech Recognition (ASR).
+Additionally, there are ASR-specific metrics, which are detailed in Table 2.
+Table 2: An overview of the metrics employed for evaluating ASR methods.
+Metric Formula Description
+WER $\mathrm{\frac{S+D+I}{N}=\frac{S+D+I}{H+S+D}}.$ The word error rate (WER) serves as a frequently utilized metric to assess the performance of Automatic Speech Recognition (ASR).
 It is computed by determining the ratio of incorrectly recognized words to the overall number of processed words [45, 17, 46].
@@
 Instead of WER, the character error rate (CER) has been employed, while adhering to the same evaluation principle.
-PESQ and MOS-LQO
-$\mathrm{MOS-LQO}=0.999+\frac{4.999-0.999}{1+e^{-1.4945.PESQ+4.6607}}$
-PESQ serves as an objective technique for evaluating the perceived quality of speech [47].
+PESQ and MOS-LQO $\mathrm{MOS-LQO}=0.999+\frac{4.999-0.999}{1+e^{-1.4945.PESQ+4.6607}}$ PESQ serves as an objective technique for evaluating the perceived quality of speech [47].
 The assessment involves assigning numerical scores within the range of -0.5 to 4.5.
-Additionally, a correlation can be established between MOS and PESQ scores, giving rise to a novel evaluation metric termed the mean opinion score-listening quality objective (MOS-LQO), also identified as PESQ Rec.862.1.
+Additionally, a correlation can be established between MOS and perceptual evaluation of speech quality (PESQ) scores, giving rise to a novel evaluation metric termed the mean opinion score-listening quality objective (MOS-LQO), also identified as PESQ Rec.862.1.
 [22]
-RTF
-RTF=Total Processing TimeTotal $\mathrm{RTF=\frac{\text{Total Processing Time}}{\text{Total Duration}}}$
-[…]
-RTF serves as a standard metric to assess the processing time cost of an ASR system. It represents the average processing time required for one second of speech
-RER
-$\frac{(E_{\text{baseline}}-E_{\text{proposed}})}{E_{\text{baseline}}}\times 100\%$
-The relative error rate (RER) expresses the percentage error rate achieved by the proposed DL model compared to the baseline.
+RTF RTF=Total Processing TimeTotal $\mathrm{RTF=\frac{\text{Total Processing Time}}{\text{Total Duration}}}$Ac RTF serves as a standard metric to assess the processing time cost of an ASR system.
+It represents the average processing time required for one second of speech
+RER $\frac{(E_{\text{baseline}}-E_{\text{proposed}})}{E_{\text{baseline}}}\times 100\%$ The relative error rate (RER) expresses the percentage error rate achieved by the proposed DL model compared to the baseline.
 $E_{\text{baseline}}$ is the error rate of the baseline model.
 $E_{\text{proposed}}$ is the error rate of the proposed model or method.
-D
-$\frac{\sum_{i=1}^{n}\Big{(}1-\frac{\sum_{j=1}^{n}a_{ij}\cdot|i-j|}{\max(|i-1|,|i-2|,\ldots,|i-n|)}\Big{)}}{n}\hskip 28.45274pt$
-Diagonal centrality of an attention matrix (D) is defined as the mean value across the centrality of all its rows.
+D $\frac{\sum_{i=1}^{n}\Big{(}1-\frac{\sum_{j=1}^{n}a_{ij}\cdot|i-j|}{\max(|i-1|,|i-2|,\ldots,|i-n|)}\Big{)}}{n}\hskip 28.45274pt$ Diagonal centrality of an attention matrix (D) is defined as the mean value across the centrality of all its rows.
 where $j$ represents the index of each column, $n$ signifies the length of the input sequence, $a_{ij}$ denotes the attention weight between the $i$-th and $j$-th elements of the input sequence, and $|i-j|$ signifies the distance between the $i$-th and $j$-th elements of the input sequence [48].
@@
 Each utterance is stored as a one-second WAVE format file with 16-bit single-channel at 16KHz rate.
-It involves 2,618 speakers AURORA4 [57] Compare front-ends for large vocabulary recognition performance.
+It involves 2,618 speakers Aurora-4 [57] Compare front-ends for large vocabulary recognition performance.
 Aurora-4 is a speech recognition dataset derived from the WSJ corpus, offering four conditions (Clean, channel, noisy, channel+noisy) with two microphone types and six noise types, totaling 4,620 utterances per set.
@@
 Traditional statistical LMs, such as backoff n-gram LMs, have been widely used due to their simplicity and reliability.
-However, bidirectional encoder representations from transformers (BERT), which utilize attention models, have shown better contextual understanding compared to single-direction LMs, as demonstrated in the work of Devlin et al.
+However, bidirectional encoder representations from transformers (BERT) [60], which utilize attention models, have shown better contextual understanding compared to single-direction LMs, as demonstrated in the work of Devlin et al.
 [61].
-In terms of AM, deep learning-based models like the deep neural network-hidden Markov model (DNN-HMM) and the connectionist temporal classification (CTC) have made significant advancements.
+In terms of AM, deep learning-based models like the deep neural network-hidden Markov model (DNN-HMM) and the connectionist temporal classification (CTC) have made significant advancements [60].
 DNN-HMM models have been extensively studied in ASR research, while CTC is an end-to-end training method that does not require pre-alignment and only needs input and output sequences.
@@
 Figure 4: Overview of advanced DL-driven ASR algorithms and their commonly utilized models.
+The figures in this section are designed to fulfill two key objectives: firstly, to visually explain the principles behind advanced DL techniques, namely Transformers, DTL, FL, and DRL, as depicted in Figures 5, 8, 9, and 10, respectively; and secondly, to present practical examples of implementing widely recognized techniques.
+These include a convolutional neural network (CNN)-based Transformer illustrated in Figure 6, a particularly effective ASR application of source separation using a Transformer, highlighted in Figure 7, and a S2S-based ASR model depicted in Figure 11.
+This dual presentation provides young researchers and developers with valuable, practical insights on how to implement these advanced DL techniques in their projects.
+By doing so, it bridges the gap between theory and practice, encouraging the application and extension of these examples in new and innovative ways.
 3.1 Transformer-based ASR
@@
 The standard Transformer model incorporates the query–key–value (QKV) attention mechanism.
-In this setup, given matrix representations of queries $\mathbf{Q}\in\mathbb{R}^{N\times D_{k}}$, keys $\mathbf{K}\in\mathbb{R}^{M\times D_{k}}$, and values $\mathbf{V}\in\mathbb{R}^{M\times D_{v}}$, the scaled dot-product attention is defined as:
-$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{k}}}\right)\mathbf{V}$
+In this setup, given matrix representations of queries $\mathbf{Q}\in\mathbb{R}^{N\times D_{k}}$, keys $\mathbf{K}\in\mathbb{R}^{M\times D_{k}}$, and values $\mathbf{V}\in\mathbb{R}^{M\times D_{v}}$, the scaled dot-product attention is defined as formulated in Equation 2:
+$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{k}}}\right)\mathbf{V}$ (2)
 Here, $N$ and $M$ represent the lengths of queries and keys (or values), and $D_{k}$ and $D_{v}$ denote the dimensions of keys (or queries) and values.
@@
 Within the Transformer architecture, three attention mechanisms exist based on the source of queries and key–value pairs:
-  • Self-attention: In the Transformer encoder, queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$ all equal the outputs of the previous layer, denoted as $\mathbf{X}$ in Eq. (2).
+  • Self-attention: In the Transformer encoder, the queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) are all equal to the outputs of the previous layer, considered as $\mathbf{X}$, or to the initial embeddings in the case of the first layer.
   • Masked Self-attention: In the Transformer decoder, self-attention is constrained, allowing queries at each position to attend only to key–value pairs up to and including that position.
@@
 In the subsequent subsections, we delve into a comprehensive review and detailed analysis of several cutting-edge techniques within each of these categories.
+Table 4 summarises the most recent Transformer-based ASR techniques used in AM and LM domains.
 Table 4: Summary of some proposed work in Transformer-based ASR.
@@
 In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
-Ref.
-Based on
-Speech recognition task
-Transformer
-AM/LM
-Result with metric
-[55]
-CNN
-Solve the problem of code-switching
-Multi-head attention
-LM
-RER= 10.2%
-[56]
-VGGnet
-Compress ASR parameters and speeds up the inference time
-Low-rank multi-head attention
-AM
-CER= 13.09%
-[57]
-DNN-HMM
-Improve ASR
-Attention
-AM
-RER= 4.7$\downarrow$
-[62]
-Emformer
-Large scale ASR
-Attention
-AM
-RERR= 26%
-[63]
-TRUNet
-Sound source separation
-TNet
-AM
-PESQ= 0.22$\uparrow$
-[64]
-MHSA
-Improve speech/ASR
-[…]
-AM
-PESQ= 0.96$\uparrow$
-[58]
-HMM
-Improve ASR
-Acoustic Encoder
-AM
-Acc= 96%
-[65]
-RNN-T
-Acoustic re-scoring scenario
-Transformer- Transducer
-AM
-Acc= 8$\uparrow$
-[66]
-CTC
-ASR, ST, Acoustic ED
-All-in-one
-AM
-WER=0.3$\uparrow$
-[67]
-CNN
-Speech recognition with low latency, reduced frame rate, and streamability.
-Transformer-Transducer
-AM
-WER= 3.6%
-[68]
-CTC alignment
-Retrieve the acoustic embedding at the token level for better ASR
-Attention
-AM
+Ref. Based on Speech recognition task Transformer AM/LM Result with metric
+[55] CNN Solve the problem of code-switching Multi-head attention LM RER= 10.2%
+[56] VGGnet Compress ASR parameters and speeds up the inference time Low-rank multi-head attention AM CER= 13.09%
+[57] DNN-HMM Improve ASR Attention AM RER= 4.7$\downarrow$
+[63] Emformer Large scale ASR Attention AM RERR= 26%
+[64] TRUNet Sound source separation TNet AM PESQ= 0.22$\uparrow$
+[65] MHSA Improve speech/ASR D2Net AM PESQ= 0.96$\uparrow$
+[58] HMM Improve ASR Acoustic Encoder AM Acc= 96%
+[66] RNN-T Acoustic re-scoring scenario Transformer- Transducer AM Acc= 8$\uparrow$
+[67] CTC ASR, ST, Acoustic ED All-in-one AM WER=0.3$\uparrow$
+[68] CNN Speech recognition with low latency, reduced frame rate, and streamability. Transformer-Transducer AM WER= 3.6%
 […]
 WER= 2.3%
-[69]
-RNN-LSTM
-Improve the efficiency of end-to-end ASR
-Attention
-LM
-CER=1.98%
-[59]
-CTC Alignment
-Enhance the performance of end-to-end ASR
-Autoregressive Transformer
-AM
-RTF= 0.0134
+[70] RNN-LSTM Improve the efficiency of end-to-end ASR Attention LM CER=1.98%
+[59] CTC Alignment Enhance the performance of end-to-end ASR Autoregressive Transformer AM RTF= 0.0134
 WER= 2.7%
 3.1.1 Acoustic domain
-The study [57] reveals the Transformer model’s increased susceptibility to input sparsity compared to the convolutional neural network (CNN).
+Figure 5: Three forms of end-to-end Transformers models: (a) attention, (b) RNN-Transducer, and (c) basic CTC [71].
+The study [57] reveals the Transformer model’s increased susceptibility to input sparsity compared to the CNN.
 The authors analyze the performance decline, attributing it to the Transformer’s structural characteristics.
 Additionally, they introduce a novel regularization method to enhance the Transformer’s resilience to input sparsity.
-This method directly regulates attention weights through silence label information in forced-alignment, offering the advantage of not requiring extra module training and excessive computation.
+This method directly regulates attention weights , i,e.
+output of Figure 5 (a), through silence label information in forced-alignment, offering the advantage of not requiring extra module training and excessive computation.
 The paper [50] addresses a limitation in Transformer-based end-to-end modeling for ASR tasks, where intermediate features from multiple input streams may lack diversity.
@@
 The model incorporates FMix data augmentation, involving the creation of a binary mask from a randomly sampled complex matrix with a low-pass filter.
-SE-Trans achieves outstanding performance in ESR tasks, proven through evaluations on DCASE challenge databases, underscoring its robustness and versatility in environmental sound recognition.
-Automated audio captioning (AAC) involves generating textual descriptions for audio recordings, covering sound events, acoustic scenes, and event relationships.
-Current AAC systems typically employ an encoder-decoder architecture, with the decoder crafting captions based on extracted audio features.
+SE-Trans achieves outstanding performance in ASR tasks, proven through evaluations on DCASE challenge databases, underscoring its robustness and versatility in environmental sound recognition.
+[…]
+AAC involves generating textual descriptions for audio recordings, covering sound events, acoustic scenes, and event relationships.
+Current automated audio captioning (AAC) systems typically employ an encoder-decoder architecture, with the decoder crafting captions based on extracted audio features.
 Chen et al.
@@
 […]
-TRUNet proposed in [63] represents an innovative approach to end-to-end multi-channel reverberant sound source separation.
-The model incorporates a recurrent-U network that directly estimates multi-channel filters from input spectra, enabling the exploitation of spatial and spectro-temporal diversity in sound source separation.
-In Figure 7, the block diagram illustrates the architecture of TRUNet’s transformer.
-The transformer network (TNet) encompasses three variations: (i) TNet–Cat, which concatenates multi-channel spectra, treating them as a single input.
-This approach allows for the direct utilization of spatial information between channels.
-(ii) TNet–RealImag, utilizing two separate transformer stacks for real and imaginary parts, respectively.
-Queries and keys are computed from the multi-channel spectra.
-Despite this, the method may not fully exploit spatial information, such as phase differences, directly.
-(iii) TNet–MagPhase, analogous to TNet–RealImag, but employing spectral magnitude and spectral phase instead of real and imaginary parts.
-This variation proves superior in extracting spatial information from complex-valued spectra, resulting in maximum enhanced performance in sound source separation when employing TRUNet-MagPhase architecture.
+[…]
+In recent times, dual-path networks have demonstrated effective results in many speech processing tasks such as speech separation and SE.
+In light of this, Wang Ke and colleagues incorporated the Transformer into the structure of dual-path networks, presenting a time-domain SE model called two-stage Transformer-based neural network (TSTNN).
+This model significantly enhances the performance of SE [76].
+Some research findings suggest that the dot-product self-attention may not be essential for Transformer models.
+Similarly, the paper [65] introduces D2Net, a denoising and dereverberation network for challenging single-channel mixture speech in complex acoustic environments.
+D2Net incorporates a two-branch encoder (TBE) for feature extraction and fusion, along with a global-local dual-path Transformer (GLDPT) featuring local dense synthesizer attention to enhance local information perception.
 Figure 7: An example of source separation scheme based on Transformer [64].
-[…]
 The study proposed by Gong et al.
@@
 Shareef et al.
-in [58] propose a collaborative training method for acoustic encoders in Arabic ASR systems for speech-impaired children, achieving a 10% relative accuracy improvement on phoneme alignment in the output sequence.
+in [58] propose a collaborative training method for acoustic encoders in Arabic DTL systems for speech-impaired children, achieving a 10% relative accuracy improvement on phoneme alignment in the output sequence.
 Pioneering in recognizing impaired children’s Arabic speech.
@@
 Using co-distillation with an auxiliary task, collaborative training achieves up to 11% relative WER improvement on LibriSpeech corpus.
-Transducer models, in the context of ASR, map input sequences (acoustic features) to output sequences (transcriptions).
+Transducer models (Figure 5 (b)), in the context of ASR, map input sequences (acoustic features) to output sequences (transcriptions).
 Unlike traditional S2S models, transducers can handle variable-length input and output sequences more efficiently.
@@
 The authors claim that variable masking is the viable choice in acoustic rescoring scenarios.
-Similarly, to adapt the Transformer for streaming ASR, the authors in [67] employ the Transducer framework for streamable alignments.
+Similarly, to adapt the Transformer for streaming DTL, the authors in [68] employ the Transducer framework for streamable alignments.
 Using a unidirectional Transformer with interleaved convolution layers for audio encoding, they model future context and gradually downsample input to reduce computation cost, while limiting history context length.
 Moving on, the work [67] introduces an all-in-one AM based on the Transformer architecture, combined with the CTC to ensure a sequential arrangement and utilize timing details.
-It addresses ASR, audio tagging (AT), and acoustic ED simultaneously.
+It addresses ASR, AT, and acoustic ED simultaneously.
 The model demonstrates superior performance, showcasing its suitability for comprehensive acoustic scene transcription.
@@
 It significantly reduces parameters, boosting training speed by over 50% and inference time by 1.35$\times$ compared to baseline.
-Experiments show better generalization, lower error rates, and outperformance existing schemes without external language or acoustic models.
-Growing demand for on-device ASR systems prompts interest in model compression.compression.
+Experiments show better generalization, lower error rates, and outperformance of existing schemes without external language or acoustic models.
+Growing demand for on-device ASR systems prompts interest in model compression.
 3.1.2 Language domain
 Self-attention models, such as Transformers, excel in speech recognition and reveal an important pattern.
-As upper self-attention layers are replaced with feed-forward layers, resembling CLDNN architecture in [48], experiments on wall street journal (WSJ) and switchboard (SWBD) datasets show no performance drop and minor gains.
+As upper self-attention layers are replaced with feed-forward layers, resembling convolutional long-short term deep neural network (CLDNN) architecture in [48], experiments on wall street journal (WSJ) and switchboard (SWBD) datasets show no performance drop and minor gains.
 The novel proposed metric of attention matrix diagonality indicates increased diagonality in lower to upper encoder self-attention layers.
@@
 On the LibriSpeech benchmark dataset, significant reductions in WERs are observed, particularly on noisier evaluation dataset portions, outperforming baseline models and approaching the performance of Transformer-XL neural language model re-scoring with 6-gram.
-Figure 8: Some end-to-end models: Basic CTC , RNN-Transducer, and attention architectures [78].
 […]
-CTC is an architecture and principle commonly used in S2S tasks (Figure 8), such as ASR.
+CTC is an architecture and principle commonly used in S2S tasks (Figure 5 (c)), such as ASR.
 It enables alignment-free training by introducing a blank symbol and allowing variable-length alignments between input and output sequences.
@@
 CTC is particularly effective in tasks with variable-length outputs, making it well-suited for applications like speech recognition where the duration of spoken words may vary.
-Figure Figure 8 (a) illustrates the operational principles and components of CTC.
 This latter has been used in many ASR schemes, for example, Deng et al.
@@
 The inclusion of a CTC branch aids in the encoder’s convergence during training and considers all potential time boundaries in beam searching.
-The encoder is initiated with wav2vec2.0, and the introduction of a one-cross decoder (OCD) mitigates reliance on acoustic representations, enabling initialization with pretrained DistilGPT2 and overcoming the constraint of conditioning on acoustic features.
+The encoder is initiated with wav2vec2.0, and the introduction of a one-cross decoder mitigates reliance on acoustic representations, enabling initialization with pretrained DistilGPT2 and overcoming the constraint of conditioning on acoustic features.
 […]
@@
 The study introduces a new dataset, yielding SOTA results with a low character error rate.
-4.2 TL-based ASR
+3.2 DTL-based ASR
 Overall, DTL consists of training a DL model on a specific domain (or task) and then transferring the acquired knowledge to a new, similar domain (or task).
@@
 [86] introduce an audio pyramid Transformer with an attention tree structure, with four branches, to reduce computational complexity in fine-grained audio spectrogram processing.
-It proposes a DA transfer learning approach for weakly supervised sound event detection, a sub-field of ASR, enhancing localization performance by aligning feature distributions between frame and clip domains with a DA detection loss.
+It proposes a DA transfer learning approach for weakly supervised sound ED, a sub-field of ASR, enhancing localization performance by aligning feature distributions between frame and clip domains with a DA detection loss.
 3.2.2 Language domain
@@
 Song et al.
-[87] present a novel learning-to-rescore (L2RS) approach, which relies on two main components: (i) utilizing diverse textual data from SOTA NLP models, such as BERT, and (ii) automatically determining their weights to rescore the N-best lists for ASR systems.
+[88] present L2RS approach, which relies on two main components: (i) utilizing diverse textual data from SOTA NLP models, such as BERT, and (ii) automatically determining their weights to rescore the N-best lists for ASR systems.
 Recent advancements in S2S models have shown promising results for training monolingual ASR systems.
@@
 Additionally, joint training of these architectures in a multi-task hybrid approach has been explored, demonstrating improved overall performance.
-For instance, the architecture illustrated in Fig.
-8 (a) comprises S2S layers.
+For instance, the architecture illustrated in Figure 5 (a) comprises S2S layers.
 The encoder network of S2S consists of a series of RNNs that generate embedding vectors, while the RNN decoder utilizes these vectors to produce final results.
 The RNN also benefits from prior predictions ($P_{i},i=0,\dots,n$), enhancing the accuracy of subsequent predictions.
-Moving on, a novel TL-based approach that enhances end-to-end speech recognition has been proposed in [88].
+Moving on, a novel DTL-based approach that enhances end-to-end speech recognition has been proposed in [89].
 The novelty lies in applying DTL through multilingual training and multi-task learning at two levels.
-The initial stage utilizes nonnegative matrix factorization (NMF), instead of a bottleneck layer, and multilingual training for high-level feature extraction.
+The initial stage utilizes non-negative matrix factorization, instead of a bottleneck layer, and multilingual training for high-level feature extraction.
 The subsequent stage employs joint CTC-attention models on these features, where the CTC was transferred to the target attention-based model.
@@
 Further optimization is needed for standard end-to-end training.
-In addition, Integrating both AM and LM methodologies has the potential to enhance or construct an effective DTL-based ASR model, as demonstrated in [46, 89, 90, 91, 92, 93].
+In addition, integrating both AM and LM methodologies has the potential to enhance or construct an effective DTL-based DTL model, as demonstrated in [46, 90, 91, 92, 93, 94].
 Table 5: In contemporary cutting-edge frameworks, diverse pre-trained models are utilized for distinct tasks within the field.
@@
 In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
-Scheme
-Based on
-Speech recognition task ($\mathbb{T}_{T}$)
-AM/LM
-Adaptation
-Result with metric
-[84]
-ConvNeXt-Tiny
-Audio classification
-AM
-DA
-mAP= 0.471
-[85]
-APT
-Sound event detection
-AM
-DA
-F1= 79.6%
-[49]
-BERT (Jasper)
-Speech-to-text
-LM
-TL
-WER= 14%
-[54]
-DistilGPT2
-Improve ASR
-Both
-Fine-tuning
-CER= 4.6%
-[93]
-XLRS Wave2vec
-Improve ASR in low resource language
-Both
-Fine-tuning
-5.6% WER $\downarrow$
-[82]
-XLRS Wave2vec
-Improve ASR for children’s speech
-AM
-TL
-WER = 4.86%
-[90]
-S2S
-Speaker adaptation
-Both
-Features norm.
+Scheme Based on Speech recognition task ($\mathbb{T}_{T}$) AM/LM Adaptation Result with metric
+[85] ConvNeXt-Tiny Audio classification AM DA mAP= 0.471
+[86] APT Sound event detection AM DA F1= 79.6%
+[49] BERT (Jasper) Speech-to-text LM DTL WER= 14%
+[54] DistilGPT2 Improve ASR Both Fine-tuning CER= 4.6%
+[94] XLRS Wave2vec Improve ASR in low resource language Both Fine-tuning 5.6% WER $\downarrow$
+[83] XLRS Wave2vec Improve ASR for children’s speech AM DTL WER = 4.86%
 […]
-[83]
-Transformer
-Multi-dialect model aids recognizing diverse speech dialects effectively
-AM
-Multi-task learning
-Acc= 100%
-[94]
-S2S
-Enhancing the existing multilingual S2S model.
-LM
-DTL
-4%CER $\downarrow$ 6% WER
-[95]
-PaSST
-Audio tagging and event detection
-AM
-Fine-tuning
-F1= 64.85%
-[81]
-Wav2vec
-WSJ data speech
-AM
-affine transform.
+[84] Transformer Multi-dialect model aids recognizing diverse speech dialects effectively AM Multi-task learning Acc= 100%
+[95] S2S Enhancing the existing multilingual S2S model. LM DTL 4%CER $\downarrow$ 6% WER
+[96] PaSST Audio tagging and event detection AM Fine-tuning F1= 64.85%
 […]
-[96]
-ARoBERT
-Spoken language understanding
-LM
-Fine-tuning
-F1-score=92.56%
+[97] ARoBERT Spoken language understanding LM Fine-tuning F1-score=92.56%
 Abbreviations: Transformer (T)
@@
 The server then aggregates these updates to improve the global model.
-  • Horizontal federated fearning (HFL): In HFL, clients train a shared global model using their respective datasets, characterized by the same feature space but different sample spaces.
+  • Horizontal federated learning (HFL): In HFL, clients train a shared global model using their respective datasets, characterized by the same feature space but different sample spaces.
 Each client utilizes a local AI model, and their updates are aggregated by a central server without exposing raw data.
@@
 The objective function minimizes a global loss across all parties’ datasets [98].
-  • Vertical federated learning (VFL): VFL trains models on datasets sharing the same sample space but having different feature spaces.
+[…]
+    VFL: vertical federated learning (VFL) trains models on datasets sharing the same sample space but having different feature spaces.
 Through entity data alignment (EDA) and encrypted model trained (EMT), VFL allows clients to cooperatively train models without sharing raw data.
@@
 Guliani et al.
-[99] proposed a strategy to compensate non-independent and identically distributed (non-IID) data in federated training of ASR systems.
+[100] proposed a strategy to compensate non-identically distributed (non-IID) data in federated training of ASR systems.
 The proposed strategy involved random client data sampling, which resulted in a cost-quality trade-off.
@@
 [105] tackled a challenging and realistic ASR federated experimental setup with clients having heterogeneous data distributions, featuring thousands of different speakers, acoustic environments, and noises.
-Their empirical study focused on attention-based S2S End-to-End (E2E) ASR models, evaluating three aggregation weighting strategies: standard FedAvg, loss-based aggregation, and a novel WER-based aggregation.
+Their empirical study focused on attention-based S2S end-to-end ASR models, evaluating three aggregation weighting strategies: standard FedAvg, loss-based aggregation, and a novel WER-based aggregation.
 Table 6 summarizes the most recent FL-based ASR techniques.
@@
 In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
-Scheme
-Based on
-Speech recognition task
-FL technique
-Metric and result
-[98]
-S2S
-Improve ASR
-FedAvg
-WER = 6%
-[99]
-End-to-end RNN-T
-ASR on non-IDD data
-FedAvg
-WER=
-[100]
-CNN+Transformer extractor
-ASR on non-IDD data
-DecoupleFL
-2.3- 3.4% WER $\downarrow$ compared with FedAvg
-[101]
-LSTM
-ASR on non-IDD data
-CAFT
-WER = 15.13%
-[102]
-wav2vec 2.0
-Improve ASR
-FedAvg
-WER= 10.92%
+Scheme Based on Speech recognition task FL technique Metric and result
+[99] S2S Improve ASR FedAvg WER = 6%
+[100] End-to-end RNN-T ASR on non-IDD data FedAvg WER=
+[101] CNN+Transformer extractor ASR on non-IDD data DecoupleFL 2.3- 3.4% WER $\downarrow$ compared with FedAvg
+[102] LSTM ASR on non-IDD data CAFT WER = 15.13%
+[103] wav2vec 2.0 Improve ASR FedAvg WER= 10.92%
 EER= 5-20%
-[103]
-QCNN and RNN
-Improve privacy-preservation in ASR
-VFL
-Accuracy = 95.12%
-[104]
-S2S
-ASR on heterogeneous data distributions
-FedAvg
-WER= 19.98-23.86%
-[105]
-S2S
-ASR on private and unlabelled user data.
-FedNST
-WER= 22.5%
-[106]
-Wav2vec 2.0 and Whisper
-ASR & KWS for child exploitation settings
-FedAvg
-WER = 11-25%
-[107]
-Kaldi and backoff n-gram
-Improve privacy-preservation in ASR
-Merging models
-WER= 17.7%
-[108]
-TDNN
-Improve privacy-preservation in ASR
-Aggregation
-EER = 1-2%.
-[109]
-Non-Streaming & Streaming Conformer
-Reduce client ASR model size
-Federated Dropout
-6-22% $\downarrow$ Client size model; 34-3% WER $\downarrow$
-4.4 RL-based ASR
+[104] QCNN and RNN Improve privacy-preservation in ASR VFL Accuracy = 95.12%
+[105] S2S ASR on heterogeneous data distributions FedAvg WER= 19.98-23.86%
+[106] S2S ASR on private and unlabelled user data. FedNST WER= 22.5%
+[107] Wav2vec 2.0 and Whisper ASR & KWS for child exploitation settings FedAvg WER = 11-25%
+[108] Kaldi and backoff n-gram Improve privacy-preservation in ASR Merging models WER= 17.7%
+[109] TDNN Improve privacy-preservation in ASR Aggregation EER = 1-2%.
+[110] Non-Streaming & Streaming Conformer Reduce client ASR model size Federated Dropout 6-22% $\downarrow$ Client size model; 34-3% WER $\downarrow$
+3.4 DRL-based ASR
 […]
-RL is a ML paradigm where an agent learns optimal decision-making by interacting with an environment.
+DRL is a ML paradigm where an agent learns optimal decision-making by interacting with an environment.
 The agent receives feedback in the form of rewards or penalties, adapting its behavior to maximize cumulative reward over time through a trial-and-error process.
-reinforcement learning (RL) involves several key concepts, as defined in the following terms: Environment model, serving as a representation of contextual dynamics; State (s), denoting the current situation perceived by the agent; Observation (o), a subset of the state directly perceived by the agent; Action (a), the decision made by the agent in response to the environment; Policy ($\pi$), describing how the agent converts environmental conditions into actions; Agent, the entity making decisions based on current states and past experiences; Reward, numerical values assigned by the environment to the agent based on state-action interactions.
-Figure 11 illustrate the principle of RL.
-Figure 11: RL principle.
+DRL involves several key concepts, as defined in the following terms: Environment model, serving as a representation of contextual dynamics; State ($s$), denoting the current situation perceived by the agent; Observation ($o$), a subset of the state directly perceived by the agent; Action ($a$), the decision made by the agent in response to the environment; Policy ($\pi$), describing how the agent converts environmental conditions into actions; Agent, the entity making decisions based on current states and past experiences; Reward, numerical values assigned by the environment to the agent based on state-action interactions.
+Figure 10 illustrates the principle of DRL.
+Figure 10: DRL principle.
 Markov decision process (MDP) is a fundamental framework for dynamic and stochastic decision-making, characterized by state space $S$, action space $\mathbb{A}$, transition probabilities $\mathbb{P}$, and a reward function $R$.
-The primary objective in an MDP is to identify an optimal policy $\pi^{*}$ maximizing the expected discounted total reward over time, expressed as $\pi^{*}=\max_{\pi}\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}(s_{t},\pi(s_{t}))]$, where $\gamma$ is the discount factor.
+The primary objective in an MDP is to identify an optimal policy $\pi^{*}$ maximizing the expected discounted total reward over time, expressed as:
+$\pi^{*}=\max_{\pi}\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}(s_{t},\pi(s_{t}))],$ (4)
+where $\gamma$ is the discount factor.
 MDPs find extensive applications in addressing uncertainties in intelligent systems within dynamic wireless environments, including spectrum management, cognitive radios, and wireless security.
-In the field of ASR, RL has primarily been proposed to tackle discrepancies between training and testing phases.
+In the field of ASR, DRL has primarily been proposed to tackle discrepancies between the training and testing phases.
 Two main discrepancies leading to potential performance deterioration have been identified: 1) The conventional use of the cross-entropy criterion maximizes log-likelihood during training, while performance is assessed by WER, not log-likelihood; 2) The teacher-forcing method, which relies on ground truth during training, implies that the model has never encountered its own predictions before testing.
-RL addresses these discrepancies by bridging the gap between the training and testing phases.
-Several RL-based approaches for ASR have been proposed [110, 111, 112, 113, 114, 115, 116, 117, 118, 119].
-For example, in [110], the authors introduced a RL-based optimization method for the S2S ASR task called Self-critical sequence training (SCST).
+DRL addresses these discrepancies by bridging the gap between the training and testing phases.
+Several DRL-based approaches for ASR have been proposed [111, 112, 113, 114, 115, 116, 117, 118, 119, 120].
+For example, in [111], the authors introduced a DRL-based optimization method for the S2S ASR task called self-critical sequence training (SCST).
 This method can be conceptualized as a sequential decision model, depicted in Figure 11.
@@
 The framework is illustrated in Figure 11.
-In ASR, RL has been primarily proposed to address mismatches between training and testing phases.
-Two main mismatches leading to potential performance degradation are identified: 1) The commonly used cross-entropy criterion maximizes log-likelihood during training, while performance is evaluated by WER, not log-likelihood; 2) The teacher-forcing method, relying on ground truth during training, implies that the model has never encountered its own predictions before testing.
-RL bridges the gap between the training and testing phases, addressing these mismatches.
-Several RL-based ASR approaches have proposed [110, 111, 112, 113, 114, 115, 116, 117, 118, 119].
-For instance, In [110], the authors presented a RL-based optimization method for S2S ASR task called self-critical sequence training (SCST).
-This can be viewed as a sequential decision model as shown in Fig.
-12.
-The whole encoder-decoder neural network can be viewed as an agent.
-In each time step $t$, acoustic feature $x_{t}$ and previous $Y_{t-1}$ prediction are concatenated as current state $s_{t}$.
-The output token is the action at that will update the generated hypotheses sequence.
-After comparing it with ground truth sequence $Y^{*}$, a reward $r_{t}$ of this time step is calculated.
-SCST associates the training loss and WER using WER-related reward function, which considers the intermediate reward at each token generation step.
-Furthermore, SCST utilizes the test-time beam search algorithm to sample a set of hypotheses for reward normalization.
-As a result, the high-reward hypotheses that outperform the current test-time system are given positive weights, while the low-reward hypotheses are given negative weights.
-The entire encoder-decoder neural network is treated as an agent.
-At each time step $t$, the current state $s_{t}$ is formed by concatenating the acoustic feature $x_{t}$ and the previous prediction $Y_{t-1}$.
-The output token serves as the action, updating the generated hypotheses sequence.
-SCST associates training loss and WER using a WER-related reward function, calculating the reward $r_{t}$ at each token generation step by comparing it with the ground truth sequence $Y^{*}$.
-SCST uses the test-time beam search algorithm to sample hypotheses for reward normalization, assigning positive weights to high-reward hypotheses that outperform the current test-time system and negative weights to low-reward hypotheses.
-The framework is illustrated in Figure 12.
-Figure 12: SCST sequential decision model of ASR [110].
-In [111], the authors developed a RL framework for speech recognition systems using the policy gradient method.
-They introduced a RL method within this framework, incorporating user feedback through hypothesis selection.
+Figure 11: Example of sequential decision model of DRL-based ASR [111].
+In [112], the authors developed a DRL framework for speech recognition systems using the policy gradient method.
+They introduced a DRL method within this framework, incorporating user feedback through hypothesis selection.
 Tjandra et al.
-[112, 113] also employed policy gradient RL to train a S2S ASR model.
-In [114], the authors constructed a generic RL-based AutoML system.
+[113, 114] also employed policy gradient DRL to train a S2S ASR model.
+In [115], the authors constructed a generic DRL-based AutoML system.
 This system automatically optimizes per-layer compression ratios for a SOTA attention-based end-to-end ASR model, which consists of multiple LSTM layers.
@@
 Shen et al.
-[116] suggested employing RL to optimize a speech enhancement model based on recognition results, aiming to directly enhance ASR performance.
+[117] suggested employing DRL to optimize a SE model based on recognition results, aiming to directly enhance ASR performance.
 AutoML-based low-rank factorization (LRF) achieves up to 3.7× speedup.
 In the shade of this, Mehrotra et al.
-in their work [115] propose an iterative AutoML-based LRF that employs RL for the iterative search, surpassing 5× compression without degrading WERs, advancing ASR.
-Table 8 summarizes the recent RL-based ASR techniques.
-Table 8: Summary of recent proposed works in RL-based ASR.
+in their work [116] propose an iterative AutoML-based LRF that employs DRL for the iterative search, surpassing 5× compression without degrading WERs, advancing ASR.
+Table 7 summarizes the recent DRL-based ASR techniques.
+Table 7: Summary of recent proposed works in DRL-based ASR.
 The symbol ($\uparrow$) result increase, whereas ($\downarrow$) signifies result decrease.
 In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
-Scheme
-Model-based
-ASR Tasks
-RL technique
-Metric and result
-[110]
-S2S Conformer
-Improve ASR
-Policy gradient
-8.7%-7.8% WER $\downarrow$ over Baseline model
-[111]
-DNN-HMM
-Improve ASR for AM
-Policy gradient
-WER=23.82-25.43%
-[112]
-S2S
-Improve ASR FOR am
-Policy gradient
-CER=6.10%
-[113]
-S2S
-Improve ASR for AM
-Policy gradient
-CER=6.10%
-[114]
-End-to-end encoder-attention- decoder
-Improve ASR Model compression for AM
-Policy gradient
-Up to $\sim$3x compression; WER=8.06%
-[115]
-End-to-end encoder-attention- decoder
-Improve ASR Model compression for AM
-Policy gradient
-Up to $\sim$5x compression; WER=8.19%
-[116]
-CD-DNN-HMM AM & SRI LM
-Speech enhancement for ASR for AM and LM
-Q-learning
-12.40% and 19.23% CER $\downarrow$ at 5 and 0 dB SNR conditions.
-[117]
-LSTM
-Improve ASR for dialogue state tracking
-DQN
-Acc= 3.1$\uparrow$
-[118]
-S2S
-Improve ASR for AM
-Policy gradient
-CER=8.7%
-[119]
-Wav2vec 2.0
-Improve ASR for AM
-Policy gradient
-4% WER $\downarrow$
-DTL for acoustic models (AMs) Two types of deep learning-based AMs commonly used in ASR are the end-to-end model and the layered deep neural network-hidden Markov model (DNN-HMM) model.
-The DNN component is responsible for extracting high-level features, such as MFCCs and HMM lexical sequences, from acoustic signals.
-These features are then decoded into transcripts.
-The DNN takes acoustic characteristics as input and produces context-dependent lexical units (tri-phonemes) that are used as input for the downstream HMM component.
-In contrast, the end-to-end model is a DNN-based technique that directly takes acoustic characteristics (features) as input and outputs the recognition rate without the need for separate components like the DNN and HMM.
-A typical end-to-end framework for voice recognition is illustrated in Figure 13.
-The neural network in this framework generates embeddings from the input features, which are then passed through a stack of recurrent layers.
-These recurrent layers analyze patterns based on prior and current input information to produce the final output.
-The network is trained using back-propagation with the CTC loss [mridha2021study].
-Regarding domain transfer learning (DTL), three main strategies have been commonly employed with AMs.
-Figure 13: An example of end-to-end source model for DTL-based ASR [mridha2021study].
-4.4.1 Feature normalisation based-DTL
-The concept behind linear transformation is to normalize speech characteristics through a linear mapping process.
-This involves adding a transformation network or layer to an existing network.
-The transformation network is an effective method for adapting neural networks.
-Typically, the last hidden layer is designed to act as a bottleneck, reducing the number of parameters to be adjusted by using fewer neurons.
-Figure 14 illustrates this architecture.
-The transformation network can be either a linear input network or a linear output network.
-Figure 14: The basics of transformation in DTL involve estimating the weights connected to the links within the dashed rectangles, while leaving the remaining weights unchanged.
-In the case of mono-task DTL, feature normalization is applied as shown in Figure (a).
-For multi-task DTL, feature normalization is also performed, as illustrated in Figure (b).
-Assuming that the last hidden network (LHN) functions as a feature extractor and the output layer serves as a discriminative source model ($M_{S}$), the weights of the linear transformation matrix $W_{L}$ in the output layer correspond to the parameters of the target model, $M_{T}$.
-The representation of $M_{T}$ can be expressed as follows [huang2015maximum]:
-$\mathrm{M_{T}=softmax(W_{L}M_{S})}.$ (4)
-The activation of the last hidden layer in $M_{S}$ can serve as the extracted feature representation for the hidden layers in $M_{T}$.
-Transforming the model parameters using a transformation matrix $W_{LHN}$ to generate an adapted set of model parameters is equivalent to incorporating an extended last hidden layer after the existing last hidden layer [huang2015maximum]:
-$\mathrm{M_{T}=softmax(W_{LHN}W_{L}M_{S})}.$ (5)
-Several ASR research studies have utilized linear transformation strategies to improve performance.
-In one example [elaraby2016deep], the authors aimed to enhance a computer-aided language learner (CAPL) system for teaching Arabic pronunciation for Quran recitation regulations.
-They implemented various improvements, including speaker adaptive training (SAT), integrating a hybrid DNN-HMM model, combining the hybrid DNN with minimum phone error (MPE), and using a grammar-based decoding graph during testing.
-Another study [mimura2016joint] employed a multi-target learning approach to optimize the output of a denoising auto-encoder (DAE) and the input of a DNN.
-The DAE was trained in the first stage to reduce input-related errors propagated to the DNN.
-Then, a unified network of DAE and DNN was fine-tuned for phone state ASR, with an additional target of input voice augmentation applied to the DAE.
-Additionally, in [ma2017approaches], an adaptation layer was introduced for fine-tuning, and non-linearity was incorporated to learn a more complex function than a linear transformation in the softmax layer.
-During fine-tuning, adjustments were made to the cluster softmax and NNadapt layers while keeping the other layers unchanged.
-Table 6 provides a summary of additional schemes that employ linear transformation techniques, along with their performance details.
-4.4.2 Conservative training for DTL
-The conservative training approach is commonly used for accent adaptation in ASR.
-It is an efficient method that requires only a limited amount of spoken data to achieve satisfactory results.
-However, it can lead to an excessive number of parameters, which may disrupt the model’s structure.
-To address this, KL-Divergence (KLD) is widely employed as a DNN-based DTL algorithm for ASR.
-KLD regularization provides a mathematical framework for DL training, aiming to minimize the loss and make the output distributions of the source model $M_{S}$ and target model $M_{T}$ more similar.
-The KL-divergence prevents overfitting and helps to keep the adapted $M_{T}$ close to the domain of $M_{S}$ [89, 90].
-In terms of model-based DTL with KLD-regularization, assuming the loss functions for training DNNs in the source and target domains are $\mathbb{D}_{S}$ and $\mathbb{D}_{T}$, respectively, the conservative approach can be summarized mathematically as follows [89, 90]:
-$\mathbb{D}_{T}^{KLD}=(1-\rho)\mathbb{D}_{S}+\frac{\rho}{N}\sum P_{S}(x_{T}/x_{T})logP_{T}(y_{T}/x_{T}),$ (6)
-In the context of DTL-based ASR, the conservative training approach with KLD-regularization is utilized to build a target AM.
-This approach involves training DNNs on speech samples collected from the target domain $\mathbb{D}_{T}$, denoted as ($x_{T},y_{T}$), where $N$ represents the number of speech samples in $\mathbb{D}_{T}$.
-The hyper-parameter $\rho$ controls the transfer ratio from the source domain $\mathbb{D}_{S}$.
-An example of applying this DTL-based ASR approach with conservative training and KLD-regularization can be found in [89] for building a target AM.
-Additionally, in [90], both KLD and LHN techniques were employed for speaker adaptation using a pre-trained seq2seq ASR model.
-Table 6 summarizes other techniques that employ conservative training along with their corresponding performance details.
-4.4.3 Subspace-based DTL
-Subspace-based DTL approaches in ASR utilize unsupervised techniques such as PCA, SVD, and NMF for data dimensionality reduction.
-These techniques aim to identify a subspace of model parameters or transformations that capture essential information.
-PCA maps high-dimensional data to lower-dimensional subspaces while preserving correlation and maximizing variance.
-SVD, similar to PCA, condenses networks by selecting a specific number of singular values.
-However, SVD introduces nonlinearity in the weight matrix, which can lead to information loss when constructing a linear projection layer.
-NMF algorithms require at least one nonnegative matrix and express the target matrix as a weighted sum of base matrix columns, making it more interpretable than SVD and PCA.
-In the study [88], Convex Nonnegative Matrix Factorization (CNMF), a variant of NMF, is employed to extract high-level features.
-Subsequently, DTL is applied to both high and low-level features through multilingual training and multi-task learning.
-These techniques yield significant performance improvements compared to state-of-the-art ASR studies.
-Other studies utilizing subspace techniques and their respective performance details can be found in Table 6.
-LDA-based DTL In the case of discrete data collection, generative probabilistic models, such as the latent Dirichlet allocation (LDA), are employed.
-Typically, LDA is a hierarchical Bayesian model with three levels, where each item in a collection is represented as a finite mixture over a set of underlying topics.
-Each topic, in turn, is modeled as an infinite mixture of topic probabilities.
-To capture the relationship between words and create language models for specific documents, topic model-based techniques, such as LDA, have been utilized as described in [song2019topic].
-In the work presented in [hentschel2018feature], LDA features are transformed using a linear layer consisting of a weight matrix and a bias vector.
-These transformed features are then utilized as inputs in the LHN (Local Hidden Node) during the training and evaluation of the network.
-4.4.4 NNLM-based DTL
-Neural network language models (NN-LMs) generally outperform count-based LM models in automatic speech recognition (ASR) across various tasks.
-Specifically, when applied to N-best rescoring, NN-LMs achieve WER as mentioned in [hentschel2018feature].
-Adapting NN-LMs to new domains poses a research challenge, and current approaches can be categorized as either model-based or feature-based adaptations.
-In feature-based adaptation, auxiliary features are incorporated into the input of an NN-LM, while model-based adaptation involves fine-tuning and adapting network layers.
-The authors of [91] propose a recurrent neural network-based LM (RNN-LM) approach where both types of adaptation are explored.
-As an illustrative example, Figures 15 (a) and (b) provide detailed explanations of the adopted RNN-based DTL in the study.
-Furthermore, the authors in [92] suggest a DNN-based model for LM modification in ASR, employing a factorized time-delay neural network (TDNN-F).
-Specifically, the TDNN-F model is trained using a combination of cross-entropy and lattice-free maximum mutual information objective functions (LF-MMI).
-The effectiveness of TDNN-F is demonstrated in the recognition of English child speech [92].
-Figure 15: DTL-based RNNLM with different adaptation techniques, where the out-of-vocabulary (OOV) node represents an input word that does not belong to the specified vocabulary but can be included in the input.
-Similarly, out-of-shortlist (OOS) nodes can also be included in the output: (a) RNNLM with LHN adaptation layer.
-(b) RNNLM with feature-based adaptation layer.
-LSTM-based DTL Generally, the NNLM models used in ASR are still trained on a sentence-level corpus, despite the attempts to train them at the document level.
-This is due to various factors; for example, a more extended context may not be relevant for enhancing next-word prediction in conventional ASR systems.
-It is also challenging to gather training data representing extended session contexts in many conversational circumstances.
-Long-span models are becoming more common in scenarios where they are beneficial.
-Long-span models will likely help scenarios, such as transcriptions of conversations and meetings, human-to-human communication, and document production by voice [parthasarathy2019long].
-LSTM models are widely employed, and their architectures are well-suited to variable-length sequences.
-Therefore, they can exploit extreme long-range dependencies without using n-gram approximation.
-For instance, by employing equal context, the authors in [tuske2018investigation] demonstrate that the deep 4-gram LSTM outperforms big interpolated count models by performing considerably better backing off and smoothing.
-In another example, the central part of a shared encoder is constructed using BLSTM [94].
-Cross-domain ASR
-4.4.5 Cross-language DTL
-Cross-language DTL is a particular application of cross-modality DTL.
-It serves as a commonly employed method for constructing ASR models for low resource languages by leveraging models trained on other languages.
-This approach is based on the assumption that phoneme features can be shared across languages.
-Additionally, a generic ASR model can be adapted to a specific narrow domain using DTL techniques.
-To address the issue of data sparsity, various knowledge transfer methods are explored in [liu2019investigation] with the assistance of high-resource languages.
-These methods include DTL and fine-tuning, where a well-trained neural network initializes the parameters of the LHN.
-Progressive neural networks (Prognets) are also examined, as they are resistant to the forgetting effect and excel at knowledge transfer due to the presence of lateral connections in the network architecture.
-Furthermore, the utilization of cross-lingual DNNs is explored, where bottleneck features are extracted to enhance the effectiveness of the ASR system.
-Recent approaches related to ASR using cross-language DTL and their respective performances are summarized in Tables 6 and 9.
-Table 9: A summary of the recent ASR-based cross-language DTL technique.
-Whereas the marks ($\uparrow$) and ($\downarrow$) indicate improvement and reduction, respectively.
-If many scenarios has been conducted in one metric, only the best result is mentioned.
-Scheme
-Model-based
-ASR Tasks ($\mathbb{T}_{T}$)
-Characteristic
-Performance
-[yusuf2019low]
-EDML
-Framework of performing the DTL that reduces the impact of the prevalence of out-of vocabulary terms.
-query-by-example task
-[…]
-[liu2019investigation]
-Prognets
-Improving ASR scheme quality by overcoming the data sparsity problems by means of high-resource languages.
-Fine-tuning
-LHN adapt.
-[…]
-[feng2019low]
-FNN and CNN architectures
-Indo-European speech samples used to improve the identification of African languages.
-Using PLP coeff.
-Fine-tuning
-[…]
-[sahraeian2018cross]
-DNN
-Weighted averaging schemes are used to combine the ensemble’s constituents, with the combination weights being trained to minimize the cross-entropy objective function.
-Weights interpolation
-[…]
-[wilkinson2020semi]
-CNN-GMM-HMM
-Fully-automatic segmentation, semi-supervised training of ASR systems for five-lingual code-switched speech
-Semi-supervised
-[…]
-4.4.6 DTL-based ASR for emotion recognition:
-Combining language information with acoustic data has been shown to enhance the accuracy of speech emotion recognition (SER).
-Therefore, integrating both systems can be advantageous to enhance the capability of ASR systems to handle emotional speech and provide linguistic input to SER systems.
-Figure 16 illustrates a hybrid ASR-SER system [fayek2016deep], where a spectrogram is inputted into shared convolutional layers, followed by specialized layers that have shared levels to facilitate interaction between the two systems.
-This integration allows for improved performance in handling emotional speech and leveraging linguistic features.
-Figure 16: A linguistic-paralinguistic hybrid system [fayek2016deep]. (a) ASR-SER Hybrid system. (b) An example of DTL-based ASR-SER hybrid system.
-The approach presented in [tits2018asr] utilizes the internal representation of a speech-to-text system to explore the connection between valence/arousal and different modalities through the use of DTL.
-A speech-to-text or ASR system learns to map audio speech signals to their corresponding transcriptions.
-By employing DTL, the proposed method can estimate valence and arousal using features learned from an ASR task.
-This approach offers the advantage of combining large datasets of speech with transcriptions with smaller datasets annotated with emotional dimensions.
-In a similar vein, the work described in [ananthram2020multi] fine-tunes a TDNN-based speaker recognition model for the task of emotion detection using the crema-D multi-modal emotion dataset and canonical label clustering.
-By adapting the model using fine-tuning, the authors aim to improve its performance on emotion detection.
-The study presented in [boateng2020speech] focuses on extracting features from audio segments with extreme positive and negative ratings, as well as the ending of the audio.
-They employ the peak-end rule and a DTL approach to extract acoustic features.
-The authors utilize a pre-trained CNN speech model called YAMNet and a linear SVM for binary classification of partner valence.
-4.4.7 Cross-corpus SER (CC-SER)
-In the context of Speech Emotion Recognition (SER), it is typically assumed that speech utterances in the training and testing domains are recorded under the same conditions.
-However, in real-world scenarios, speech data is often collected from diverse environments or devices, resulting in a domain discrepancy that adversely affects recognition performance.
-As a solution, researchers have recently investigated the problem of cross-corpus SER (CC-SER) and explored various DTL models.
-For example, [song2019transfer] proposes a transfer linear subspace learning (TLSL) scheme to develop a CC-SER framework.
-This approach facilitates the learning of shared feature space for both the source and target domains.
-The similarity between different corpora is estimated using a nearest-neighbor graph algorithm.
-Additionally, a feature grouping method is devised to partition emotional features into highly transferable parts (HTP) and low transferable parts (LTP).
-In the case of unsupervised CC-SER examined in [liu2018unsupervised], only the training data is annotated.
-The authors introduce a domain-adaptive subspace learning (DoSL) technique to learn a projection matrix that transforms the source and target speech data from the initial domain to the labeled domain.
-This allows the classifier trained on the labeled source domain data to effectively predict the emotional states of the unlabeled target domain data.
-Furthermore, [liu2021transfer] presents an improvement to the DoSL-based CC-SER method by introducing transfer subspace learning (TRaSL).
-In another study by [luo2019cross], a semi-supervised CC-SER approach is proposed using non-negative matrix factorization (NMF).
-This approach incorporates training corpus labels into NMF and seeks a latent low-rank feature space where the differences in conditional and marginal distributions between the two corpora can be simultaneously minimized.
-Advancing further, [zhang2019transfer] proposes a transfer sparse discriminant subspace learning (TSDSL) method to discover a shared feature subspace among multiple corpora.
-This is achieved by incorporating the $\ell_{2,1}$-norm penalty and discriminative learning, facilitating the identification of the most discriminative characteristics across the corpora.
-Similarly, [luo2020nonnegative] introduces a non-negative matrix factorization-based transfer subspace learning (NMFTSL) scheme.
-The goal is to minimize the distances between the marginal and conditional distributions in the common subspace.
-The distances are estimated using the maximum mean discrepancy (MMD) criterion.
-In [zhang2021cross], a joint transfer subspace learning and regression (JTSLR) technique is employed.
-It learns a latent subspace using discriminative MMD as the discrepancy metric, followed by modeling the relationships between features and annotations using a regression function in the latent subspace.
-A label graph is utilized to enhance knowledge transfer from the source domain (SD) data to the target domain (TD) data.
-Similarly, [chen2019target] presents a target-adapted subspace learning (TaSL) approach for CC-SER.
-It aims to find a projection subspace that enables more accurate regression of labels from the features.
-This effectively bridges the gap in feature distributions between the TD and SD.
-The projection matrix is optimized by combining $\ell_{1}$-norm and $\ell_{2,1}$-norm penalty terms with other regularization terms.
-Furthermore, [zhao2021cross] employs sparse subspace transfer learning (SSTL) to develop a CC-SER technique.
-It learns a robust common subspace projection using discriminative subspace learning and transfers knowledge from the source corpus to the target corpus through sparse reconstruction based on $\ell_{2,1}$-norm.
-The target samples are suitably represented as linear combinations of the SD data.
-On a different note, [braunschweiler2021study] investigates the impact of cross-corpus data complementation and data augmentation on the performance of SER models.
-The study focuses on six emotional speech corpora, considering factors such as single and multiple speakers, as well as variations in emotion style (natural, elicited, and acted).
-4.5 Adversarial TL-based ASR
-In many cases, the source model is trained using multilingual training, where abundant speech data is available in multiple languages [88, 94].
-Multilingual training involves shared hidden layers (SHL) and language-specific layers or classifier layers for different languages.
-The SHL of the source model acts as a feature converter, transforming language-specific features into a common feature space [liu2019investigation].
-However, there might be language-dependent features present in the common feature space, which hinders effective cross-lingual knowledge transfer.
-To address this issue, language-adversarial training is employed.
-Adversarial training helps in creating a language-invariant feature space.
-Once the source model is prepared, the first $n$ SHL can be transferred to the target model of an unknown language.
-In [yi2018language], the authors propose language-adversarial transfer learning as a solution to mitigate the performance degradation of the target model caused by shared features that may contain unnecessary language-dependent information.
-Fig.
-17 illustrates the architecture of the suggested language-adversarial transfer learning method [yi2018language].
-The source model, also known as the adversarial SHL model, is shown on the left, while the target model is depicted on the right.
-The presence of an additional language discriminator in the SHL model is denoted.
-The fully connected layer is represented as FC.
-The gradient reversal layer (GRL) ensures that the feature distributions across all languages are made as similar as possible for the language discriminator.
-The output labels of the language discriminator indicate the languages.
-Figure 17: An example of proposed model architecture language-adversarial TL for limited ASR resource [yi2018language].
-Senones refers to feature cluster’s name, representing similar acoustic states/events.
-To enhance the effectiveness of ASR in low-resource scenarios, a combination of semi-supervised training and language adversarial TL is explored in [kumar2021exploration].
-The research presented in [yi2020adversarial] suggests utilizing adversarial transfer learning to improve punctuation prediction performance.
-Specifically, a pre-trained BERT model is employed to transfer bidirectional representations to punctuation prediction models.
-The proposed approach is applied to the ASR task as the target task.
-Table 10 provides a summary of the performance achieved in recent studies involving adversarial TL for ASR.
-Table 10: A summary of the recent ASR-based adversarial-language TL technique.
-Whereas the marks ($\uparrow$) and ($\downarrow$) indicate improvement and reduction, respectively.
-If many scenarios has been conducted in one metric, only the best result is mentioned.
-Scheme
-Model-based
-ASR Tasks ($\mathbb{T}_{T}$)
-Characteristic
-Performance
-[kumar2021exploration]
-Adversarial SHL-Mode
-( SincNet-CNN-LiGRU )
-Use three Indian languages ( Hindi, Marathi, and Bengali ) cross-lingual to improve Hindi ASR
-Semi-supervised
-WER=5.5%
-(25.65% WER $\downarrow$)
-[yi2018language]
-SHL Model
-Improve the performance of low-resource ASR
-Cross-lingual
-10.1 % WER $\downarrow$
-[yi2020adversarial]
-BERT
-Improve the performance of punctuation predicting
-Multi-task
-9.4% F1-score $\uparrow$
-4.6 DTL-based ASR for medical diagnosis
-DTL-based ASR has made significant advancements in the field of medicine, particularly in the early detection of diseases.
-These advancements have been observed in various medical domains, as outlined in Table 11, which includes:
-4.6.1 Heart sound classification
-The components of the heart sound encompass various elements.
-The first (S1) and second (S2) heart sounds are considered normal, whereas the third (S3) and fourth (S4) heart sounds are often associated with murmurs, and ejection clicks typically indicate certain illnesses or abnormalities.
-Koike et al.
-[koike2020audio] introduced a novel DTL approach based on Probabilistic Audio Neural Networks (PANNs), which involves pre-training a model on a large-scale audio dataset for the purpose of classifying heart sounds.
-Another approach for heart sound classification was proposed by Boulares et al.
-[boulares2020transfer].
-Their method utilizes DTL on the Pascal public dataset, without any denoising or cleaning procedures, to establish an experimental benchmark.
-The primary objective is to provide a foundation of experimental results that can serve as a starting point for future research on cardiovascular disease (CVD) recognition using phonocardiogram (PCG)-based cardiac cycle vibration sounds.
-This proposed scheme addresses the absence of a CVD recognition benchmark and the lack of objective comparability among classification results, which tend to vary significantly.
-4.6.2 Parkinson disease detection
-Parkinson’s disease (PD) is a progressive neurodegenerative condition with a global impact.
-Accurate diagnosis of PD is crucial for improving daily activities and extending patients’ lives.
-However, predicting symptom changes and their impact on patients’ lives is challenging due to the variability in symptoms and disease progression.
-Traditional PD detection methods are often manual and require specialized expertise.
-Language impairment and atrophy are common early symptoms in over 90% of PD patients, characterized by reduced voice volume, monotonous and rapid speech, and eventual loss of audibility.
-Karaman et al.
-[43] proposed a robust automated PD detection approach using DTL-based ASR, where pre-trained models like SqueezeNet1_1, ResNet101, and DenseNet161 were fine-tuned and retrained.
-This scheme showed promising results in PD detection.
-To address the scarcity of speech data for PD and the distribution inconsistency among subjects, Li et al.
-[42] introduced a two-step unsupervised DTL algorithm called two-step sparse transfer learning (TSTL).
-This algorithm helps extract useful information from large amounts of unlabeled speech data, align the distribution of training and test sets, and preserve the original sample structure simultaneously.
-Another strategy proposed by Qing et al.
-[45] involves enhancing ASR for Parkinson’s patients using a pre-trained long short-term memory (LSTM) neural network model.
-To mitigate overfitting and reduce WER, the scheme employs frequency spectrogram masking data augmentation.
-4.6.3 Other medical diagnosis
-In the field of speech-based depression prediction, Harati et al.
-[harati2021speech] proposed a DTL method that utilizes a lightweight encoder and transfers only the encoder weights.
-This approach simplifies the runtime model for depression prediction.
-For individuals with dysarthria, Xiong et al.
-[xiong2020source] developed an improved DTL framework for robust personalized speech recognition models.
-They adapted the CNN-TDNN-F ASR AM onto the target dysarthric speakers using neural network weight adaptation.
-In the context of dysarthria speaking identification, Takashima et al.
-[takashima2019knowledge] proposed a method that transfers two types of knowledge from different datasets: the language-dependent characteristics of unimpaired speech and the language-independent characteristics of dysarthric speech.
-They focused on Japanese people with articulation disorders.
-Additionally, Sertolli et al.
-[sertolli2021representation] presented a novel feature representation for health states identification using an end-to-end DTL-based ASR framework.
-They utilized ASR DNNs as feature extractors, combined multiple feature representations using compact bilinear pooling (CBP), and employed an optimized RNN classifier for inference.
-Table 11: A summary of a DTL-based ASR technique in medical diagnosis, whereas the marks ($\uparrow$) and ($\downarrow$) indicate the improvement and reduction, respectively.
-If many scenarios have been conducted in one metric, only the best result is mentioned.
-Scheme
-Model-based
-ASR Tasks ($\mathbb{T}_{T}$) $\uparrow$
-DTL Type
-Performance
-[43]
-DenseNet-161
-PD detection
-Fine-tuning
-Accuracy= 91.17%
-[42]
-TSTL-based CSC&SF
-PD speech diagnosis
-Unsupervised
-Accuracy= 97.50%
-[45]
-Proposed four layers
-PD speech
-Fine-tuning
-[…]
-[koike2020audio]
-PANN CNN14
-Heart sound classification
-Fine-tuning
-UAR= 89.7%
-[boulares2020transfer]
-InceptionResNet-v2
-PCG-based CVD classification
-Fine-tuning
-Accuracy= 0.89%
-[harati2021speech]
-EH-AC
-Depression prediction
-LHN (encoder weights)
-27% AUC $\uparrow$
-[xiong2020source]
-CNN-TDNN-F
-Dysarthric speech
-Neural weight adapter
-[…]
-[takashima2019knowledge]
-LAS
-Dysarthric Speech
-Multilingual
-[…]
-[sertolli2021representation]
-Wav2Letter and
-DeepSpeech
-Health states classification
-Transductive
-UAR= 73.0% (8.6% $\uparrow$)
-4.7 DTL-based ASR attacks and security
-Adversarial examples are created by introducing small perturbations or noise to valid audio files or speech characteristics in order to manipulate or deceive ASR systems.
-These perturbations, although often imperceptible to the human auditory system and may be perceived as background noise by the ASR model, can lead to correct or incorrect classification of the inputs.
-For instance, by introducing a slight disturbance to the speech "At the still point, there the dance is," an ASR system may produce the incorrect transcription "At the tail point, there the tense is" [hu2019adversarial].
-The concept behind attacking ASR systems lies in the vulnerability of these models to adversarial examples, which has prompted speech researchers to explore the creation of such examples.
-By generating adversarial examples for different representations of speech in the time or frequency domain, which capture various speech features and can be used as inputs to neural networks, researchers can manipulate ASR performance, either improving it or decreasing it.
-DTL plays a key role in achieving transferability, whereby adversarial examples crafted to attack a source model can also be effective in attacking target models that classify the same type of data.
-In this context, adversarial attacks on DTL-based ASR models can be categorized into two groups, as illustrated in Fig.
-18.
-Figure 18: Possible adversarial attacks in DTL-based ASR schemes.
-4.7.1 Positive adversarial attacks
-Encompassing various techniques that aim to enhance or ensure the efficacy of current ASR systems, these methods strive to improve the robustness and performance of acoustic models.
-In [sun2018training], the authors introduced a methodology that combines natural data with adversarial data to train a resilient acoustic model.
-Specifically, they focused on utilizing MFCC features and employed a gradient-based approach to generate adversarial MFCC features, taking into account the network model and input parameters for each mini-batch.
-By applying the teacher/student training concept, the neural network was trained using a combination of natural data and the generated adversarial data.
-The effectiveness of the proposed approach was validated through experiments conducted on the CHiME-4 and Aurora-4 tasks, utilizing a customized convolutional neural network (CNN).
-4.7.2 Negative adversarial attacks
-Encompassing various techniques that aim to diminish or undermine the effectiveness of existing ASR systems, these methods pose potential threats to the performance and robustness of acoustic models.
-In [abdullah2021hear], the authors propose a framework that utilizes the Google (Phone) model as the source model and investigates the impact of adversarial attacks on the target model (DeepSpeech 1).
-The adversarial attacks are applied to the audio waveform, specifically after the signal decomposition and thresholding processes, and the resulting manipulated input is fed into the ASR source model.
-According to [hu2019adversarial], the adversarial attack models can be categorized into two types based on the adversary’s objectives, knowledge, and background, as illustrated in Figure 18:
-  • Adversary knowledge: It is divided into two types: white-box attack, which assumes the adversary has complete knowledge of $M_{T}$ including its architecture, training weights, and parameters; and black-box attack, which assumes the adversary has no access to $M_{T}$ and only knows its output like a regular user.
-  • Adversarial specificity: It is divided into two types: non-targeted attack, which aims to make the adversarial example’s $M_{T}$ predict any incorrect class with the sole goal of compromising the ASR algorithm, and targeted attack, which aims to deceive $M_{T}$ into assigning the adversarial example to a specific class selected by the attacker, imposing specific instructions on the ASR scheme.
-Moving on, Table 12 shows a summary of DTL-based adversarial models for existing works.
-Table 12: A summary of the recent DTL-based ASR for adversarial attacks. Whereas the marks ($\uparrow$)and ($\downarrow$), indicate improvement and reduction respectively.
-Scheme
-Model-based
-ASR Tasks ($\mathbb{T}_{T}$)
-Target object
-Adversary knowledge
-Adversarial specificity
-Performance
-[sun2018training]
-Aurora-4
-Boost
-MFCC
-White-box
-Targeted
-23% WER $\downarrow$
-[schonherr2018adversarial]
-DNN-HMM (Kaldi)
-Boost
-Waveform
-White-box
-Targeted
-Accuracy= 98 %
-[zelasko2021adversarial]
-DeepSpeech
-Fool
-Waveform
-White-box
-Targeted
-[…]
-[subramanian2020study]
-VGG13
-Fool (Dense_mel)
-Mel-spectrogram
-White-box
-Non-targeted
-SNR=29.06 dB
-[carlini2018audio]
-DeepSpeech
-Fool (Speech-to-Text)
-Waveform
-White-box
-Targeted
-Attack success rate= 100%
-[abdullah2021hear]
-Google
-Fool
-(Deep-Speech 1)
-Waveform
-Black-box
-Non-targeted
-Attack success rate=87%
-[kwon2019selective]
-DeepSpeech
-Fool (victime DeepSpeech)
-Mel-frequency cepstrum
-White-box
-Targeted
-Attack success rate=91.67%)
-4.8 DTL-based ASR for other applications
-DTL-based ASR has been applied in various fields.
-In the work by Boes et al.
-[12], DTL is explored for audio tagging and sound event detection tasks, where pre-trained auditory and visual features are incorporated into a baseline system using feature fusion.
-Arora et al.
-[arora2017study] use DTL to address the lack of annotated databases for audio event detection.
-Wang et al.
-[wang2020cross] propose an environment adaptation technique using DTL for deep speech enhancement models.
-Chen et al.
-[chen2018transfer] employ DTL for wearable devices to evaluate social speech in natural daily situations.
-Wu et al.
-[wu2020self] investigate self-supervised pre-trained speech for speech translation, while Zhu et al.
-[zhu2021conwst] propose a self-supervised bidirectional distillation system for low-resource speech translation.
-DTL is used in the speaker verification field by Hong et al.
-[hong2017transfer] for discriminative learning, and in the detection of marine mammal sounds by Lu et al.
-[lu2021detection] to classify different species using DTL with the AlexNet pre-trained model.
-Lastly, Azizah et al.
-[azizah2020hierarchical] employ hierarchical DTL for multilingual text-to-speech synthesis using DNNs for low-resource languages.
-5 Open Issues and of Key challenges
-Integrating advanced techniques like DTL, FL, and RL into ASR systems presents exciting opportunities but comes with its set of challenges.
+Scheme Model-based ASR Tasks DRL technique Metric and result
+[111] S2S conformer Improve ASR Policy gradient 8.7%-7.8% WER $\downarrow$ over Baseline model
+[112] DNN-HMM Improve ASR for AM Policy gradient WER=23.82-25.43%
+[113] S2S Improve ASR for AM Policy gradient CER=6.10%
+[114] S2S Improve ASR for AM Policy gradient CER=6.10%
+[115] End-to-end encoder-attention- decoder Improve ASR Model compression for AM Policy gradient Up to $\sim$3x compression; WER=8.06%
+[116] End-to-end encoder-attention- decoder Improve ASR Model compression for AM Policy gradient Up to $\sim$5x compression; WER=8.19%
+[117] CD-DNN-HMM AM & SRI LM Speech enhancement for ASR for AM and LM Q-learning 12.40% and 19.23% CER $\downarrow$ at 5 and 0 dB SNR conditions.
+[118] LSTM Improve ASR for dialogue state tracking DQN Acc= 3.1$\uparrow$
+[119] S2S Improve ASR for AM Policy gradient CER=8.7%
+[120] Wav2vec 2.0 Improve ASR for AM Policy gradient 4% WER $\downarrow$
+4 Open Issues and Key challenges
+Integrating advanced techniques like DTL, FL, and DRL into ASR systems presents exciting opportunities but comes with its set of challenges.
 This section delves into the distinct challenges associated with each approach, emphasizing the critical areas that demand attention and innovation.
-5.1 DTL and domain adaptation
+4.1 Transformer-based
+Transformer-based ASR systems have revolutionized the field of speech processing with their superior ability to handle sequential data.
+However, their deployment in the acoustic and language domains presents unique challenges and open issues.
+For the acoustic domain, one of the key challenges in the AM is the handling of long audio sequences.
+Transformers require significant memory and computational resources, making it difficult to process lengthy audio inputs efficiently.
+Additionally, the variability in speech, such as accents, speech disorders, and background noise, can affect the robustness of transformer-based ASR systems.
+Developing models that can generalize across these variations without substantial data for each scenario remains an open issue.
+For the language domain, the major challenge in the LM lies in capturing the nuances of different languages and dialects.
+The transformer’s reliance on large amounts of training data poses a challenge for low-resource languages.
+Furthermore, language models need to understand context deeply to accurately predict words in continuous speech, which can be particularly challenging in languages with rich morphology.
+For both domains, integrating acoustic and language models in transformer-based ASR systems to work seamlessly is complex.
+Achieving real-time processing speeds while maintaining high accuracy is a persistent challenge.
+Furthermore, the interpretability of these models is limited, making it difficult to diagnose and correct errors in ASR predictions.
+Addressing the trade-off between model complexity and the ability to deploy these systems on devices with limited computing power is also a critical challenge.
+4.2 DTL and DA-based
 This section discusses challenges and concepts related to DTL and DA in speech recognition, including distribution shift, feature space adaptation, label distribution shift, catastrophic forgetting, domain-invariant feature learning, sample selection bias, and hyperparameter optimization.
@@
 Finding the optimal hyperparameters typically involves solving complex optimization problems using techniques like grid search, random search, or Bayesian optimization on a validation set.
-MMD^2[f(X_S), f(X_T)] = ∥1nS∑_i=1^n_Sϕ(x_S_i) - 1nT∑_j=1^n_Tϕ(x_T_j)^2
-where $\phi(\cdot)$ represents the mapping to a reproducing kernel Hilbert space (RKHS), and $n_{S}$, $n_{T}$ are the numbers of samples in the source and target domains, respectively.
 Moreover, the unification of DTL in ASR poses a challenge due to the varied mathematical formulations used in different studies.
@@
 4.3 FL-based
-FL has significant potential for ASR systems, particularly in enhancing privacy and personalization.
-However, deploying this technology in ASR also comes with a set of challenges.
+FL has significant potential for DTL systems, particularly in enhancing privacy and personalization.
+However, deploying this technology in DTL also comes with a set of challenges.
 Typically, in FL, data is inherently decentralized and can vary significantly across devices.
 This heterogeneity in speech data—due to differences in accents, dialects, languages, and background noise—can make it challenging to train a model that performs well across all nodes.
-Ensuring robustness and generalization of the ASR model under these conditions is a complex task.
+Ensuring robustness and generalization of the DTL model under these conditions is a complex task.
 Moving on, FL requires periodic communication between the central server and the devices to update the model.
-For ASR systems, where models can be quite large, this can result in substantial communication overhead.
+For DTL systems, where models can be quite large, this can result in substantial communication overhead.
 Optimizing the efficiency of these updates, in terms of both bandwidth usage and energy consumption, especially on mobile devices, is a significant challenge.
@@
 Scalability issues, including managing updates from such a large and potentially unreliable network of devices, ensuring consistent model improvements, and handling devices joining or leaving the network, are significant technical hurdles.
-Lastly, in FL, the distribution of data across devices is often non-identically distributed (non-IID).
+Lastly, in FL, the distribution of data across devices is often non-IID.
 This means that the speech data on one device might be very different from that on another, leading to challenges in training a model that generalizes well across all devices.
 Overcoming the bias introduced by non-IID data is a major challenge in FL for ASR.
-5.3 RL-based
-Using RL in ASR systems offers promising avenues for improvement but also presents several challenges.
-Specifically, one of the primary challenges in applying RL to ASR is the issue of sparse and delayed rewards.
+4.4 DRL-based
+Using DRL in ASR systems offers promising avenues for improvement but also presents several challenges.
+Specifically, one of the primary challenges in applying DRL to ASR is the issue of sparse and delayed rewards.
 In many ASR tasks, the system only receives feedback (rewards or penalties) after processing lengthy sequences of speech, making it difficult to attribute the reward to specific actions or decisions.
 This delay complicates the learning process, as the model struggles to identify which actions led to successful outcomes.
-Moreover, balancing exploration, trying new actions to discover their effects, with exploitation, using known actions that yield the best results, is a critical challenge in RL.
+Moreover, balancing exploration, trying new actions to discover their effects, with exploitation, using known actions that yield the best results, is a critical challenge in DRL.
 In the context of ASR, this means the system must balance between adhering to known speech patterns and exploring new patterns or interpretations.
 Overemphasis on exploration can lead to inaccurate transcriptions, while excessive exploitation may prevent the model from adapting to new speakers or accents.
-Additionally, RL models typically require a significant amount of interaction data to learn effectively.
+Additionally, DRL models typically require a significant amount of interaction data to learn effectively.
 In ASR, obtaining large volumes of labeled speech data, especially with user feedback, can be challenging and expensive.
-Additionally, RL algorithms can be sample-inefficient, meaning they need a lot of data before they start performing well, which can be a bottleneck in practical applications.
+Additionally, DRL algorithms can be sample-inefficient, meaning they need a lot of data before they start performing well, which can be a bottleneck in practical applications.
 Moving forward, most ASR systems are built using supervised learning techniques that rely on vast amounts of annotated data.
-Integrating RL into these systems poses technical challenges, as it requires a different training paradigm that focuses on learning from user interactions and feedback rather than static datasets.
-Besides, using RL in ASR often involves collecting and analyzing user feedback and interactions to improve the model.
+Integrating DRL into these systems poses technical challenges, as it requires a different training paradigm that focuses on learning from user interactions and feedback rather than static datasets.
+Besides, using DRL in ASR often involves collecting and analyzing user feedback and interactions to improve the model.
 This raises concerns about user privacy and data security, as sensitive information might be inadvertently captured and used for training.
@@
 Poorly designed reward functions can lead to suboptimal learning outcomes or unintended behaviors.
-Lestly, ASR systems are used in a wide range of environments, from quiet offices to noisy streets.
-RL models need to adapt to these varying conditions, but training them to handle such diversity can be complex.
-The environment’s variability requires models that can generalize well across different acoustic conditions, which remains a challenge for RL-based ASR systems.
+Lastly, ASR systems are used in a wide range of environments, from quiet offices to noisy streets.
+DRL models need to adapt to these varying conditions, but training them to handle such diversity can be complex.
+The environment’s variability requires models that can generalize well across different acoustic conditions, which remains a challenge for DRL-based ASR systems.
 5 Future directions
 5.1 Personalized data augmentation for dysarthric and older people
-While ASR technologies have advanced significantly, especially in recognizing typical speech patterns, they still struggle to accurately identify speech from individuals with dysarthria or older adults [120].
+While DTL technologies have advanced significantly, especially in recognizing typical speech patterns, they still struggle to accurately identify speech from individuals with dysarthria or older adults [121].
 Gathering extensive datasets from these groups is challenging due to mobility limitations often associated with these populations.
 In this context, personalized data augmentation plays a crucial role [122, 123].
-Personalized data augmentation tailores the training process to accommodate the unique speech patterns and challenges associated with these groups.
-Dysarthria, a motor speech disorder, and the natural aging process can lead to speech that deviates from the normative models typically used to train ASR systems, making accurate recognition difficult [123].
+Personalized data augmentation tailors the training process to accommodate the unique speech patterns and challenges associated with these groups.
+Dysarthria, a motor speech disorder, and the natural aging process can lead to speech that deviates from the normative models typically used to train DTL systems, making accurate recognition difficult [124].
 Personalized data augmentation introduces a wider range of speech variations into the training dataset, including those specific to dysarthric speakers or older adults.
 This can include variations in speech rate, pitch, articulation, and clarity.
-By training on this augmented dataset, the ASR system learns to recognize and accurately transcribe speech that exhibits these characteristics [124].
-Moreover, this helps the ASR models generalize better to unseen examples of speech from dysarthric speakers or older adults.
+By training on this augmented dataset, the DTL system learns to recognize and accurately transcribe speech that exhibits these characteristics [125].
+Moreover, this helps the DTL models generalize better to unseen examples of speech from dysarthric speakers or older adults.
 This enhanced generalization is crucial for real-world applications where the system encounters a wide range of speech variations.
@@
 […]
-MTL enhances the performance of ASR systems by leveraging the inherent relatedness of multiple learning tasks to improve the generalization of the primary ASR task.
+MTL enhances the performance of DTL systems by leveraging the inherent relatedness of multiple learning tasks to improve the generalization of the primary ASR task.
 This approach allows the ASR model to learn shared representations that capture underlying patterns across different but related tasks, leading to several key benefits [129].
-Typically, MTL encourages the ASR system to learn representations that are beneficial across multiple tasks.
+Typically, multitask learning (MTL) encourages the ASR system to learn representations that are beneficial across multiple tasks.
 This can lead to more robust feature extraction, as the model is not optimized solely for transcribing speech but also for other related tasks, such as speaker identification or emotion recognition.
 This shared learning process helps in capturing a broader range of speech characteristics, which can improve the ASR system’s ability to handle varied speech inputs.
-Moving one, By simultaneously learning related tasks, MTL acts as a form of regularization, reducing the risk of overfitting on the primary ASR task.
+Moving on, by simultaneously learning related tasks, MTL acts as a form of regularization, reducing the risk of overfitting on the primary ASR task.
 This is because the model must find a solution that performs well across all tasks, preventing it from relying too heavily on noise or idiosyncrasies specific to the training data of the main task.
@@
 […]
-FMTL extends the concept of FL by allowing each client to learn a personalized model that addresses its specific task, while still benefiting from collaboration with other clients.This approach recognizes the heterogeneity in clients’ data distributions and tasks.
+FMTL extends the concept of FL by allowing each client to learn a personalized model that addresses its specific task, while still benefiting from collaboration with other clients.
+This approach recognizes the heterogeneity in clients’ data distributions and tasks.
 Mathematically and compared with FL (Equation 3), federated multi-task learning (FMTL) can be formulated as:
@@
 The diversity of inputs helps in diluting the impact of any adversarial data introduced to compromise the model.
-End-to-End ASR models Continuing to develop and refine end-to-end deep learning models that directly map speech inputs to text outputs without the need for intermediate representations (like phonetic transcriptions) [mamyrbayev2023hybrid].
-Advances in models such as Transformer and Conformer architectures, which can capture long-range dependencies in speech, are promising for improving ASR accuracy and efficiency [liu2023sfa].
-Specifically, end-to-end models in ASR aim to directly convert speech input into text output using a single neural network architecture.
-This simplifies the ASR pipeline by bypassing traditional stages such as acoustic, pronunciation, and language modeling [yang2023attention].
-An end-to-end ASR model maps an input sequence of speech features $X=(x_{1},x_{2},\ldots,x_{T})$ to an output sequence of tokens $Y=(y_{1},y_{2},\ldots,y_{N})$.
-The model function $f$ with parameters $\theta$ aims to minimize the difference between the predicted sequence and the ground truth.
-Two common loss functions are used [almadhor2023e2e]:
-  • CTC Loss: Connectionist Temporal Classification (CTC) introduces a ’blank’ token for aligning sequences. The CTC loss is given by:
-    $\mathcal{L}_{\text{CTC}}(\theta)=-\sum_{(X,Y)\in\mathcal{D}}\log P(Y|X;\theta)$
-    where $\mathcal{D}$ is the dataset, and $P(Y|X;\theta)$ is the sequence probability.
-  • Attention-based Seq2Seq Loss: This uses an encoder-decoder architecture with attention to predict each token. The loss is the negative log-likelihood:
-    $\mathcal{L}_{\text{Seq2Seq}}(\theta)=-\sum_{(X,Y)\in\mathcal{D}}\log P(Y|X;\theta)$
-Besides, model architectures can be described as follows:
-  • Recurrent Neural Networks (RNNs) including Long Short-Term Memory (LSTM) networks for temporal dependencies [vander2023using].
-  • Convolutional Neural Networks (CNNs) for capturing local patterns in speech features.
-  • Transformer Models utilize self-attention mechanisms for parallel processing and superior performance [wang2022optimizing].
-Gradient-based methods like stochastic gradient descent (SGD) or Adam are used to minimize the loss function, optimizing the model parameters $\theta$ to improve the speech-to-text mapping.
-However, end-to-end models face challenges such as data scarcity for low-resource languages, computational resource demands, and integrating external language models for linguistic improvements [qu2023emphasizing].
-6.4 Domain-specific language modeling
-[…]
-[…]
-$P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}$ (8)
-Additionally, data augmentation plays a crucial role in ASR-based DA.
-Typically, data augmentation in domain-specific ASR involves generating synthetic training examples by altering existing recordings or using text-to-speech (TTS) systems to create new audio samples from domain-specific texts.
-Mathematically, augmentation techniques can include time stretching, pitch shifting, adding noise, or simulating room acoustics, each represented as transformations $T$ applied to the original audio signal $x(t)$:
-$x^{\prime}(t)=T(x(t);\phi)$ (9)
-where $\phi$ represents the parameters of the transformation.
-Moving on, regularization techniques are applied to prevent overfitting to the domain-specific data, ensuring the model maintains generalization capabilities.
-Techniques such as dropout, L1/L2 regularization, or elastic net involve adding terms to the loss function or modifying the optimization process to penalize large weights or complex models.
-For instance, L2 regularization adds a penalty equal to the square of the magnitude of coefficients:
-$L^{\prime}(D;\theta)=L(D;\theta)+\lambda\|\theta\|^{2}$ (10)
-where $\lambda$ is a regularization parameter.
-While DSLM offers substantial benefits, challenges remain, such as data scarcity in niche domains, the need for continual model updates to keep pace with evolving language use, and ensuring privacy and security in sensitive domains like healthcare and finance.
-Addressing these challenges through innovative approaches in model training, data augmentation, and privacy-preserving technologies will be crucial for the advancement of domain-specific ASR systems.
-Privacy preservation With the advancements in ML, DL, and DTL, ASR systems have become more practical and scalable.
-However, these systems also introduce serious privacy concerns due to the abundance of sensitive acoustic and textual information in speech data.
-While open-source and offline ASR systems can mitigate privacy risks, online DTL-based systems can amplify these threats.
-Additionally, the transcription performance of offline and open-source ASR systems is typically inferior to cloud-based ASR systems, especially in real-world scenarios [ahmed2020preech].
-In this context, the $D_{S}$ data may contain sensitive information that needs to be protected.
-Thus, preserving users’ privacy during the knowledge transfer from the source domain (SD) to the target domain (TD) becomes a critical issue.
-To address this challenge, future research efforts should focus on integrating effective security and privacy protection strategies.
-Examples include decentralized DTL approaches utilizing blockchain technology [ul2020decentralized, wang2021enabling] and federated DTL methods [zhang2021federated, maurya2021federated].
-These approaches aim to ensure privacy-preserving knowledge transfer in DTL-based ASR systems.
-6.5 Interpretation of DTL models
-DTL-based ASR models, despite their success, are often considered as "black box" systems lacking interpretability.
-This lack of interpretability raises doubts about the credibility and repeatability of their decisions, making it crucial to explain the reasoning behind their predictions.
-The field of explainable and interpretable ML/DL has gained increasing interest in various applications, including speech processing.
-Several studies have explored the interpretability of DTL models.
-In one study [ramakrishnan2016towards], an agent was designed to explain how it learns a new task using prior common knowledge, aiming to enhance users’ trust and acceptance of the system results.
-Another study [kim2019the] defined interpretable features in a DTL algorithm and examined the relationship between the SD and TD in the task, focusing on the interpretability of the pretrained DTL model.
-The work by Lee et al.
-[lee2021interpretable] introduced a knowledge distillation approach that generated interpretable embedding procedure (IEP) knowledge based on PCA and transferred it to the student network using a message passing neural network.
-This approach enhanced interpretability while maintaining accuracy through multi-task learning.
-Additionally, Carr et al.
-proposed an interpretable staged TL (iSTL) scheme for accurate and explainable classification of optical coherence tomography (OCT) scans with a small sample size [carr2021interpretable].
-iSTL outperformed DTL techniques on unseen data, utilizing clinical features for predictions with interpretable attention maps.
 Delving deeper into techniques for FL distillation (optimizing model size) within FL frameworks is essential.
@@
 It is imperative to investigate the trade-offs associated with reducing model size while maintaining performance metrics like WER.
-Developing strategies to strike a balance between downsizing models and preserving satisfactory performance levels within federated learning environments is crucial.
-6.6 Recent RL techniques for ASR
+Developing strategies to strike a balance between downsizing models and preserving satisfactory performance levels within FL environments is crucial.
+5.4 Recent DRL techniques for ASR
 Exploring incremental DRL approaches [137, 138, 139] in DRL-based ASR systems, could be very interesting.
@@
 This capability not only fosters greater resilience and adaptability in speech recognition capabilities but also offers potential applications in scenarios where real-time adaptation to changing conditions is crucial, such as in noisy environments or with varying speaker accents.
-Moreover, incremental RL can potentially lead to more efficient use of computational resources, as the model only needs to focus on learning from new data, rather than reprocessing the entire dataset.
+Moreover, incremental DRL can potentially lead to more efficient use of computational resources, as the model only needs to focus on learning from new data, rather than reprocessing the entire dataset.
 Further research in this area could unlock new possibilities for ASR systems to evolve and improve over time, ultimately enhancing their usability and effectiveness in diverse real-world settings.
-Although some ASR schemes based on RL have been proposed, there remains a notable scarcity in the application of RL techniques to enhance ASR methods.
-While policy gradient and Q-learning are commonly employed, the realm of RL encompasses various subcategories such as double deep Q-network (DDQN), actor-critic (AC), deep deterministic policy gradien (DDPG), and more [139], which hold promise for advancing ASR with innovative approaches.
-Researchers are encouraged to delve into these diverse DRL-based methods to further enrich the field of ASR for both AM and LM fields.
+Although some DTL schemes based on DRL have been proposed, there remains a notable scarcity in the application of DRL techniques to enhance DTL methods.
+While policy gradient and Q-learning are commonly employed, the realm of DRL encompasses various subcategories such as double deep Q-network (DDQN), actor-critic (AC), deep deterministic policy gradien (DDPG), and more [140], which hold promise for advancing DTL with innovative approaches.
+Researchers are encouraged to delve into these diverse DRL-based methods to further enrich the field of DTL for both AM and LM fields.
 5.5 Online DTL
-Online DTL combines the principles of TL and online learning with deep neural networks, enabling models to adapt in real-time to new tasks or data distributions.
+Online DTL combines the principles of DTL and online learning with DNNs, enabling models to adapt in real-time to new tasks or data distributions.
 This approach is beneficial in dynamic environments where data arrives sequentially.
-DL models, specifically Deep Neural Networks (DNNs), learn through optimizing the weights $\theta$ to minimize a loss function $L$, which measures the discrepancy between predicted outputs $\hat{y}$ and true outputs $y$: $\theta^{*}=\arg\min_{\theta}L(D;\theta)$.
-On the other hand, TL improves learning in a new target task through the transfer of knowledge from a related source task, adapting a pre-trained model $\theta_{S}$ on $D_{S}$ to a $D_{T}$: $\theta_{T}^{*}=\arg\min_{\theta}L(D_{T};\theta_{T})$.
+DL models, specifically DNNs, learn through optimizing the weights $\theta$ to minimize a loss function $L$, which measures the discrepancy between predicted outputs $\hat{y}$ and true outputs $y$: $\theta^{*}=\arg\min_{\theta}L(D;\theta)$.
+On the other hand, DTL improves learning in a new target task through the transfer of knowledge from a related source task, adapting a pre-trained model $\theta_{S}$ on $D_{S}$ to a $D_{T}$: $\theta_{T}^{*}=\arg\min_{\theta}L(D_{T};\theta_{T})$.
 In this regard, online learning updates the model incrementally as new data $(x_{t},y_{t})$ arrives:
@@
 Where $\alpha_{t}$ is the learning rate, and $\nabla_{\theta}L$ is the gradient of the loss with respect to $\theta$.
-Moving on, online TL integrates these concepts to continuously adapt a deep learning model to new tasks or data streams, often involving techniques such as feature extraction, fine-tuning, model adaptation, and continual learning.
+Moving on, online DTL integrates these concepts to continuously adapt a deep learning model to new tasks or data streams, often involving techniques such as feature extraction, fine-tuning, model adaptation, and continual learning.
 The adaptation process at each time step $t$ can be viewed as $\theta_{t+1}^{*}=\arg\min_{\theta}L(D_{T_{t}};\theta_{T_{t}})$.
@@
 Furthering the discourse, online DTL augmented with extreme learning machines introduces a novel framework [144].
-Addressing the challenge of limited data in the $D_{T}$, the technique of transfer learning with lag (TLL), rooted in shallow neural network embeddings, has been applied.
+Addressing the challenge of limited data in the $D_{T}$, the technique of DTL with lag, rooted in shallow neural network embeddings, has been applied.
 This method ensures the continuity of knowledge transfer, notwithstanding fluctuations in the feature set.
@@
 […]
-Similarly, the LM aspect of LLM based on DTL can enhance ASR by leveraging TL.
+Similarly, the LM aspect of LLM based on DTL can enhance ASR by leveraging DTL.
 Pre-training LLM on vast text corpora equips it with extensive language representations, aiding in addressing diverse language challenges.
@@
 These advanced techniques address issues related to DA, privacy preservation, and dynamic decision-making, thereby enhancing ASR performance and reducing computational costs.
-This survey has provided a comprehensive review of DTL, FL, and RL-based ASR frameworks, offering insights into the latest developments and helping researchers and professionals understand current challenges.
-Additionally, the integration of transformers, powerful DL models, has been explored for their ability to capture complex dependencies in ASR sequences.
+This survey has provided a comprehensive review of DTL, FL, and DRL-based DTL frameworks, offering insights into the latest developments and helping researchers and professionals understand current challenges.
+Additionally, the integration of Transformers, powerful DL models, has been explored for their ability to capture complex dependencies in DTL sequences.
 By presenting a structured taxonomy and conducting critical analyses, this paper has shed light on the strengths and weaknesses of existing frameworks, as well as highlighted ongoing challenges.
-Moving forward, further research is needed to overcome these challenges and unlock the full potential of advanced DL techniques in ASR.
-Future work should focus on refining existing approaches, addressing privacy concerns in FL, improving RL algorithms for ASR optimization, and exploring innovative ways to leverage transformers for more efficient and accurate speech recognition.
+Moving forward, further research is needed to overcome these challenges and unlock the full potential of advanced DL techniques in DTL.
+Future work should focus on refining existing approaches, addressing privacy concerns in FL, improving DRL algorithms for DTL optimization, and exploring innovative ways to leverage Transformers for more efficient and accurate speech recognition.
 By continuing to innovate and collaborate across disciplines, we can push the boundaries of ASR technology and realize its transformative impact on various fields, including healthcare, communication, and accessibility.